### v1

Abstract
A central challenge in neuroscience is understanding how the brain’s complex nonlinear dynamics give rise to emergent properties underlying adaptive cognition and behavior across spatial and temporal scales. These dynamics exhibit scale-free and multifractal properties, shaping the adaptive reconfiguration of neural networks. However, conventional neuroimaging models remain constrained by assumptions of linearity and stationarity, limiting their capacity to fully characterize these intricate processes. Transformer-based architectures, known for capturing long-range dependencies in sequential data, naturally align with the brain’s intrinsically hierarchical and temporal organization. Here, we introduce Multi-Band Brain Net (MBBN), a transformer-based deep learning framework that explicitly models frequency-specific spatiotemporal brain dynamics from fMRI by integrating scale-free network principles with frequency-resolved multi-band self-attention mechanisms.
Trained on three large-scale neuroimaging cohorts (UK Biobank, ABCD, ABIDE) totaling 45,951 individuals, MBBN reveals previously undetectable frequency-dependent intrinsic network interactions, shedding new light on distinct connectivity disruptions across psychiatric conditions (ADHD, ASD, depression). This extensive validation demonstrates robust generalizability and highlights core neural principles conserved across populations and developmental stages.
Specifically, MBBN achieves up to 30.59% higher predictive accuracy than state-of-the-art deep learning and conventional connectome-based methods, demonstrating the advantage of frequency-informed spatiotemporal modeling in capturing latent neural computations. Furthermore, MBBN’s attention-based interpretability reveals novel frequency-specific biomarkers for neurodevelopmental disorders, offering mechanistic insights into the hierarchical organization of brain function.
By providing a principled, interpretable framework for dissecting spatiotemporal learning mechanisms, MBBN advances methodological rigor and provides new insight into how dynamic neural computations underpin cognitive function and psychiatric vulnerability, with implications for brain decoding, cognitive neuroscience, and precision psychiatry.
keywords:
Spatiotemporal learning, Transformer, fMRI, Knowledge-guided Deep Neural Network, Neural Dynamics
1 Introduction
Understanding the brain’s dynamic processes is essential for unraveling the mechanisms underlying adaptive cognition and behavior. These processes influence critical functions such as learning, decision-making, and social interactions, and their dysregulation is closely linked to mental health conditions ([1, 2, 3, 4]). While functional magnetic resonance imaging (fMRI) has been instrumental in probing brain activity, many conventional analytical methods remain constrained by assumptions of linearity, stationary, static connectivity. Approaches ranging from mass-univariate statistical tests and seed-based correlations to static functional connectivity maps often treat the brain as a relatively fixed system, thus failing to capture the temporal fluctuations, nonlinear interactions, and adaptive reconfigurations that characterize real neural dynamics. Recent advances in artificial intelligence (AI), particularly transformer-based architectures, have begun to address these shortcomings by modeling brain activity as a continuously evolving network of interdependent processes ([5, 6, 7, 8]).
However, a critical gap remains in understanding and modeling two fundamental properties of neural dynamics: scale-free dynamics and multifractal structures. These properties are essential for comprehending the brain’s hierarchical complexity and adaptability ([9, 10, 2]). Scale-free dynamics, characterized by power-law relationships (Equation 1), reveal how neural activity maintains consistent patterns across different timescales and modalities. These dynamics play crucial roles in brain maturation and state transitions, with temporal variability serving as a key regulatory mechanism ([11, 12]). The power-law exponent ($\beta$) varies across brain regions and states, providing valuable insights into neural network heterogeneity and serving as a potential marker for neurological and psychiatric disorders ([13, 14]).
Multifractality extends this framework by incorporating multiple scaling dimensions and power-law behaviors ([15, 16, 17]), offering a sophisticated approach to understanding the brain’s nested, hierarchical organization ([18, 19]). This property manifests in distinct frequency-dependent patterns that reflect neural adaptability and functional specialization ([20]). Evidence from fMRI, ECoG, and EEG studies suggests that these multifractal principles drive frequency-specific changes in connectivity and topological organization, particularly in properties such as small-worldness and modularity ([21, 22]). Understanding these patterns is crucial for developing more effective diagnostic and therapeutic strategies.
To address these challenges, we introduce the Multi-Band Brain Net (MBBN), a novel framework based on Bidirectional Encoder Representations from Transformers (BERT) that integrates scale-free property-based frequency decomposition with advanced self-attention mechanisms and network theory concepts like communicability (Figure 1). MBBN divides fMRI BOLD signals into distinct frequency bands, enabling the capture of band-specific functional and topological properties. This approach provides a neurobiologically informed framework for modeling the brain’s multifractal and scale-free dynamics. The model features two key strengths: Firstly, its self-attention weighted connectivity dynamically models functional connectivity by capturing complex, nonlinear interactions and time-varying relationships between brain regions, thereby overcoming the limitations of traditional static, correlation-based linear representations. Secondly, a communicability-inspired pretraining loss function optimizes for biologically meaningful connectivity patterns. By leveraging these advances, MBBN not only provides deeper insights into brain function but also identifies frequency-specific markers for neurological and psychiatric disorders, offering a powerful new tool for clinical neuroscience research and applications.
$log(power)\propto-\beta\cdot log(frequency)$ (1)
Figure 1: Multi-Band Brain Net procedure
A.
The MBBN module processes each timeseries signal, divided by frequency, within specific frequency ranges. In the temporal module (denoted as BERT), BERT updates the [CLS (Classification)] token, which is then passed through a classifier to generate a prediction $\hat{y_{\nu}}$ ($\nu$ represents the frequency range). The spatial module employs a multi-head attention mechanism to compute attention matrices representing the interactions between regions of interest (ROIs). B. The MBBN modules are applied to timeseries signals across three frequency ranges (ultra-low, low, and high frequency). Parameter sharing is implemented in the temporal module, while spatial modules have independent parameters for each frequency range. The final output of MBBN is computed as the average of the predictions across all frequency ranges, along with attention matrices specific to each frequency range. C. The frequency decomposition process is demonstrated using a representative sample from the UKB dataset. Although the knee frequencies ($f_{1}$, $f_{2}$) may vary across individuals, the overall pattern of frequency division remains consistent. The blue line represents the original power spectral density (PSD) of the fMRI signal. The pink, yellow-green, and yellow lines show the fitted power-law functions for the high, ultra-low, and low-frequency ranges, respectively. The dotted red and green lines indicate the frequency boundaries: ultra-low to low ($f_{1}$), and low to high ($f_{2}$) frequencies, respectively.
2 Results
2.1 Comparison of Model Performance Across Various Downstream Tasks
Figure 2: Comparative Analysis of fMRI Prediction Models. Across datasets and tasks, the proposed MBBN demonstrates superior predictive performance compared to baseline models including state-of-the-art models in the field (e.g., Brain Network Transformer ([5])), as well as vanilla BERT (without multi-fractality division). [CC] indicates models based on Pearson’s correlation-based connectivity; [TD] represents, models accounting for temporal dynamics. The x-axis denotes the models, and the y-axis represents performance metrics, with AUROC used for classification tasks and MAE for regression tasks. (A–B) show results for sex prediction, (C–D) for fluid intelligence prediction, (E–F) for depression prediction, G for ADHD prediction, and H for ASD prediction. Specifically: (A, B, C, G) correspond to the ABCD dataset using the HCP-MMP1 atlas. (D, E, F) correspond to the UKB dataset using the HCP-MMP1 atlas. H correspond to the ABIDE dataset using the HCP-MMP1 atlas.
To evaluate model performance, we trained MBBN from scratch on multiple downstream tasks spanning biological, cognitive, and clinical domains, using both the HCP-MMP1 and Schaefer atlases across the UKB and ABCD datasets. MBBN consistently outperformed baseline models across all tasks. For biological phenotype prediction (sex classification), MBBN achieved notable gains of up to 11.6% on UKB and 18.88% on ABCD using the HCP-MMP1 atlas, with similar improvements using the Schaefer atlas (9.06% on UKB, 14.78% on ABCD). In cognitive assessment tasks, fluid intelligence prediction showed substantial improvements on the ABCD dataset (up to 18.88% with HCP-MMP1, 8.03% with Schaefer). For clinical applications, depression prediction on the UKB dataset demonstrated meaningful gains (up to 9.42% with HCP-MMP1, 6.98% with Schaefer). Comprehensive performance comparisons, including task-specific metrics and statistical analyses, are detailed in Appendix.
2.2 Communicability-based masking loss improves fine-tuning performance over random masking loss.
A key component of our methodology was a pretraining loss function designed to mask nodes with high communicability within the brain network. This strategy was predicated on the hypothesis that selectively reconstructing these highly influential nodes during pretraining would better represent the complex communication patterns within the brain. To validate this hypothesis and evaluate the downstream impact of this pretraining strategy, MBBN was subsequently fine-tuned on two clinically relevant classification tasks: ADHD classification using the ABCD dataset, and ASD classification using the ABIDE dataset. These datasets and tasks were chosen for their established clinical significance and high prevalence rates.
Fine-tuning experiments revealed a clear advantage for the MBBN-high model, which employed a pretraining strategy that selectively masked nodes with high communicability. This approach resulted in significant performance improvements compared to baselines: a 5.55% gain in ADHD classification accuracy and a remarkable 30.59% gain in ASD classification accuracy. In contrast, models using random node masking (MBBN-random) or masking nodes with low communicability (MBBN-low) showed inferior performance, even compared to the MBBN model trained from scratch (MBBN-from scratch). These findings underscore the critical role of communicability-based pretraining in effectively leveraging the underlying network structure of the data to enhance predictive performance. Complete results and statistical details are available in Appendix.
2.3 Key connectivity for mental illness classification revealed by attention mechanisms
Figure 3: Self-Attention-Based Connectivity Learned from Resting FMRI Distinguishing ADHD and ASD from Healthy Controls. For presentational purposes, this figure highlights statistically significant connections with the top 10 largest effect sizes that differentiate individuals with ADHD (A-C) or ASD (D-F) from healthy controls (HC) (FDR corrected P’s ¡ 0.05). [Red: Higher in the disorder group; Blue: Lower in the disorder group.]. A detailed list of the top 100 connectivity is provided in the Appendix F. (A-C) show ADHD connectivity patterns compared to HC across frequency bands: high (A), low (B), and ultralow (C). (D-F) illustrate ASD patterns compared to HC: high (D), low (E), and ultralow (F). The HCP-MMP1 atlas was used.
We utilized a Grad-CAM-based attribution score to identify self-attention-weighted-connectivity with large effect sizes distinguishing healthy controls from individuals with mental illness and analyzed their roles. Details are described in Section 3.9.
In the ABCD test dataset (HCP-MMP1 atlas), frequency-dependent differences in functional connectivity patterns were identified between the ADHD and HC groups. ADHD showed reduced connectivity in high and low frequencies, while exhibiting increased connectivity in the ultralow-frequency range (Figure 3). These findings reveal frequency-specific neural connectivity alterations in ADHD. The results of the t-test distinguishing between the HC and ADHD groups showed that 6.55% of the statistically significant connections ($p<0.05$) were in the high-frequency range, with the 75% quantile value of Cohen’s d score being 0.71. In contrast, the low-frequency range accounted for only 1.67%, with the 75% quantile value of Cohen’s d being 0.17. The ultralow frequency range accounted for 8.39%, with the 75% quantile value of Cohen’s d being 0.71.
In the high-frequency range (Figure 3 A), distinct connectivity patterns emerged involving key regions such as ParaHippocampal Area 3 and VentroMedial Visual Area, highlighting potential disruptions in visual processing and memory formation. Similarly, the low-frequency range (Figure 3 B) showed significant differences centered around Area PH (Parahippocampal Area) and Primary Auditory Cortex , pointing to altered neural system how sound information is integrated and processed in memoryn. In the ultralow-frequency range (Figure 3 C), connectivity changes in VentroMedial Visual Area 3 and Area TF suggest impairments related to visual processing and attention regulation.
In the ABIDE test dataset (HCP-MMP1 atlas), the results of the t-test distinguishing between the HC and ASD groups showed that 7.88% of the statistically significant connections ($p<0.05$) were in the high-frequency range, with the 75% quantile value of Cohen’s d score being 1.19. In contrast, the low-frequency range accounted for only 0.24%, with the 75% quantile value of Cohen’s d being 0.74. The ultralow frequency range accounted for 0.91%, with the 75% quantile value of Cohen’s d being 0.84.
ASD-related connectivity patterns exhibited localized differences compared to healthy controls (HC), with varying effects across frequency ranges (Figure 3). In the high-frequency range (Figure 3 D), connectivity differences were observed in regions such as PGi (PG inferior) and 47m (Area 47 medial), potentially linked to challenges in socio-emotional and executive functions. In the low-frequency range (Figure 3 E), IFSa showed prominent differences, indicating possible impairments in executive and communication functions. Finally, in the ultralow-frequency range (Figure 3 F), connectivity alterations in Lateral Area 7P and the Insular Granular Complex suggest disruptions in sensory and emotional processing.
In both the ABCD and ABIDE datasets, analysis of false positive and false negative subjects (based on predicted values) showed no statistically significant connections in the self-attention-based connectivity analysis.
3 Methods
3.1 Preprocessing
The ABCD, UKB, and ABIDE datasets underwent initial preprocessing prior to analysis. In the ABCD study, a bandpass filter ranging from 0.009 to 0.08 Hz was applied to harmonize multi-site data [23]. In the UKB study, a bandpass filter ranging from 0.008 to 0.1 Hz was used to minimize the effects of heart rate variability and hemodynamic response delays [24]. In the ABIDE study, a finite impulse response (FIR) filter with a passband of 0.01–0.1 Hz [25] was employed to facilitate multi-site data integration. For the present study, additional preprocessing was performed on the resting-state BOLD data. The following preprocessing procedures were applied to each participant’s resting-state BOLD data: First, skull-stripping, slice-timing correction, and spatial normalization to MNI space were conducted using fMRIPrep ([26]). Subsequently, head motion artifacts and signals from white matter were removed using component-based noise correction (CompCor) [27]. The mean time-series signals from predefined regions of interest (ROIs) were extracted using the Nilearn package, with HCP-MMP1 symmetric ([28]) and Schaefer 400 ([29]) atlases employed for ROI definitions. Participants whose data contained voxels with NaN values, typically arising from the regression of unreliable white-matter signals, were excluded from the analysis. Additionally, to minimize spurious signal fluctuations due to scanner instability, the first 20 volumes of each time-series were discarded.
For deep neural network training, uniform sequence lengths across all participants were required. To standardize the sequence length, it was set to the smallest multiple of 8 that was greater than or equal to the shortest sequence length in the dataset, optimizing computational efficiency for GPU-based matrix operations. The final sequence lengths for the datasets were as follows: UK Biobank (UKB) 464, Adolescent Brain Cognitive Development (ABCD) 348, and Autism Brain Imaging Data Exchange (ABIDE) 280.
3.2 Dataset
Demographic and diagnosis distribution details for the three datasets (UKB, ABCD, ABIDE) are summarized in Table 1. The UK Biobank (UKB) dataset, which includes 41,283 participants, was used for pretraining due to its large sample size and broad demographic representation. This extensive dataset enabled the model to capture generalizable brain dynamics patterns. The Adolescent Brain Cognitive Development (ABCD) dataset, consisting of 4,527 participants (including individuals with ADHD and healthy controls), and the Autism Brain Imaging Data Exchange (ABIDE) dataset, which includes 141 participants (comprising individuals with ASD and healthy controls), were employed for fine-tuning to capture disorder-specific patterns.
ADHD diagnosis was based on CBCL attention and ADHD scores exceeding a T-score threshold of 65, while healthy controls were defined as individuals with no diagnosed mental health disorders ([30]). This classification ensured a distinct separation between clinical and non-clinical groups, facilitating model training and evaluation.
Features UKB ABCD ABIDE
Gender (M/F) 40,699 (21,556/19,143) 8,896 (4,662/4,234) 141 (114/27)
Age (year) 54.9$\pm$7.5 - 14.3$\pm$3.6
Age (month) - 119.2$\pm$7.5 -
Depression (ph9 score) 11.5$\pm$3.5 - -
Diagnosis (disorder/controls) - 141 (77/64) -
Fluid intelligence (z-score) - 0.01$\pm$0.91 -
Table 1: Features of the UKB, ABCD, and ABIDE subjects. Disorder denotes ADHD in the ABCD dataset and ASD in the ABIDE dataset.
3.3 Experimental Settings
To address the imbalance in clinical variables within the dataset, stratified sampling was employed to ensure balanced distributions of target variables (ADHD, depression, ASD) across the training, validation, and test sets. A similar approach was used for the biological phenotype (sex), which showed minimal imbalance. To enhance model robustness and reduce overfitting, training was conducted using three random seeds, and the average performance across these runs was reported.
Gradient clipping with a norm of one and gradient accumulation were applied to stabilize training and improve predictive accuracy. To accelerate the training process, Automatic Mixed Precision (AMP) was utilized. The BERT model architecture comprised 8 hidden layers and 8 or 12 attention heads, depending on the number of regions of interest (ROIs) in the atlas used. The number of attention heads was chosen based on the constraint that it should be a divisor of the total number of ROIs. Hyperparameter optimization included testing learning rates ([1e-5, 1e-2]), weight decay ([1e-3, 1e-2]), learning rate policy (step, SGDR), and optimizers (Adam, AdamW, RMSprop).
For computational efficiency, pretraining was performed on a single A100 GPU, while training from scratch and fine-tuning were conducted using a single NVIDIA RTX 3090 GPU. The number of epochs was set to 100 for training from scratch. Efficient pretraining was achieved by determining the optimal number of epochs using weightwatcher ([31]), which identifies well-trained layers based on statistical learning theory. Optimal pretraining epochs were set to 1,000 for the HCP-MMP1 atlas and 400 for the Schaefer 400 atlas. Additional methodological details are provided in Appendix A.2.
3.4 Dividing frequencies
3.4.1 Frequency Dividing Strategy
$Power(f)=\frac{A\cdot f_{1}^{2}}{f^{2}+f_{1}^{2}}$ (2)
Let $w(f)$ be defined as:
$w(f)=\begin{cases}0&\text{if }\log(f)<\log(f_{2})-s\\ 1&\text{if }\log(f)>\log(f_{2})+s\\ \frac{1}{2}(1-\cos(\frac{\pi(\log(f)-\log(f_{2})+s)}{2s}))&\text{otherwise}\end{cases}$ (3)
Then the spline multifractal function is:
$Power(f)=A\cdot f^{\beta_{low}(1-w(f))+\beta_{high}w(f)}$ (4)
where:
  • $A$ is the amplitude.
  • $\beta_{low}$ is the low-frequency scaling exponent.
  • $\beta_{high}$ is the high-frequency scaling exponent.
  • $f_{2}$ is the knee frequency.
  • $s$ is the smoothness parameter.
To segment the resting-state fMRI data into distinct frequency bands, we applied both Lorentzian and spline multifractal equations to the power spectral density (PSD). These methods allowed us to identify knee frequencies ($f_{1}$, $f_{2}$) that demarcate the transitions between frequency bands and reveal their scale-free properties. The Lorentzian equation (Equation 2) was used to estimate $f_{1}$, the first knee frequency, by fitting the entire PSD. For frequencies above $f_{1}$, the spline multifractal equation (Equation 4) was applied to identify $f_{2}$, the second knee frequency. To ensure a smooth transition between different frequency regimes, we implemented cubic spline smoothing in the multifractal equation, preventing artifacts that could arise from abrupt transitions at the knee frequencies. The smoothing weight was adapted based on methodologies from ([32, 33]). The transition function $w(f)$ plays a crucial role in interpolating between two distinct power-law regimes. When $x\ll f_{2}$, $w(f)\approx 0$, ensuring that the function follows the lower-frequency power-law behavior characterized by $\beta_{\text{low}}$. Conversely, when $x\gg f_{2}$, $w(f)\approx 1$, shifting the function to the higher-frequency power-law behavior defined by $\beta_{\text{high}}$. In the intermediate range, $w(f)$ smoothly transitions between these regimes, with the parameter $s$ controlling the sharpness of the transition. This formulation enables a continuous and differentiable shift between scaling regimes while preserving the structural properties of the signal. These knee frequencies were used to segment the time-series data into three frequency bands: ultra-low frequency (below $f_{1}$), low frequency (between $f_{1}$ and $f_{2}$), and high frequency (above $f_{2}$). The segmentation process was implemented using the Finite impulse response (FIR) and Boxcar filter from the nitime package. The optimal filter types (FIR and Boxcar) were determined through systematic hyperparameter tuning on validation sets, evaluating their performance impact on the overall model. The Lorentzian function, having only two parameters, was efficiently fitted using the curve fit function from the SciPy package, which is well-suited for low-dimensional fitting. In contrast, the spline multifractal function, which involves five parameters, was fitted using the iminuit package, which is optimized for high-dimensional parameter estimation. In all three datasets, ANOVA analysis of $f_{1}$ and $f_{2}$ computed separately for each dataset, revealed a statistically significant difference between the two values ($p<0.001$), indicating that $f_{1}$ and $f_{2}$ reflect distinct frequency characteristics.
To assess the scale-free properties of these frequency bands, we computed $\beta$ values for each range, as shown in Equation 1. The results, summarized in Table 3, demonstrated significant differences in $\beta$ values across bands ($p<0.001$), supporting the hypothesis that each frequency component exhibits distinct scale-free dynamics. These results highlight that different frequency bands in the brain exhibit unique functional and topological properties. Specifically, lower frequencies are associated with long-range, global integration of brain activity, while higher frequencies correspond to more localized processing and rapid adaptation ([34, 35]). This finding reinforces the concept of multifractal behavior in brain signals, which has been previously observed in ECoG and EEG recordings ([20]). These $\beta$ values, denoted as $\beta_{ultralow}$, $\beta_{low}$, $\beta_{high}$, correspond to the ultra-low, low, and high-frequency bands, respectively.
3.4.2 Frequency Divison and Physiological Relevance
The values of $f_{1}$ and $f_{2}$ varied slightly across datasets (Table 2), yet consistently fell within specific ranges, reinforcing their physiological relevance. $f_{1}$ (0.046 Hz for ABCD, 0.062 Hz UKB datasets, and 0.018 Hz for ABIDE) corresponds to the time constant of an exponentially decaying function ([36]). Meanwhile, $f_{2}$ marks the transition to higher-frequency brain processes. $f_{2}$ was 0.078Hz for ABCD, 0.096Hz for UKB, and 0.030Hz for ABIDE. In ABCD and UKB, the $f_{2}$ values approaches the upper bound of conventional resting-state fMRI bandpass filtering (0.009-0.08Hz for ABCD, 0.008-0.1Hz for UKB). However, contemporary fMRI preprocessing pipelines commonly employ the Butterworth filter, which applies a gradual roll-off rather than an abrupt cutoff at upper bound (0.08Hz for ABCD, 0.1 Hz for UKB). This design minimizes phase distortion while effectively suppressing physiological noise from respiratory and cardiac ([37]). Consequently, the proximity of $f_{2}$ to upper bound of bandpass filtering does not imply a loss of meaningful higher-frequency information, as these signals are gradually attenuated rather than completely eliminated.
The significant differences in $\beta$ values across frequency bands further validate the frequency-dependent multifractal properties of fMRI signals. This demonstrates the potential of frequency-specific analyses to uncover multi-scale brain network organization and adaptive complexity in brain activity. Additionally, the reliability of $f_{1}$ and $f_{2}$ across different datasets emphasizes the generalizability of this approach, making it a powerful tool for further research into brain dynamics.
Dataset $f_{1}$ $f_{2}$
ABCD 0.0456 $\pm$ 0.0106 0.0784 $\pm$ 0.0151
UKB 0.0624 $\pm$ 0.0113 0.0961 $\pm$ 0.0157
ABIDE 0.0182 $\pm$ 0.0121 0.0299 $\pm$ 0.0256
Table 2: Different $f_{1}$ and $f_{2}$ among datasets
Dataset $\beta_{ultralow}$ $\beta_{low}$ $\beta_{high}$
ABCD 0.2442$\pm$0.2991 1.7845$\pm$0.3836 1.1595$\pm$0.4260
UKB 0.0209$\pm$0.2185 2.8959$\pm$0.4562 1.1232$\pm$0.4107
ABIDE 0.5805$\pm$0.3925 2.5939$\pm$0.3923 0.9563$\pm$0.2228
Table 3: Different $\beta$ value among frequency ranges and datasets
The spline multifractal equation and the Lorentzian equation showed superior fitting performance across all datasets as depicted in Table 4. The consistent and high goodness-of-fit metrics across different datasets ($R^{2}>0.66$ for all cases) demonstrate the effectiveness of our dual-equation fitting approach, validating its reliability for characterizing the frequency-domain properties of resting-state fMRI data.
Dataset Lorentzian Equation Spline Multifractal Equation
Mean $R^{2}$ Mean $R^{2}$
ABCD 0.7423 $\pm$ 0.0765 0.6822 $\pm$ 0.2375
UKB 0.6605 $\pm$ 0.0786 0.8292 $\pm$ 0.0854
ABIDE 0.7435 $\pm$ 0.1018 0.8309 $\pm$ 0.1019
Table 4: Goodness of Fit ($R^{2}$) Statistics for Lorentzian and Spline Multifractal Equations Across Datasets
  • Note: Values are presented as mean±$\pm$±standard deviation. $R^{2}$ values indicate the proportion of variance explained by each model.
3.5 Model Design
3.5.1 Model Architecture
Our model consists of two primary modules: the temporal module and the spatial module. These modules work synergistically to encode both the temporal and spatial characteristics of frequency-divided timeseries data, as shown in Figure 1 A. To ensure an effective balance between generalization and specificity, a parameter-sharing strategy is employed, as outlined below.
The temporal module (depicted as a blue oval in Figure 1 A) is designed to capture the temporal dynamics of the data. For each frequency-divided timeseries (ultra-low, low, and high), the data are processed independently using a shared BERT encoder, reflecting the temporal similarities across frequency bands. The BERT encoder outputs a sequence of hidden states for each timeseries, with the final hidden state corresponding to the [CLS] token passed to a classifier. This process produces logits ($\hat{y_{ultralow}}$, $\hat{y_{low}}$, and $\hat{y_{high}}$), which represent the model’s predictions for each frequency range. The hidden dimension size of the MBBN temporal module is set to match the number of ROIs. Due to the constraints imposed by the BERT backbone in MBBN, the number of attention heads must be a divisor of the hidden dimension size. Consequently, when using the Schaefer atlas, the number of attention heads is set to 8, whereas for the HCP MMP1 atlas, it is set to 12.
The spatial module (depicted as a yellow-green oval in Figure 1 A) focuses on the spatial connectivity among regions of interest (ROIs). Unlike the temporal module, the spatial module uses a frequency-specific approach to capture the distinct spatial patterns of each frequency band. Specifically, a multi-head attention mechanism is applied to the transposed timeseries data, generating a set of attention matrices for each frequency band. These attention matrices—reflecting the distinct focus of different heads—are averaged to produce a single attention matrix per band. These averaged matrices represent the spatial connectivity patterns, which are subsequently used to compute the spatial loss (Equation 9). In practice, the MBBN spatial module receives the transposed input of the temporal module, so the number of attention heads must divide the sequence length. For example, we set 12 heads for ABCD (sequence length is 348) and 8 heads for both UKB (464) and ABIDE (280).
Although the temporal and spatial modules do not directly exchange hidden states, they remain closely coupled through the composite loss function. The total loss combines cross-entropy (or task-specific) loss with the spatial loss, so backpropagation updates parameters across the entire network. As a result, while the BERT-based temporal module primarily absorbs gradient updates from the mask and task objectives, and the attention-based spatial module is mostly driven by spatial loss, their optimization processes are fundamentally interdependent. Changes in one module can indirectly shift the parameter landscape of the other, ensuring a shared optimization of temporal and spatial patterns.
This architectural design supports module specialization—BERT focuses on temporal dependencies, while the attention module models spatial relationships—without direct interference. Yet both modules ultimately contribute to the network’s overall optimization, facilitating a more integrated representation of brain dynamics. Additionally, by treating these modules as partly separable, the framework allows for independent evaluation and targeted enhancements in each module, preserving their complementary roles in the final model.
As shown in Figure 1 B, the temporal and spatial modules operate on each frequency range, generating corresponding logits and attention matrices. To obtain the final prediction, the logits from the three frequency ranges are averaged and compared with the true label ($y_{true}$) using a task-specific loss function ($Loss_{task}$). In parallel, a spatial loss is computed to enforce distinct spatial connectivity patterns across frequency bands. The total loss function, which combines these two components, is defined as:
$Loss_{total}=Loss_{task}(y_{true},\frac{\hat{y_{ultralow}}+\hat{y_{low}}+\hat{y_{high}}}{3})+\lambda\cdot Loss_{spatial}$ (5)
In Equation 5, $Loss_{task}$ represents the cross-entropy loss for classification tasks or L1 loss for regression tasks. The $Loss_{spatial}$ term, as described in Section 9, emphasizes the distinct spatial connectivity patterns inherent to each frequency band. $\lambda$ was determined by hyperparameter tuning. The chosen $\lambda$ value and the filter type (described in Section 3.4.1) were shown in Table 5.
Dataset Filter Type Spatial Factor $\lambda$
Schaefer atlas
ABCD Boxcar 10
UKB Boxcar 1
ABIDE Boxcar 100
HCPMMP1 atlas
ABCD FIR 10
UKB Boxcar 1
ABIDE Boxcar 100
Table 5: Optimal filter types and $\lambda$ determined through hyperparameter tuning
3.5.2 Parameter Sharing
To enhance model performance, we adopted a multi-view learning approach, which is well-suited for frequency-divided timeseries data. We treat the ultra-low, low, and high frequency time series as three ‘views’ of the same underlying data, where different representations of the same underlying phenomenon are considered. Multi-view data enables the model to leverage shared information across views, allowing it to capture complementary aspects of the data.
For the temporal module, we utilized a parameter-sharing strategy. Given that temporal dynamics exhibited minimal differences across frequency bands (as indicated by similar autocorrelation structures), a single shared BERT encoder was used to process timeseries data from all frequency bands. This approach allows the model to exploit common temporal patterns across frequency bands, leading to a robust and generalized representation of temporal dynamics.
In contrast, the spatial module required a frequency-specific approach. Previous studies ([38]) and our analysis have shown that spatial connectivity patterns differ significantly across frequency bands. Therefore, we implemented frequency-specific parameters for the spatial module. Self-attention-weighted connectivity was calculated independently for each frequency range, with each frequency band using its own set of parameters to capture the unique spatial relationships among ROIs. This design ensures that the model preserves the distinct connectivity characteristics of each frequency band while maintaining the interpretability of spatial attention.
This parameter-sharing strategy strikes a balance between generalization and specificity. Shared temporal parameters capture global patterns common across frequency bands, while frequency-specific spatial parameters emphasize the unique spatial relationships within each frequency band. By integrating this approach with task-specific and spatial loss functions (Equation 5), the model effectively captures the multi-level organization of frequency-divided timeseries data, facilitating the identification of relevant biomarkers across multiple scales.
3.6 Temporal BERT Module
3.6.1 Data Encoding
To capture temporal dependencies across regions of interest (ROIs), the frequency-divided time-series data were sequentially input into the BERT model. Each signal at a single time point was treated as a distinct token, with the token length corresponding to the total number of ROIs. This encoding strategy effectively represented both the temporal and spatial dynamics of the fMRI data, ensuring that these features were preserved for downstream analysis.
3.6.2 Using Original Form of BERT
FMRI BOLD signal data consists of long sequences, typically ranging from 280 to 464 tokens, necessitating a model that can handle such extended sequences. BERT, originally developed for natural language processing (NLP) tasks, is particularly well-suited for this purpose due to its ability to capture complex contextual relationships within sequences ([39]). This makes BERT an ideal model for modeling the intricate temporal dependencies inherent in fMRI data.
In contrast, models like GPT are optimized for text generation tasks and do not emphasize inter-sequence contextual relationships, which are critical for our analysis. Unlike GPT’s causal masking—which is tailored to text generation and unidirectional context—BERT’s bidirectional attention allows it to capture long-range temporal dependencies from both past and future time points simultaneously, making it more suitable for modeling complex fMRI time series. Although some previous studies have used convolutional filters as a preprocessing step before entering signals into BERT ([40]), we chose to bypass this step. Although convolutional preprocessing can be computationally efficient, it may sacrifice the preservation of fine-grained temporal dynamics within the fMRI data, which could impact model performance (see Appendix D). By directly inputting the time-series data into the original BERT structure, we aimed to retain the full complexity of the temporal dynamics, allowing the model to effectively capture both short-term and long-term dependencies.
3.7 Spatial Attention Module
3.7.1 Self-Attention-Weighted Connectivity
Traditional methods of functional connectivity based on statistical correlations assume linear relationships and static connectivity ([41]). However, these approaches do not account for the nonlinear and time-varying nature of brain connectivity. To address this limitation, we employed a dynamic method that captures the complex, evolving interactions within resting-state functional networks. This method provides a more accurate representation of the diverse and temporal relationships among brain regions.
Given that spatial connectivity reflects the relationships between brain regions, we hypothesized that a self-attention mechanism, as used in transformer models ([42]), would be a suitable alternative. Self-attention mechanisms are designed to capture relationships between tokens (or regions, in our case), and the attention matrices they generate are diverse, reflecting the multifaceted nature of these relationships. While self-attention with multiple parallel attention heads is referred to as multi-head attention, our implementation specifically employed the multi-head attention mechanism to capture diverse representational information simultaneously. Multi-head attention mechanisms allow neural networks to focus on different aspects of input data simultaneously, similar to how different neural circuits in the brain can process the same sensory information through specialized pathways. In this approach, the input is projected into multiple representation spaces where separate attention computations occur in parallel, enabling the model to capture diverse relationships within the data. This parallel processing architecture has proven remarkably effective for complex tasks like language understanding, where different heads can specialize in detecting various linguistic patterns such as syntactic relationships, semantic connections, or long-range dependencies. In our transformer model architecture, multiple attention heads are employed, with each head generating its own $N_{ROI}\times N_{ROI}$ attention matrix, where $N_{ROI}$ denotes number of ROIs. These attention matrices differ across heads as each head learns to capture different aspects of the relationships between ROIs. For example, one head might focus more strongly on frontal-temporal relationships while another might emphasize occipital-parietal connections. The final attention-based connectivity is then computed by averaging these head-specific attention matrices. To more effectively capture the complex dynamic relationships between ROIs compared to traditional correlation-based methods, we defined self-attention-weighted connectivity as the mean of the attention matrices (Attention matrix $\in\mathbb{R}^{(N_{ROI}\times N_{ROI})}$), where each matrix represents an estimate of spatial connectivity between ROIs.
3.7.2 Spatial Loss
The three frequency-divided components, separated by frequencies ($f_{1}$, $f_{2}$), exhibit more distinct differences in connectivity patterns ([38]). To highlight these differences, we introduced a spatial loss term designed to encourage different spatial connectivity patterns for each frequency range. Specifically, the loss term maximizes the sum of L1 losses between attention matrices derived from different frequency ranges (Equation 9). The use of the negative logarithm in the loss function magnifies smaller differences between frequency bands, ensuring that even subtle distinctions are emphasized during training. This approach enables the model to learn unique spatial connectivity patterns for each frequency band, reflecting the potentially distinct contributions of these bands to brain activity. Moreover, this spatial loss term works in conjunction with other loss functions in the model to strike a balance between generalization and specificity.
$Loss_{high-low}=Loss_{L1}(attmat_{high},attmat_{low})$ (6)
$Loss_{high-ultralow}=Loss_{L1}(attmat_{high},attmat_{ultralow})$ (7)
$Loss_{ultralow-low}=Loss_{L1}(attmat_{ultralow},attmat_{low})$ (8)
$Loss_{spatial}=-log(Loss_{high-low}+Loss_{high-ultralow}+Loss_{ultralow-low})$ (9)
3.8 Pretraining Strategy Based on Communicability
3.8.1 Masking Strategy and Pretraining Loss
The goal of pretraining is to enable the model to learn the underlying structure of brain signal data in a self-supervised manner. To achieve this, we incorporated the concept of communicability into the pretraining process. A detailed definition of communicability used in this study is provided in Appendix A.1.
Our pretraining strategy draws inspiration from masking techniques commonly used in language models to capture contextual relationships and data patterns ([39]). We hypothesized that masking nodes with high communicability would encourage the model to infer the activity of these influential nodes based on the surrounding network dynamics. This approach highlights the role of high-communicability hubs in shaping the overall structure of resting-state networks.
To validate this hypothesis, we compared three masking strategies:
  • Masking nodes with high communicability.
  • Masking nodes with low communicability.
  • Randomly masking regions of interest (ROIs).
Detailed model performance results (Appendix E.1) show that masking high-communicability nodes led to the most effective learning of the data structure, resulting in superior performance in downstream tasks.
Figure 4: MBBN Pretraining Workflow. A. Each frequency-divided timeseries is processed using the MBBN pretraining module, where the temporal module extracts hidden states using a BERT encoder, and the spatial module computes attention matrices via a multi-head attention mechanism. B. Masking strategies: spatial masking nullifies the time series data of the top $k$ nodes with the highest communicability scores, while temporal masking zeros out ROI signals within predefined time windows. C. Pretraining is conducted separately for high, low, and ultralow frequency bands. Parameter sharing is applied in the temporal module but not in the spatial module. Outputs include hidden states and averaged attention matrices for each frequency range.
Pretraining involved two complementary masking strategies: temporal masking and spatial masking (Figure 4). Spatial masking involved nullifying the timeseries data of the top $k$ nodes with the highest communicability scores, where $k$ is less than the total number of ROIs. Temporal masking, on the other hand, initialized all ROI signals to zero within predefined time windows. Both strategies were applied to each frequency band (ultralow, low, and high), denoted as $\nu$.
The masking loss ($Loss_{mask}$) was calculated as the sum of L1 losses between the original BOLD signals ($BOLD_{\nu}$) and their masked counterparts ($maskedBOLD_{\nu}$):
$Loss_{mask}=\sum\limits_{\nu}Loss_{L1}(BOLD_{\nu},maskedBOLD_{\nu})$ (10)
The pretraining loss was the combination of the masking loss and the spatial loss (Equation 9), which enables the model to simultaneously learn node relationships and recover masked regions. $\lambda$ employed in the pretraining loss function was maintained at the same value as that used in the total loss calculation (Equation 5) during from-scratch training.
$Loss_{pretrain}=Loss_{mask}+\lambda\cdot Loss_{spatial}$ (11)
3.8.2 Temporal Masking Optimization
In temporal masking, rather than predicting a single time point, the model is tasked with predicting entire time windows. This strategy makes the task more challenging, encouraging the model to learn temporal dependencies more effectively. Masking was implemented by setting ROI signals to zero within predefined time windows.
To optimize temporal masking, we empirically tested various configurations by adjusting two parameters:
  • $N$: the length of the masked time window.
  • $k$: the distance between consecutive masked windows.
The best performance in downstream tasks was achieved with $N=20$ and $k=1$, where $N$ ranged from 5 to 45 in multiples of 5, and $k$ was an integer between 1 and 5.
3.8.3 Implementation Details and Dataset
Pretraining was conducted using the UK Biobank dataset, with a sequence length of 464. During finetuning, padding was applied to align shorter sequences ($\tau$) to this fixed length, with padding added equally to both ends of the sequence ($(464-\tau)/2$). The outputs from the pretraining process, which include hidden states and attention matrices for each frequency band, were used to initialize the MBBN model for downstream tasks. Figure 4 illustrates the pretraining workflow, including the temporal and spatial masking procedures. Temporal masking is represented by lavender rectangles, while spatial masking is depicted with purple lines.
3.9 Interpretability
MBBN computes attention matrices between Regions of Interest (ROIs) across different frequency ranges. These matrices serve as estimates of functional connectivity, making it crucial to analyze individual connections to understand their contributions to phenotype classification. This analysis offers insights into the neural mechanisms underlying the phenotype of interest.
Traditional attribution methods are not directly applicable to this model, as the relationship between attention matrices and target values is indirect. To address this challenge, we developed a novel approach inspired by GradCAM ([43]), specifically tailored to the spatial attention module described in Section 3.7.
GradCAM typically highlights important regions by computing the gradient of the loss for the target class and weighting it by the output feature maps. Similarly, we compute heatmaps by multiplying the activation values (attention matrices) with their respective gradients, as described in Equation 12. This method highlights specific functional connectivity patterns that are relevant to mental disorders. In Equation 12, $f$ refers to the spatial attention module, ${Loss}_{spatial}$ corresponds to the loss term from Equation 9, $k$ represents the frequency range (high, low, ultralow), and $input_{k}$ denotes the fMRI BOLD data for frequency range $k$ ($Heatmap_{k}\in\mathbb{R}^{(N_{ROI}\times N_{ROI})}$).
To interpret the model’s predictions, we selected the best-performing model based on the validation dataset and applied it to the test dataset. Heatmaps were generated for true positive subjects. Analysis for false positive and false negative subjects was conducted to validate the method. For inter-group comparisons, we conducted t-tests to identify statistically significant connections ($p<0.05$). For multiple comparison correction, we applied False Discovery Rate (FDR) (Benjamin-Hochberg method) to control for Type I errors. Then we calculated the effect size using Cohen’s d ([44]). The sign of the t-statistic indicated which group had a higher attribution score for a given connection. For example, in the classification of ADHD using the ABCD dataset, a positive t-statistic suggested that a connection contributed more significantly to the Healthy Control (HC) group than to the ADHD group.
Statistically significant connections ($p<0.05$) and regions were analyzed, but the number of connections was too large to handle effectively. To address this, we selected only a few top connections based on the absolute values of Cohen’s d for further evaluation of their impact. An exhaustive list of all significant connections can be found in Appendix F.
$Heatmap_{k}=\frac{\partial{Loss}_{spatial}}{\partial f_{k}(input_{k})}\odot f_{k}(input_{k})$ (12)
3.10 Performance Comparison with Baseline Models
To evaluate the effectiveness of MBBN in learning spatiotemporal dynamics, we compared its performance with baseline models, including functional connectivity-based models (XGBoost [45], BNT [5]) and temporal dynamics models (BolT ([6]), vanilla BERT ([39])). These models were selected based on their relevance to brain dynamics interpretation, either through static functional connectivity or sequential processing of temporal data. While functional connectivity-based models primarily rely on static representations, temporal dynamics models capture sequential patterns but do not explicitly integrate spatial relationships. MBBN addresses this gap by combining both spatiotemporal characteristics using its frequency-dividing encoder and spatial attention module. Given that MBBN is an extension of vanilla BERT with additional architectural enhancements, a direct comparison between the two provides a robust evaluation of these modifications. While MBBN demonstrated superior performance across several downstream tasks, including the ABCD ADHD classification task, with significant improvements in accuracy and AUC scores, not all tasks showed a statistically significant advantage. Detailed performance comparison is described in Appendix E.1.
To ensure a fair evaluation, we conducted statistical tests (paired t-tests) to compare the performance of MBBN with baseline models on each task. The p-values for these comparisons indicated that, although MBBN did not outperform all baseline models in every metric, the differences were not statistically significant ($p>0.05$) in cases where baseline models showed slightly better performance. These results suggest that MBBN encodes information at a level comparable to or better than existing models while providing the additional advantage of capturing spatiotemporal dynamics, which baseline models are unable to incorporate.
4 Discussion
Discussions should be brief and focused. In some disciplines use of Discussion or ‘Conclusion’ is interchangeable. It is not mandatory to use both. Some journals prefer a section ‘Results and Discussion’ followed by a section ‘Conclusion’. Please refer to Journal-level guidance for any specific requirements. This study introduces MBBN, a novel computational framework that integrates neuroscientific principles with advanced deep learning methodologies to analyze brain dynamics. MBBN employs a scale-free, frequency-based decomposition technique that encodes distinct frequency components as individual features, effectively capturing the multi-fractal dynamics of brain activity. Additionally, the model leverages a communicability-based masking strategy during pretraining, embedding the underlying network topology into its learned representations. Through a self-attention mechanism, MBBN combines biologically informed spatial relationships with temporal context, enabling the nuanced modeling of connectivity patterns and their evolution across multiple frequency ranges. This unified framework provides a robust platform for accurate, interpretable, and personalized predictions of cognitive and clinical outcomes.
Compared to previous transformer-based models for brain imaging data (e.g., [8, 7, 5, 6]), MBBN uniquely integrates the multi-fractal ([46, 47, 48]) and scale-free properties of neural signals ([36, 49]), essential for understanding the hierarchical and dynamic nature of brain activity ([50, 51, 52]). Unlike models relying solely on static or sequential representations ([5, 6]), or employing simplistic masking strategies ([7, 8]), MBBN’s biologically grounded approach—decomposing multifractal neural signals into interpretable frequency components and utilizing a communicability-driven masking scheme—enables it to effectively capture and reconstruct critical network hubs. By doing so, MBBN provides a more nuanced and biologically relevant representation of spatiotemporal connectivity.
Crucially, correlation analysis confirms that the connectivity derived from MBBN’s self-attention mechanism represents a novel measure, capturing information distinct from and complementary to traditional functional connectivity (FC). To elucidate this, we directly compared attention-based connectivity with FC. This revealed a strong negative correlation when analyzing raw, non-frequency-divided signals (Spearman $\rho\approx-0.6$), indicating that attention-based connectivity captures fundamentally different, complementary information rather than merely replicating static FC patterns. Of note, the correlations became considerably weaker within specific frequency bands (ranging from $\rho\approx-0.1\backsim-0.2$) (Supplementary Table 2). These findings strongly suggest that MBBN successfully extracts unique, frequency-dependent connectivity patterns that are distinct from both traditional FC and the aggregated signal. This divergence not only validates our attention mechanism as providing a novel and meaningful measure of neural interaction but also underscores the importance of frequency-specific analysis for uncovering nuanced brain dynamics potentially missed by conventional methods. The biological plausibility of MBBN is further supported by its incorporation of the scale-free property, characterized by a power-law exponent ($\beta$), which differentiates brain states and regions based on their temporal correlation profiles ([2, 53, 54, 55, 56, 57]). Variations in $\beta$ values are associated with functional specialization and pathological alterations ([13, 14, 9, 58]), underscoring the importance of integrating these intrinsic properties into computational models. By incorporating communicability-based masking, MBBN systematically accounts for graph-structured brain data, addressing the limitations of random masking strategies ([59, 60]) and enhancing understanding of network-level interactions ([61, 62]).
When applied to graph-structured functional brain data, random masking fails to consider node importance and the non-Euclidean nature of the data ([59, 60]), leading to uncertainties in reconstructing masked information. MBBN overcomes these limitations through a communicability-based masking strategy, which quantifies information flow across all possible network paths and emphasizes the role of high-communicability hubs in shaping network dynamics ([61, 62]). By selectively masking these critical hubs and reconstructing their activity through interactions with surrounding nodes, MBBN learns structural dependencies intrinsic to neural systems, enhancing both interpretability and robustness. This approach establishes MBBN as a foundational framework for advancing neuroimaging analysis.
MBBN demonstrated superior performance across diverse datasets, including UK Biobank, ABCD, and ABIDE, underscoring its robustness across different age groups and developmental stages. While UK Biobank includes primarily middle-aged adults, ABCD and ABIDE focus on children and adolescents during critical brain maturation periods. Unlike most functional brain transformer models ([8, 7, 5]), which are rarely validated on heterogeneous cohorts, MBBN effectively handles lifespan variability, suggesting that it captures fundamental principles of brain organization. This enables accurate predictions of clinical and cognitive outcomes across ages and provides insights into developmental trajectories and age-specific neurological markers.
Our findings revealed distinct, frequency-specific connectivity alterations distinguishing individuals with ADHD and ASD from healthy controls. In ADHD, problematic connectivity was observed in regions essential for visual and audio processing, memory formation, and attention regulation with disruptions varying significantly across frequency ranges.
In the high-frequency range, we observed significant reductions in connectivity between the Parahippocampal area 3 and multiple early visual regions, as well as between the Ventromedial visual area 2 and primary/third visual cortices. These findings imply impaired integration between visual information and memory or contextual processing systems, which may be linked to difficulties in sustaining attention, interpreting visual scenes, and integrating incoming stimuli in ADHD ([63, 64]). In the low-frequency range, decreased connectivity emerged between the Parahippocampal area and anterior temporal, perirhinal, and orbitofrontal regions—networks typically involved in episodic memory, semantic integration, and affective regulation ([65, 66]). Additionally, weakened connections between the primary auditory cortex and multimodal association areas point to disruptions in sensory processing ([67]). These alterations suggest that ADHD may be characterized by deficits in slower, long-range communication necessary for high-level cognitive integration and stable internal state representation. Conversely, the ultralow-frequency range showed increased connectivity in ADHD, particularly between the Ventromedial visual area 3 and spatial attention regions (e.g., Lateral intraparietal area ventral, medial area 7A), as well as between emotion-related regions such as Area TF (the posterior subdivision of the Parahippocampal cortex), Area 25 (Subgenual cingulate cortex), and the anterior cingulate cortex (s32). This pattern may reflect excessive internal processing, emotional dysregulation, and inefficient attentional gating ([68]). Together, these findings indicate that ADHD involves a dynamic reconfiguration of functional brain networks across multiple timescales—with reduced integration in fast and intermediate-range communication, and hyperconnectivity in the slowest neural rhythms—potentially contributing to the core symptoms of inattention, emotional lability, and sensory processing anomalies.
Our frequency-specific analysis of the ABIDE dataset for ASD prediction revealed divergent patterns of functional connectivity alterations across spectral bands, underscoring the multi-timescale nature of network disruptions in autism spectrum disorder. In the high-frequency range, the ASD group exhibited markedly increased connectivity, particularly involving Area PGs (Posterior superior parietal lobule ), Area 8 anterior ventral (within the prefrontal cortex), and Area 8Ad (Anterior dorsal prefrontal cortex) with regions implicated in executive control and higher-order social cognition (e.g., Area p32, Area 9 Middle, Dorsal Area 24d). These patterns suggest hyperconnectivity in frontoparietal and cingulo-opercular circuits, potentially reflecting compensatory or maladaptive neural synchronization related to cognitive control and social attention ([69, 70]). In the low-frequency range, increased connectivity was observed between multiple sensory and associative areas (e.g., Area PH, Primary Sensory Cortex, Auditory and ParaBelt Complexes) and Area posterior 9–46 ventral. This may reflect overintegration or inefficient filtering in low-frequency sensory-cognitive loops, contributing to the sensory hypersensitivity and perceptual integration challenges frequently reported in ASD ([71, 72]). Conversely, the ultralow-frequency range showed decreased connectivity in the ASD group compared to controls, notably in connections between Parahippocampal area 1, Area PH, and auditory complexes with Area 7PC (Precuneus) and Area Frontal Opercular 5. These connections are implicated in high-level integrative functions such as spatial awareness, self-referential processing, and multimodal integration. Reduced ultralow-frequency connectivity may reflect a breakdown in global neural coordination and internal state regulation, core deficits commonly implicated in autism ([73, 74]). Together, these findings reveal a frequency-dependent reconfiguration of functional brain networks in ASD, with fronto-executive hyperconnectivity at higher frequencies and reduced long-range integrative coupling at the slowest rhythms—highlighting potential biomarkers and targets for intervention across multiple neurophysiological timescales.
Connectivity patterns observed across frequency bands revealed a transition from local to more widespread disruptions in both ADHD and ASD. In ADHD, high-frequency alterations were primarily local, focusing on visual and memory-related circuits, while ultralow-frequency changes spanned broader fronto-parietal and cingulo-limbic networks. Similarly, ASD showed local hyperconnectivity in high and low-frequency bands, especially in frontal and sensory areas, with broader hypoconnectivity emerging only in the ultralow-frequency range. These results suggest that while both disorders exhibit frequency-specific connectivity alterations, the spatial scale of disruption expands with slower neural oscillations.
In sum, MBBN’s multifaceted and biologically informed approach to modeling spatiotemporal brain dynamics represents a significant advancement over existing transformer-based frameworks. Its comprehensive integration of neuroscientific principles, self-attention mechanisms, and frequency-specific analyses positions MBBN as a promising foundation for more accurate, interpretable, and clinically meaningful neuroscience research.
Data and Code Availability
The code is available on Github (https://github.com/Transconnectome/MBBN). The UKB, ABCD, and ABIDE datasets are publicly available to researchers.
Author Contributions
Sangyoon Bae conceptualized the research, implemented the core algorithms, and conducted all experiments. Junbeom Kwon contributed to code optimization and manuscript preparation.
Declarations
Nothing to declare.
Funding
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. 2021R1C1C1006503, RS-2023-00266787, RS-2023-00265406, RS-2024-00421268), by Creative-Pioneering Researchers Program through Seoul National University(No. 200-20240057), by Semi-Supervised Learning Research Grant by SAMSUNG(No.A0426-20220118), by Identify the network of brain preparation steps for concentration Research Grant by LooxidLabs(No.339-20230001), by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)] by the MSIT(Ministry of Science, ICT), Korea, under the Global Research Support Program in the Digital Field program(RS-2024-00421268) supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation), by the National Supercomputing Center with supercomputing resources including technical support(KSC-2023-CRE-0568) and by the Ministry of Education of the Republic of Korea and the National Research Foundation of Korea (NRF-2021S1A3A2A02090597), and by Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT(MSIT, Korea) & Gwangju Metropolitan City.
Appendix A Pretraining
A.1 Communicability
Communicability encompasses not only the shortest paths but also other types of walks between nodes $p$ and $q$ within a network. A network can be represented as a graph $G=(V,E)$, where $V$ represents nodes and $E$ represents the edges connecting the nodes. Let us denote the number of nodes as $n$ and the number of edges as $m$. The adjacency matrix of the graph is $A(G)=A$, where $A_{ij}$ takes a value of 1 if nodes $i$ and $j$ are connected and 0 otherwise. The communicability between nodes $p$ and $q$ in the network, denoted as $G_{pq}$, incorporates eigenvalues and eigenvectors of the adjacency matrix. Let $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$ be the eigenvalues in non-increasing order, and let $\phi_{j}(p)$ represent the $p$-th element of the $j$-th orthonormal eigenvector corresponding to eigenvalue $\lambda_{j}$. Then, communicability is calculated as follows:
$G_{pq}=\sum_{j=1}^{\infty}\phi_{j}(p)\phi_{j}(q)e^{\lambda_{j}}$ (13)
Communicability provides a measure of how efficiently information flows between nodes by considering not just direct connections but also indirect pathways weighted by their importance. This property makes communicability particularly suitable for identifying high-influence nodes in brain networks, which are then used in our pretraining masking strategy.
A.2 Effective Pretraining: Finding the Optimal Number of Epochs
Supplementary Figure 1: Alpha Value Changes by Pretraining Epochs (HCPMMP1 Atlas, Epochs: 100-600)
Determining the optimal number of epochs for pretraining is crucial for creating an effective pretraining strategy. We utilized weightwatcher ([31]), a diagnostic tool based on traditional statistical learning theory and statistical mechanics, to identify well-trained models. Weightwatcher analyzes the norms of the weight matrix $W$ and parameters of power-law (PL) fit on the eigenvalues of $W$ to assess training quality without requiring data.
According to the Heavy-Tailed Self-Regularization (HT-SR) Theory, lower PL exponent ($\alpha$) values indicate stronger implicit self-regularization, leading to better generalization. Consistently low $\alpha$ values across layers allow effective correlation flow, enabling the gradual transformation of data correlations through each layer. As shown in previous work, layers with $\alpha$ values below 6 are deemed well-trained, while those exceeding 6 are considered poorly trained.
In the following figures, the x-axis represents the layers, and the y-axis represents the $\alpha$ values. The red dotted line marks the threshold of $\alpha=6$. The number of layers with $\alpha$ values above 6 converges to a single layer after epoch 500 (Supplementary Figure 1). We hypothesize that the epoch at which the pattern of layers with $\alpha$ values exceeding 6 stabilizes marks the optimal pretraining epoch. In our monitoring process, we utilized WeightWatcher to analyze the pretrained model at intervals of 100 epochs. We identified the critical epoch $n$ at which the number of layer with $\alpha>6$ converges. Subsequently, we conducted a series of fine-tuning experiments at 10-epoch intervals between epochs $n-100$ and $n$ to empirically determine the optimal epoch for model performance. In HCPMMP1 atlas, $n$ was 600.
Selecting these epochs ensures that most layers in the model achieve sufficient self-regularization, improving overall generalization performance during fine-tuning.
Appendix B Interpretation of Network Properties
Measure Modularity Small-worldness Parameters
FC Attention FC Attention
$\gamma$ $\lambda$ $\sigma$ $\gamma$ $\lambda$ $\sigma$
High frequency 0.0278 0.0576 4.7683 1.4239 3.3487 9.3960 0.9220 10.1913
Low frequency 0.0251 0.0701 4.8803 1.4140 3.4514 5.6272 0.9953 5.6539
Ultralow frequency 0.0350 0.0991 4.8600 1.4794 3.2850 8.3929 0.9229 9.0941
Raw (Non-dividing) 0.0630 -0.0000 7.7804 1.6143 4.8197 5.9527 0.9254 6.4329
Supplementary Table 1: Comparison of Network Properties between FC and Attention Matrix
Our analysis of network properties provides strong evidence that frequency-specific attention matrices capture meaningful connectivity patterns that differ from traditional functional connectivity (FC).
  • Modularity quantifies the strength of division of a network into modules or communities. It measures how well a network can be decomposed into distinct subgroups, with higher values indicating more distinct community structures where connections within communities are denser than connections between communities. Formally, it is defined as the fraction of edges that fall within given communities minus the expected fraction if edges were distributed at random while preserving node degrees.
  • Small-worldness parameters characterize the topological properties of a network that exhibits both high clustering and short path lengths:
    • Gamma ($\gamma$): The clustering coefficient ratio comparing the observed network to a random network with the same number of nodes and edges. Higher values indicate greater local connectivity and clustering, reflecting more efficient local information processing.
    • Lambda ($\lambda$): The characteristic path length ratio comparing the observed network to a random network. Values closer to 1 indicate efficient global integration and information transfer across the network.
    • Sigma ($\sigma$): The small-worldness index calculated as $\gamma/\lambda$. Values significantly greater than 1 indicate that the network exhibits small-world properties, characterized by high local clustering and relatively short path lengths. Small-world networks are considered optimal for information processing, combining efficient local processing with global integration.
These network metrics provide quantitative measures for assessing the topological organization of brain networks and are widely used in neuroimaging studies to characterize functional and structural connectivity patterns. Several key observations can be made from Supplementary Table 1:
  1. 1.
    Frequency-dependent modularity: The attention matrices exhibit distinct modularity and small-worldness patterns across frequency bands (high, low, and ultralow frequencies). A clear trend of increasing modularity with decreasing frequency (0.0576 → 0.0701 → 0.0991) is observed. This pattern aligns with findings from [34], which demonstrated that resting-state BOLD signals display frequency-specific network architecture, where networks that are finely subdivided in lower frequency bands become integrated into fewer networks in higher frequency bands rather than reconfigured. Such evidence supports the observation that modularity values are higher in lower frequency bands, consistent with our results.
  2. 2.
    Contrasting patterns with FC: While FC exhibits its highest modularity in the raw (unfiltered) condition (0.0630), attention-based connectivity shows essentially no modularity in the raw condition (-0.0000). This striking difference suggests that our attention mechanism captures frequency-specific network organization that is lost when analyzing the full frequency spectrum together.
  3. 3.
    Small-worldness characteristics: The small-worldness parameters ($\gamma$, $\lambda$, and $\sigma$) reveal distinct network architectures between FC and attention-based connectivity. Notably, attention matrices consistently show higher clustering ($\gamma$) and small-worldness ($\sigma$) values across frequency bands compared to FC, while maintaining more optimal path length properties ($\lambda$ closer to 1). This indicates that attention-based connectivity may better preserve the small-world architecture that is characteristic of brain networks.
  4. 4.
    Frequency-specific information: Raw data without frequency decomposition fails to exhibit meaningful modularity in the attention matrix, whereas frequency-decomposed data reveals distinct modular structures. These findings suggest effective coupling between the spatial module and frequency-dividing strategy, enabling the extraction of neurobiologically relevant patterns. This observation underscores the importance of frequency-specific analysis when employing attention-based connectivity measures to characterize brain network organization.
These findings support our claim that frequency-specific attention matrices represent a valid and interpretable new form of connectivity measure, capturing network properties that complement traditional FC measures.
Appendix C Correlation Analysis Between FC and Attention-Based Connectivity
To further investigate the relationship between functional connectivity (FC) and our proposed attention-based connectivity measure, we conducted correlation analyses across different frequency bands. Spearman correlation measures the monotonic relationship between two variables by evaluating the rank order association, making it robust to outliers and nonlinear relationships. Kendall tau is a similar non-parametric measure that assesses the ordinal association between two variables by analyzing the number of concordant and discordant pairs, providing a measure of agreement in rankings. The results are presented in Table 2.
Frequency Band Spearman Correlation Kendall Tau
Raw (Non-dividing) -0.6081 -0.4367
High frequency -0.0952 -0.0700
Low frequency -0.1459 -0.1072
Ultralow frequency -0.1973 -0.1451
Supplementary Table 2: Correlation Analysis Between FC and Attention Matrix
The correlation analysis reveals several important insights:
  1. 1.
    Complementary information: The strong negative correlation between FC and attention-based connectivity in the raw condition (Spearman $\rho=-0.6081$, Kendall $\tau=-0.4367$) indicates that these two measures capture fundamentally different aspects of brain connectivity. This negative correlation suggests that our attention-based connectivity is not merely replicating traditional FC but providing complementary information.
  2. 2.
    Frequency-specific patterns: When examining frequency-specific bands, the correlations become considerably weaker (ranging from approximately -0.10 to -0.20), suggesting that each frequency band captures unique connectivity patterns.
  3. 3.
    Methodological validity: The consistent negative correlations across all frequency bands, albeit weaker in the frequency-specific analyses, indicate that our attention mechanism systematically extracts connectivity information that differs from traditional FC. This suggests that attention-based connectivity represents a valid new form of connectivity measure with its interpretable meaning.
These correlation findings, combined with the network property analysis, provide strong evidence that our frequency-specific attention-based connectivity measure captures meaningful and interpretable connectivity patterns that complement traditional FC analysis. The method does not simply maximize differences between frequencies but extracts biologically relevant connectivity information specific to each frequency band.
Appendix D Ablations
D.1 The performance of the model when provided with timeseries data of a single frequency range.
ABCD sex classification UKB sex classification
Data AUROC ↑ Accuracy ↑ AUROC ↑ Accuracy ↑
Z-scored original signal 0.901$\pm$0.003 0.814$\pm$0.0031 0.942$\pm$0.014 0.876$\pm$0.006
High Frequency 0.888$\pm$0.007 0.797$\pm$0.011 0.933$\pm$0.007 0.857$\pm$0.003
Low frequency 0.884$\pm$0.005 0.784$\pm$0.011 0.936$\pm$0.004 0.850$\pm$0.015
Ultra-low Frequency 0.870$\pm$0.010 0.779$\pm$0.013 0.935$\pm$0.003 0.862$\pm$0.007
Divided Frequencies 0.916$\pm$0.004 0.844$\pm$0.006 0.975$\pm$0.001 0.920$\pm$0.001
Supplementary Table 3: Results of BERT models trained with single frequency range in the timeseries data extracted by HCP-MMP1 atlas
The core innovation of our model lies in encoding signals through frequency division and merging them within the modeling process. To identify the source of performance improvement, we conducted ablation studies focusing on single-frequency range data.
First, we examined whether frequency-divided timeseries data alone could capture the dominant signals essential for phenotype prediction. If successful, single-frequency data would outperform other inputs. However, as shown in Table 3, single-frequency data consistently underperformed compared to the z-scored original signal. This finding suggests that the model does not rely solely on dominant signals from specific ranges but rather leverages contributions from all three frequency ranges.
Second, we analyzed the impact of merging signals from multiple frequency ranges. Results indicate that integrating signals within the model, through parameter sharing across multi-views, significantly enhances performance compared to using frequency-separated data alone. This demonstrates the importance of combining signals at the model level rather than isolating them at the data preprocessing stage.
D.2 Treating fMRI signal as a sequence of discrete words.
ABCD sex classification UKB sex classification
Model AUROC ↑ Accuracy ↑ AUROC ↑ Accuracy ↑
MBBN with conv 0.825$\pm$0.009 0.726$\pm$0.009 0.947$\pm$0.003 0.870$\pm$0.002
MBBN without conv 0.916$\pm$0.004 0.844$\pm$0.006 0.975$\pm$0.001 0.920$\pm$0.001
Supplementary Table 4: Results of BERT models with a convolutional layer before BERT model
Determining whether fMRI data should be treated as continuous variables, akin to speech, or discrete sequences, similar to natural language, is a critical design choice. Given the high sample rate of fMRI signals, we hypothesized that they resemble written sentences more closely than speech signals.
To evaluate this, we introduced convolutional layers before feeding the input into the model, inspired by speech recognition frameworks such as Wav2Vec [40]. Table 4 reveals that treating fMRI signals as discrete word sequences yields superior performance compared to processing them as continuous variables. This result highlights the effectiveness of viewing fMRI signals as structured sequences analogous to natural language.
D.3 Frequency-dividing strategies.
ABCD sex classification UKB sex classification
Dividing Strategy AUROC ↑ Accuracy ↑ AUROC ↑ Accuracy ↑
No division 0.901$\pm$0.003 0.814$\pm$0.003 0.942$\pm$0.014 0.876$\pm$0.006
Random 2 Freq 0.903$\pm$0.004 0.818$\pm$0.004 0.969$\pm$0.003 0.911$\pm$0.003
Lorentizan-based 2 Freq 0.903$\pm$0.006 0.805$\pm$0.010 0.973$\pm$0.007 0.918$\pm$0.005
Random 3 Freq 0.900$\pm$0.006 0.803$\pm$0.008 0.972$\pm$0.002 0.918$\pm$0.003
Ours 0.916$\pm$0.004 0.844$\pm$0.006 0.975$\pm$0.001 0.920$\pm$0.001
Supplementary Table 5: Results of BERT variants trained with different frequency-dividing strategies in the timeseries data extracted by HCP-MMP1 atlas
To optimize the frequency-dividing strategy, we compared methods using Lorentzian and spline multifractal equations with alternative approaches. Results are presented in Table 5.
First, we tested random division by selecting two random cutoff frequencies from a normal distribution. Compared to this baseline (Random 3 Freq), our method (Ours) improved AUROC by 1.11%, demonstrating the effectiveness of Lorentzian and spline multifractal equations.
Second, we evaluated division into two frequency ranges instead of three. Using Lorentzian-based boundaries (Lorentzian-based 2 Freq) led to reasonable performance but was outperformed by our three-range strategy, with a 1.44% improvement in AUROC. This emphasizes the importance of high-frequency signals in predicting phenotypes.
In conclusion, dividing frequencies using Lorentzian and spline multifractal equations into three ranges yields superior performance, highlighting the biological significance of high-frequency components.
D.4 Spatial module and temporal module.
ABCD sex classification UKB sex classification
Module AUROC ↑ Accuracy ↑ AUROC ↑ Accuracy ↑
Spatial 0.553$\pm$0.007 0.540$\pm$0.018 0.856$\pm$0.004 0.778$\pm$0.004
Temporal 0.909$\pm$0.007 0.831$\pm$0.011 0.964$\pm$0.001 0.915$\pm$0.001
Spatiotemporal (MBBN) 0.916$\pm$0.004 0.844$\pm$0.006 0.975$\pm$0.001 0.921$\pm$0.006
Supplementary Table 6: Results of spatial and temporal modules of MBBN in the timeseries data extracted by HCP-MMP1 atlas
Table 6 illustrates the performance of individual spatial and temporal modules compared to the full MBBN spatiotemporal architecture on ABCD/UKB sex classification tasks. The results demonstrate that spatial and temporal modules interact synergistically through joint parameter updates during training. When used independently, both modules show diminished performance compared to the integrated spatiotemporal approach (MBBN). The spatial module achieves moderate performance, while the temporal module performs considerably worse on the ABCD dataset. These findings suggest that the neural dynamics captured in fMRI data require both spatial and temporal processing to maximize predictive power. The superior performance of the integrated MBBN architecture implies that allowing these modules to interact through shared loss functions and simultaneous parameter updates creates representations that better capture the complex spatiotemporal patterns in brain activity. This emphasizes the importance of modeling both spatial and temporal dimensions of neural data rather than treating them as independent components.
D.5 Number of parameters and FLOPs
Module AUROC ↑ Accuracy ↑ FLOPs Number of Params
Vanilla BERT (FLOPs-matched) 0.886$\pm$0.000 0.807$\pm$0.008 70,489,026,464 29,894,149
Vanilla BERT (Params-matched) 0.884$\pm$0.006 0.778$\pm$0.027 23,497,736,096 10,050,757
MBBN 0.916$\pm$0.004 0.844$\pm$0.006 70,528,086,432 11,140,695
Supplementary Table 7: Performance of vanilla BERT with parameter count or FLOPs matched to MBBN levels. Experiments were conducted on the ABCD dataset, and task was sex classification.
Table 7 demonstrates a fair comparison between MBBN and vanilla BERT models with matched computational complexity. Two variants of vanilla BERT were created: one matching MBBN’s FLOPs (70.5B vs 70.5B) and another matching MBBN’s parameter count (10.1M vs 11.1M). Despite having similar computational resources or parameter counts, MBBN significantly outperforms both Vanilla BERT variants in AUROC and accuracy. These results strongly suggest that MBBN’s superior performance cannot be attributed simply to increased model capacity or computational complexity. Rather, the architectural innovations in MBBN appear to be responsible for its enhanced effectiveness compared to vanilla BERT baselines.
Appendix E Comparison with other models
E.1 Perormance comparison
ABCD sex UKB sex
Models (data form) AUROC ↑ Accuracy ↑ AUROC ↑ Accuracy ↑
XGBOOST (CC) 0.768$\pm$0.004 0.696$\pm$0.001 0.859$\pm$003 0.768$\pm$0.005
BNT (CC) 0.895$\pm$0.001 0.814$\pm$0.010 0.950$\pm$0.004 0.864$\pm$0.017
BolT (TD) 0.883$\pm$0.001 0.795$\pm$0.001 0.950$\pm$0.000 0.878$\pm$0.001
vanilla BERT (TD) 0.884$\pm$0.006 0.778$\pm$0.027 0.942$\pm$0.014 0.876$\pm$0.006
MBBN (TD) 0.916$\pm$0.004 0.844$\pm$0.006 0.975$\pm$0.001 0.921$\pm$0.006
Supplementary Table 8: Results of models trained from scratch predicting biological phenotypes. Atlas is HCP MMP 1. CC denotes correlation-based connectivity and TD denotes temporal dynamics.
ABCD sex UKB sex
Models AUROC ↑ Accuracy ↑ AUROC ↑ Accuracy ↑
XGBOOST (CC) 0.812$\pm$0.004 0.734$\pm$0.003 0.905$\pm$0.004 0.884$\pm$0.002
BNT (CC) 0.920$\pm$0.006 0.839$\pm$0.012 0.982$\pm$0.001 0.934$\pm$0.092
BolT (TD) 0.911$\pm$0.002 0.826$\pm$0.003 0.982$\pm$0.000 0.932$\pm$0.001
vanilla BERT (TD) 0.901$\pm$0.007 0.818$\pm$0.017 0.980$\pm$0.001 0.925$\pm$0.005
MBBN (TD) 0.918$\pm$0.002 0.831$\pm$0.009 0.987$\pm$0.001 0.934$\pm$0.008
Supplementary Table 9: Results of models trained from scratch predicting biological phenotypes. Atlas is Schaefer 400, 17 Networks. CC denotes correlation-based connectivity and TD denotes temporal dynamics.
UKB depression ABCD depression
Models (data form) MAE ↓ MSE ↓ AUROC ↑ Accuracy ↑
XGBOOST (CC) 2.417$\pm$0.025 12.342$\pm$0.151 0.548$\pm$0.026 0.829$\pm$0.000
BNT (CC) 2.276$\pm$0.015 13.989$\pm$0.682 0.536$\pm$0.008 0.849$\pm$0.003
BolT (TD) 2.217$\pm$0.022 12.729$\pm$0.290 0.541$\pm$0.016 0.843$\pm$0.003
vanilla BERT (TD) 2.194$\pm$0.039 12.751$\pm$0.780 0.5147$\pm$0.014 0.847$\pm$0.003
MBBN (TD) 2.210 $\pm$0.041 13.205$\pm$0.956 0.556$\pm$0.025 0.718$\pm$0.028
Supplementary Table 10: Results of models trained from scratch predicting clinical outcomes, atlas is HCP-MMP1. CC denotes correlation-based connectivity and TD denotes temporal dynamics.
UKB depression ABCD depression
Models (data form) MAE ↓ MSE ↓ AUROC ↑ Accuracy ↑
XGBOOST (CC) 2.408$\pm$0.002 11.942$\pm$0.052 0.581$\pm$0.012 0.833$\pm$0.003
BNT (CC) 2.329$\pm$0.012 15.179$\pm$2.730 0.579$\pm$0.018 0.842$\pm$0.010
BolT (TD) 2.201$\pm$0.004 13.044$\pm$0.301 0.585$\pm$0.011 0.844$\pm$0.005
vanilla BERT (TD) 2.230$\pm$0.006 13.359$\pm$0.479 0.586$\pm$0.040 0.846$\pm$0.004
MBBN (TD) 2.249 $\pm$0.005 14.486$\pm$0.282 0.641$\pm$0.018 0.777$\pm$0.013
Supplementary Table 11: Results of models trained from scratch predicting clinical outcomes, atlas is Schaefer 400, 17 Networks. CC denotes correlation-based connectivity, and TD denotes temporal dynamics.
UKB fluid intelligence ABCD fluid intelligence
Models (data form) MAE ↓ MSE ↓ MAE ↓ MSE ↓
XGBOOST (CC) 1.609$\pm$0.009 4.078$\pm$0.086 0.787$\pm$0.002 0.987$\pm$0.033
BNT (CC) 1.581$\pm$0.023 3.993$\pm$0.516 0.793$\pm$0.008 1.061$\pm$0.093
BolT (TD) 1.589$\pm$0.002 3.973$\pm$0.006 0.728$\pm$0.004 0.845$\pm$0.010
vanilla BERT (TD) 1.597$\pm$0.015 4.081$\pm$0.094 0.722$\pm$0.005 0.838$\pm$0.036
MBBN (TD) 1.630 $\pm$0.006 4.225$\pm$0.125 0.717$\pm$0.012 0.825$\pm$0.034
Supplementary Table 12: Results of models trained from scratch predicting cognitive outcomes, atlas is HCPMMP1. CC denotes correlation-based connectivity and TD denotes temporal dynamics.
UKB fluid intelligence ABCD fluid intelligence
Models (data form) MAE ↓ MSE ↓ MAE ↓ MSE ↓
XGBOOST (CC) 1.600$\pm$0.002 4.003$\pm$0.025 0.779$\pm$0.008 0.970$\pm$0.011
BNT (CC) 1.552$\pm$0.019 4.459$\pm$2.127 0.781$\pm$0.020 0.845$\pm$0.086
BolT (TD) 1.576$\pm$0.006 3.919$\pm$0.013 0.718$\pm$0.005 0.822$\pm$0.016
vanilla BERT (TD) 1.585$\pm$0.020 3.985$\pm$0.105 0.721$\pm$0.018 0.823$\pm$0.039
MBBN (TD) 1.628$\pm$0.051 4.205$\pm$0.282 0.719$\pm$0.013 0.828$\pm$0.028
Supplementary Table 13: Results of models trained from scratch predicting cognitive outcomes, atlas is Schaefer 400, 17 Networks. CC denotes correlation-based connectivity, and TD denotes temporal dynamics.
ABCD ADHD classification ABIDE ASD classification
Models (data form) AUROC ↑ Accuracy ↑ AUROC ↑ Accuracy ↑
XGBOOST (CC) 0.598$\pm$0.010 0.570$\pm$0.005 0.531$\pm$0.020 0.487$\pm$0.006
BNT (CC) 0.634$\pm$0.008 0.606$\pm$0.014 0.575$\pm$0.013 0.567$\pm$0.019
BolT (TD) 0.603$\pm$0.005 0.592$\pm$0.010 0.650$\pm$0.064 0.598$\pm$0.057
Vanilla BERT (TD) 0.609$\pm$0.020 0.592$\pm$0.015 0.583$\pm$0.004 0.542$\pm$0.021
MBBN [from scratch] (TD) 0.598$\pm$0.009 0.589$\pm$0.024 0.654$\pm$0.134 0.521$\pm$0.059
MBBN [random] (TD) 0.603$\pm$0.032 0.589$\pm$0.026 0.601$\pm$0.180 0.534$\pm$0.082
MBBN [low] (TD) 0.615$\pm$0.009 0.602$\pm$0.011 0.583$\pm$0.084 0.547$\pm$0.098
MBBN [high] (TD) 0.628$\pm$0.001 0.609$\pm$0.011 0.672$\pm$0.079 0.646$\pm$0.078
Supplementary Table 14: Finetuning results of pre-trained BERT models with different pretraining loss (HCP-MMP1 atlas). CC denotes correlation-based connectivity, TD denotes temporal dynamics, and MBBN represents our proposed model. The [random], [low], and [high] variants indicate different pretraining strategies based on masking node communicability.
ABCD ADHD classification ABIDE ASD classification
Models (data form) AUROC ↑ Accuracy ↑ AUROC ↑ Accuracy ↑
XGBOOST (CC) 0.594$\pm$0.008 0.564$\pm$0.006 0.660$\pm$0.023 0.629$\pm$0.030
BNT (CC) 0.620$\pm$0.018 0.599$\pm$0.018 0.715$\pm$0.045 0.652$\pm$0.077
BolT (TD) 0.599$\pm$0.010 0.589$\pm$0.011 0.608$\pm$0.028 0.578$\pm$0.027
Vanilla BERT (TD) 0.604$\pm$0.017 0.600$\pm$0.008 0.664$\pm$0.148 0.667$\pm$0.059
MBBN [from scratch] (TD) 0.614$\pm$0.020 0.610$\pm$0.011 0.800$\pm$0.067 0.563$\pm$0.051
MBBN [random] (TD) 0.596$\pm$0.014 0.593$\pm$0.032 0.696$\pm$0.052 0.563$\pm$0.088
MBBN [low] (TD) 0.584$\pm$0.033 0.574$\pm$0.034 0.500$\pm$0.051 0.479$\pm$0.059
MBBN [high] (TD) 0.618$\pm$0.019 0.580$\pm$0.035 0.822$\pm$0.019 0.563$\pm$0.051
Supplementary Table 15: Finetuning results of pre-trained BERT models with different pretraining loss (Schaefer 400 atlas). CC denotes correlation-based connectivity, TD denotes temporal dynamics, and MBBN represents our proposed model. The [random], [low], and [high] variants indicate different pretraining strategies based on masking node communicability.
The tables above (Supplementary Table 8, 9, 10, 11, 12, 13) present a comparison of various models, atlases, and metrics for different downstream tasks trained from scratch. In these tables, (CC) refers to correlation-based connectivity, and (TD) represents temporal dynamics. Specifically, (CC) denotes Pearson correlation computed on temporal dynamics. Across most tasks, MBBN (highlighted in color) generally outperforms or shows comparable performance to other baseline models.
For instance, in the ABCD depression classification task, the dataset’s imbalance in the depression-HC ratio naturally leads to high regular accuracy. However, MBBN demonstrates robustness against this imbalance by achieving more balanced performance metrics compared to other models (Table 10, 11).
Tables 14 and 15 present a comparison across models and metrics for various fine-tuning tasks. Here, MBBN is compared under different pretraining strategies:
  • MBBN [from scratch]: Trained without pretraining.
  • MBBN [low]: Pretrained with masking loss that targets nodes with low communicability.
  • MBBN [random]: Pretrained with random masking loss.
  • MBBN [high]: Pretrained with masking loss that targets nodes with high communicability.
MBBN [high], highlighted in color, consistently outperforms or matches the performance of other baseline models and pretraining strategies across various tasks. This result emphasizes the effectiveness of the high-communicability masking strategy in enhancing model generalization and fine-tuning performance.
E.2 FLOPs and number of parameters
The parameter count difference between our proposed MBBN model and the baseline models (Table 16) stems from our architectural design, which processes data through three separate streams. This design choice was intentional to capture complex patterns through specialized processing pathways. While this results in a higher parameter count, our ablation studies (Appendix D.5) demonstrate that the performance improvements arise primarily from the model’s architecture rather than simply from increased capacity.
Models (data form) Number of parameters FLOPs
XGBOOST (CC) 6100 400
BNT (CC) 3,787,250 1,805,681,216
BolT (TD) 9,597,062 4,226,217,120
Vanilla BERT (TD) 10,050,757 23,497,736,096
MBBN (TD) 11,140,695 70,528,086,432
Supplementary Table 16: The table shows the number of parameters and FLOPs of baseline models used in this study and MBBN (our model).
Appendix F Interpretability
In the following tables, two regions involved in connectivity (based on HCP-MMP1) are presented, along with whether the connectivity shows higher scores in the disorder group or the healthy control (HC) group. Additionally, corrected p-values and Cohen’s d scores were rounded to three decimal places. Due to space limitations, only the top 100 connectivity pairs are displayed.
Supplementary Table 17: Top connectivity features in high-frequency range of ABCD ADHD prediction.
Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
Fourth Visual Area ParaHippocampal Area 3 $<0.001$ HC 0.654
Second Visual Area ParaHippocampal Area 3 $<0.001$ HC 0.645
Polar 10p ProStriate Area $<0.001$ HC 0.643
Primary Visual Cortex VentroMedial Visual Area 2 $<0.001$ HC 0.638
Area PFt ParaHippocampal Area 3 $<0.001$ HC 0.636
Third Visual Area VentroMedial Visual Area 2 $<0.001$ HC 0.636
Second Visual Area VentroMedial Visual Area 2 $<0.001$ HC 0.633
Second Visual Area ProStriate Area $<0.001$ HC 0.633
Area Lateral Occipital 2 VentroMedial Visual Area 2 $<0.001$ HC 0.627
Third Visual Area VentroMedial Visual Area 1 $<0.001$ HC 0.627
Area 11l ProStriate Area $<0.001$ HC 0.626
Third Visual Area ParaHippocampal Area 3 $<0.001$ HC 0.623
Primary Visual Cortex ParaHippocampal Area 3 $<0.001$ HC 0.618
Posterior OFC Complex Frontal Opercular Area 3 $<0.001$ HC 0.616
Area 2 ParaHippocampal Area 3 $<0.001$ HC 0.613
Posterior OFC Complex Anterior Agranular Insula Complex $<0.001$ HC 0.612
Third Visual Area ProStriate Area $<0.001$ HC 0.611
Fourth Visual Area VentroMedial Visual Area 2 $<0.001$ HC 0.61
Area posterior 10p Frontal Opercular Area 4 $<0.001$ HC 0.606
Area 47m Frontal Opercular Area 4 $<0.001$ HC 0.606
Posterior OFC Complex Frontal Opercular Area 4 $<0.001$ HC 0.604
Anterior IntraParietal Area VentroMedial Visual Area 1 $<0.001$ HC 0.601
Area 47l (47 lateral) Frontal Opercular Area 4 $<0.001$ HC 0.597
Area 10r Frontal Opercular Area 4 $<0.001$ HC 0.596
Area 10r Frontal Opercular Area 3 $<0.001$ HC 0.589
Area 11l Dorsal Transitional Visual Area $<0.001$ HC 0.589
Primary Visual Cortex VentroMedial Visual Area 1 $<0.001$ HC 0.589
Area TemporoParietoOccipital Junction 1 ProStriate Area $<0.001$ HC 0.587
Second Visual Area VentroMedial Visual Area 1 $<0.001$ HC 0.587
Orbital Frontal Complex Frontal Opercular Area 4 $<0.001$ HC 0.587
Area 13l Anterior Agranular Insula Complex $<0.001$ HC 0.587
Fourth Visual Area VentroMedial Visual Area 1 $<0.001$ HC 0.584
Posterior InferoTemporal complex VentroMedial Visual Area 2 $<0.001$ HC 0.582
Orbital Frontal Complex Anterior Agranular Insula Complex $<0.001$ HC 0.581
Polar 10p Area 23c $<0.001$ HC 0.578
Polar 10p VentroMedial Visual Area 1 $<0.001$ HC 0.578
Fourth Visual Area Frontal Opercular Area 3 $<0.001$ HC 0.577
Area 31pd VentroMedial Visual Area 1 $<0.001$ HC 0.575
Area Lateral Occipital 1 ProStriate Area $<0.001$ HC 0.575
Insular Granular Complex ProStriate Area $<0.001$ HC 0.574
Primary Auditory Cortex ProStriate Area $<0.001$ HC 0.574
Area 5L ParaHippocampal Area 3 $<0.001$ HC 0.572
Area 47m ProStriate Area $<0.001$ HC 0.571
Anterior IntraParietal Area ParaHippocampal Area 3 $<0.001$ HC 0.57
Area Lateral Occipital 1 ParaHippocampal Area 3 $<0.001$ HC 0.569
Area V4t ParaHippocampal Area 3 $<0.001$ HC 0.568
Area dorsal 23 a+b ParaHippocampal Area 3 $<0.001$ HC 0.567
Area 10v Frontal Opercular Area 4 $<0.001$ HC 0.567
Area V3B Anterior Ventral Insular Area $<0.001$ HC 0.565
Posterior InferoTemporal complex ParaHippocampal Area 3 $<0.001$ HC 0.564
Area 31pd ParaHippocampal Area 3 $<0.001$ HC 0.563
Area V3CD ProStriate Area $<0.001$ HC 0.562
Area V3B Anterior Agranular Insula Complex $<0.001$ HC 0.562
Area V3CD PeriSylvian Language Area $<0.001$ HC 0.562
Middle Temporal Area VentroMedial Visual Area 1 $<0.001$ HC 0.561
Medial Belt Complex ProStriate Area $<0.001$ HC 0.561
Fourth Visual Area ProStriate Area $<0.001$ HC 0.561
Area TemporoParietoOccipital Junction 2 VentroMedial Visual Area 1 $<0.001$ HC 0.56
VentroMedial Visual Area 1 VentroMedial Visual Area 2 $<0.001$ HC 0.559
Polar 10p Area V3B $<0.001$ HC 0.559
Area 52 ProStriate Area $<0.001$ HC 0.559
Third Visual Area Ventral Visual Complex $<0.001$ HC 0.557
Area Lateral Occipital 1 PeriSylvian Language Area $<0.001$ HC 0.556
Area 47l (47 lateral) ProStriate Area $<0.001$ HC 0.556
Area 10v Frontal Opercular Area 3 $<0.001$ HC 0.555
Area 10v Area 23c $<0.001$ HC 0.554
Fourth Visual Area PeriSylvian Language Area $<0.001$ HC 0.553
Area 11l Sixth Visual Area $<0.001$ HC 0.553
Area OP4-PV ParaHippocampal Area 3 $<0.001$ HC 0.553
Third Visual Area PeriSylvian Language Area $<0.001$ HC 0.552
Lateral Area 7A ParaHippocampal Area 3 $<0.001$ HC 0.552
Area 10r Middle Insular Area $<0.001$ HC 0.551
Medial Area 7P ProStriate Area $<0.001$ HC 0.55
Area PGs VentroMedial Visual Area 1 $<0.001$ HC 0.55
Area Lateral IntraParietal dorsal Primary Visual Cortex $<0.001$ HC 0.55
Area 47l (47 lateral) Anterior Ventral Insular Area $<0.001$ HC 0.549
Area TF Area TF $<0.001$ HC 0.549
Medial Area 7P ParaHippocampal Area 3 $<0.001$ HC 0.549
Primary Visual Cortex ProStriate Area $<0.001$ HC 0.549
PreCuneus Visual Area VentroMedial Visual Area 1 $<0.001$ HC 0.549
Area 7PC ParaHippocampal Area 3 $<0.001$ HC 0.549
Area Lateral Occipital 1 Anterior Ventral Insular Area $<0.001$ HC 0.548
Area IntraParietal 0 ParaHippocampal Area 3 $<0.001$ HC 0.548
Area Lateral Occipital 2 ParaHippocampal Area 3 $<0.001$ HC 0.547
Area TE1 posterior Dorsal Transitional Visual Area $<0.001$ HC 0.547
Area PGs ProStriate Area $<0.001$ HC 0.545
Area 31a ParaHippocampal Area 3 $<0.001$ HC 0.545
Area 10v VentroMedial Visual Area 1 $<0.001$ HC 0.545
Area 23d ParaHippocampal Area 3 $<0.001$ HC 0.545
Area FST ParaHippocampal Area 3 $<0.001$ HC 0.545
Area TemporoParietoOccipital Junction 3 VentroMedial Visual Area 1 $<0.001$ HC 0.543
Area 13l Frontal Opercular Area 4 $<0.001$ HC 0.543
Area Lateral Occipital 1 VentroMedial Visual Area 1 $<0.001$ HC 0.543
VentroMedial Visual Area 1 VentroMedial Visual Area 1 $<0.001$ HC 0.542
Area V3CD Anterior Ventral Insular Area $<0.001$ HC 0.542
Area V3CD ParaHippocampal Area 3 $<0.001$ HC 0.542
Area Frontal Opercular 5 ProStriate Area $<0.001$ HC 0.541
Area 31p ventral ParaHippocampal Area 3 $<0.001$ HC 0.541
Area Lateral Occipital 3 Frontal Opercular Area 3 $<0.001$ HC 0.541
Polar 10p Dorsal Transitional Visual Area $<0.001$ HC 0.541
Supplementary Table 18: Top connectivity features in low frequency range of ABCD ADHD prediction
Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
Area TG dorsal Area PH $<0.001$ HC 0.684
Area TE1 anterior Area PH $<0.001$ HC 0.669
Area anterior 47r Area PH $<0.001$ HC 0.661
Area TG Ventral Area PH $<0.001$ HC 0.65
Primary Auditory Cortex Area $<0.001$ HC 0.646
Perirhinal Ectorhinal Cortex Area PH $<0.001$ HC 0.628
Primary Auditory Cortex Area V3CD $<0.001$ HC 0.625
Area TE2 anterior Area PH $<0.001$ HC 0.62
Orbital Frontal Complex Area PH $<0.001$ HC 0.614
Primary Auditory Cortex Area anterior 10p $<0.001$ HC 0.61
Primary Auditory Cortex Area PH $<0.001$ HC 0.602
Area TG dorsal Area V4t $<0.001$ HC 0.602
Dorsal Area 24d Area PH $<0.001$ HC 0.601
Area 10d Area PH $<0.001$ HC 0.594
ParaHippocampal Area 3 Area PH $<0.001$ HC 0.593
Area anterior 47r Area V4t $<0.001$ HC 0.592
Area Area PH $<0.001$ HC 0.585
ProStriate Area Area anterior 47r $<0.001$ HC 0.585
Area TE2 posterior Area V4t $<0.001$ HC 0.585
Area TG dorsal Area $<0.001$ HC 0.579
Area V4t Area PH $<0.001$ HC 0.579
Area 45 Area $<0.001$ HC 0.577
Area 47m Area $<0.001$ HC 0.577
Frontal Opercular Area 2 Area $<0.001$ HC 0.576
Primary Auditory Cortex Area anterior 47r $<0.001$ HC 0.575
Area TG Ventral Area V4t $<0.001$ HC 0.568
Polar 10p Area PH $<0.001$ HC 0.566
ProStriate Area Area V4t $<0.001$ HC 0.564
Area 47s Area PH $<0.001$ HC 0.564
Area TG dorsal Fusiform Face Complex $<0.001$ HC 0.563
Area anterior 47r Area $<0.001$ HC 0.558
Area STGa Area PH $<0.001$ HC 0.558
Area 10v Area PH $<0.001$ HC 0.557
Primary Auditory Cortex Area $<0.001$ HC 0.557
Frontal Opercular Area 3 Area $<0.001$ HC 0.553
Area OP1-SII Area anterior 47r $<0.001$ HC 0.552
Area TemporoParietoOccipital Junction 1 Area anterior 47r $<0.001$ HC 0.552
Superior Temporal Visual Area Area anterior 47r $<0.001$ HC 0.552
Area TE2 posterior Area PH $<0.001$ HC 0.55
ProStriate Area Area $<0.001$ HC 0.547
Fusiform Face Complex Area anterior 47r $<0.001$ HC 0.542
Area STGa Area $<0.001$ HC 0.541
Area posterior 47r Area PH $<0.001$ HC 0.54
Area 25 Area PH $<0.001$ HC 0.54
Auditory 4 Complex Area V3CD $<0.001$ HC 0.539
Area 10d Area V4t $<0.001$ HC 0.539
Primary Auditory Cortex Area V4t $<0.001$ HC 0.538
Area anterior 47r Area anterior 47r $<0.001$ HC 0.538
ParaHippocampal Area 3 Area anterior 47r $<0.001$ HC 0.536
Dorsal area 6 Area PH $<0.001$ HC 0.535
Area anterior 10p Area V4t $<0.001$ HC 0.535
Area 47m Area PH $<0.001$ HC 0.535
Area 8B Area PH $<0.001$ HC 0.535
Primary Visual Cortex Area PH $<0.001$ HC 0.534
Area STSv anterior Area PH $<0.001$ HC 0.534
Area TF Area PH $<0.001$ HC 0.534
Area V6A Area PH $<0.001$ HC 0.533
Area TG dorsal Area V3CD $<0.001$ HC 0.533
Middle Temporal Area Area PH $<0.001$ HC 0.533
Area OP2-3-VS Area anterior 47r $<0.001$ HC 0.533
Area 47s Area $<0.001$ HC 0.533
Posterior OFC Complex Area PH $<0.001$ HC 0.532
Perirhinal Ectorhinal Cortex Area V4t $<0.001$ HC 0.532
Area 10v Area $<0.001$ HC 0.532
Area posterior 10p Area PH $<0.001$ HC 0.531
Area anterior 10p Area PH $<0.001$ HC 0.531
Area 47l (47 lateral) Area PH $<0.001$ HC 0.53
Area 9 Middle Area PH $<0.001$ HC 0.53
Area STSd posterior Area PH $<0.001$ HC 0.53
Area TE2 posterior Area anterior 47r $<0.001$ HC 0.529
Area TE1 anterior Area V4t $<0.001$ HC 0.529
Area PHT Area PH $<0.001$ HC 0.528
Area PFt Area anterior 47r $<0.001$ HC 0.527
Area TE1 Middle Area PH $<0.001$ HC 0.526
Area STSd anterior Area PH $<0.001$ HC 0.526
Area 47s Area $<0.001$ HC 0.525
RetroSplenial Complex Area PH $<0.001$ HC 0.525
ParaHippocampal Area 1 Area PH $<0.001$ HC 0.525
Entorhinal Cortex Area PH $<0.001$ HC 0.524
Supplementary and Cingulate Eye Field Area $<0.001$ HC 0.524
Area PGp Area PH $<0.001$ HC 0.524
Area TG dorsal Area $<0.001$ HC 0.524
Dorsal area 6 Area $<0.001$ HC 0.524
Area anterior 47r Fourth Visual Area $<0.001$ HC 0.523
Area PFcm Area PH $<0.001$ HC 0.523
Supplementary and Cingulate Eye Field Area V3CD $<0.001$ HC 0.523
Area 10r Area PH $<0.001$ HC 0.522
Area Area PH $<0.001$ HC 0.521
Eighth Visual Area Area $<0.001$ HC 0.521
Area TE2 posterior Area $<0.001$ HC 0.52
VentroMedial Visual Area 1 Polar 10p $<0.001$ HC 0.519
Ventral Visual Complex Area TG Ventral $<0.001$ HC 0.519
Area 31p ventral Area PH $<0.001$ HC 0.519
Fourth Visual Area Area PH $<0.001$ HC 0.519
ParaHippocampal Area 3 Fourth Visual Area $<0.001$ HC 0.518
Area 7m Area PH $<0.001$ HC 0.518
Area p32 Area PH $<0.001$ HC 0.517
Auditory 4 Complex Area $<0.001$ HC 0.517
Area FST Area PH $<0.001$ HC 0.517
Area anterior 9-46v Area PH $<0.001$ HC 0.516
Supplementary Table 19: Top connectivity features in ultralow frequency range of ABCD ADHD prediction
Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
Area Lateral IntraParietal ventral VentroMedial Visual Area 3 $<0.001$ ADHD 1.109
Area s32 ParaHippocampal Area 3 $<0.001$ ADHD 1.062
Area TF Area 25 $<0.001$ ADHD 1.028
Inferior 6-8 Transitional Area VentroMedial Visual Area 3 $<0.001$ ADHD 1.016
Area TF Area s32 $<0.001$ ADHD 1.007
Ventral Visual Complex Area V3B $<0.001$ ADHD 1.002
Area TA2 VentroMedial Visual Area 3 $<0.001$ ADHD 1.002
Medial Area 7A VentroMedial Visual Area 3 $<0.001$ ADHD 1.0
Area 25 Area TF $<0.001$ ADHD 0.994
Area IntraParietal 0 VentroMedial Visual Area 3 $<0.001$ ADHD 0.994
Area 25 ParaHippocampal Area 3 $<0.001$ ADHD 0.993
Premotor Eye Field VentroMedial Visual Area 3 $<0.001$ ADHD 0.99
Area TF Orbital Frontal Complex $<0.001$ ADHD 0.99
Frontal Eye Fields VentroMedial Visual Area 3 $<0.001$ ADHD 0.979
Area 13l Area TF $<0.001$ ADHD 0.975
Area Lateral IntraParietal ventral Ventral Visual Complex $<0.001$ ADHD 0.969
Ventral Area 6 VentroMedial Visual Area 3 $<0.001$ ADHD 0.968
Frontal Eye Fields Eighth Visual Area $<0.001$ ADHD 0.968
Area 10r Area TF $<0.001$ ADHD 0.963
Medial Area 7A Eighth Visual Area $<0.001$ ADHD 0.96
Lateral Area 7P VentroMedial Visual Area 3 $<0.001$ ADHD 0.959
Inferior 6-8 Transitional Area Ventral Visual Complex $<0.001$ ADHD 0.955
Premotor Eye Field Ventral Visual Complex $<0.001$ ADHD 0.955
Area posterior 9-46v VentroMedial Visual Area 3 $<0.001$ ADHD 0.955
Medial IntraParietal Area VentroMedial Visual Area 3 $<0.001$ ADHD 0.954
Area 10r ParaHippocampal Area 3 $<0.001$ ADHD 0.953
Orbital Frontal Complex Area TF $<0.001$ ADHD 0.948
Frontal Eye Fields Posterior InferoTemporal complex $<0.001$ ADHD 0.948
Ventral IntraParietal Complex VentroMedial Visual Area 3 $<0.001$ ADHD 0.945
Area a24 Area TF $<0.001$ ADHD 0.944
Area IFSp VentroMedial Visual Area 3 $<0.001$ ADHD 0.939
Area 6 anterior VentroMedial Visual Area 3 $<0.001$ ADHD 0.938
Frontal Eye Fields Ventral Visual Complex $<0.001$ ADHD 0.934
Area 8BM Area TF $<0.001$ ADHD 0.934
Area IntraParietal 1 Area TF $<0.001$ ADHD 0.934
Area s32 Area TF $<0.001$ ADHD 0.934
Posterior InferoTemporal complex Orbital Frontal Complex $<0.001$ ADHD 0.933
Area V3B VentroMedial Visual Area 3 $<0.001$ ADHD 0.932
IntraParietal Sulcus Area 1 VentroMedial Visual Area 3 $<0.001$ ADHD 0.932
Area TF Area V3B $<0.001$ ADHD 0.932
Area IFJa Posterior InferoTemporal complex $<0.001$ ADHD 0.93
IntraParietal Sulcus Area 1 Ventral Visual Complex $<0.001$ ADHD 0.927
Area a24 ParaHippocampal Area 3 $<0.001$ ADHD 0.927
Area 6 anterior Ventral Visual Complex $<0.001$ ADHD 0.926
Area 46 Eighth Visual Area $<0.001$ ADHD 0.925
Area Lateral IntraParietal ventral Fourth Visual Area $<0.001$ ADHD 0.923
Inferior 6-8 Transitional Area ParaHippocampal Area 3 $<0.001$ ADHD 0.923
Area IFJp VentroMedial Visual Area 3 $<0.001$ ADHD 0.921
Lateral Area 7P Ventral Visual Complex $<0.001$ ADHD 0.92
Inferior 6-8 Transitional Area VentroMedial Visual Area 2 $<0.001$ ADHD 0.92
Area dorsal 32 Area TF $<0.001$ ADHD 0.92
Area TF Area a24 $<0.001$ ADHD 0.919
Area IntraParietal 1 VentroMedial Visual Area 3 $<0.001$ ADHD 0.918
Posterior OFC Complex Area TF $<0.001$ ADHD 0.917
Inferior 6-8 Transitional Area Eighth Visual Area $<0.001$ ADHD 0.917
Area 10r Posterior InferoTemporal complex $<0.001$ ADHD 0.916
Area IFSp Posterior InferoTemporal complex $<0.001$ ADHD 0.916
Area 6 anterior Area Lateral Occipital 2 $<0.001$ ADHD 0.916
Anterior 24 prime VentroMedial Visual Area 3 $<0.001$ ADHD 0.913
Area IFJa Area TF $<0.001$ ADHD 0.912
Area s32 Posterior InferoTemporal complex $<0.001$ ADHD 0.911
Area IntraParietal 0 ParaHippocampal Area 3 $<0.001$ ADHD 0.91
Area 55b VentroMedial Visual Area 3 $<0.001$ ADHD 0.907
Inferior 6-8 Transitional Area Fourth Visual Area $<0.001$ ADHD 0.907
Anterior Ventral Insular Area Posterior InferoTemporal complex $<0.001$ ADHD 0.906
Area Lateral IntraParietal ventral Eighth Visual Area $<0.001$ ADHD 0.906
Area p32 prime VentroMedial Visual Area 3 $<0.001$ ADHD 0.905
Medial IntraParietal Area Ventral Visual Complex $<0.001$ ADHD 0.904
Area Lateral IntraParietal ventral VentroMedial Visual Area 1 $<0.001$ ADHD 0.904
Area p32 Area TF $<0.001$ ADHD 0.903
Area TA2 Eighth Visual Area $<0.001$ ADHD 0.902
Area 8Ad Area TF $<0.001$ ADHD 0.902
Area Lateral IntraParietal ventral VentroMedial Visual Area 2 $<0.001$ ADHD 0.902
Area 6 anterior Eighth Visual Area $<0.001$ ADHD 0.901
Premotor Eye Field Eighth Visual Area $<0.001$ ADHD 0.899
Area STSv anterior Area TF $<0.001$ ADHD 0.899
Medial Area 7A Ventral Visual Complex $<0.001$ ADHD 0.897
Posterior OFC Complex Posterior InferoTemporal complex $<0.001$ ADHD 0.896
Area 55b Posterior InferoTemporal complex $<0.001$ ADHD 0.895
Area 25 Posterior InferoTemporal complex $<0.001$ ADHD 0.894
Area PF Opercular VentroMedial Visual Area 3 $<0.001$ ADHD 0.893
Area IFJp VentroMedial Visual Area 1 $<0.001$ ADHD 0.891
Ventral IntraParietal Complex Ventral Visual Complex $<0.001$ ADHD 0.891
Anterior Agranular Insula Complex VentroMedial Visual Area 3 $<0.001$ ADHD 0.891
Area 55b Ventral Visual Complex $<0.001$ ADHD 0.89
Entorhinal Cortex ParaHippocampal Area 3 $<0.001$ ADHD 0.889
Area 46 VentroMedial Visual Area 3 $<0.001$ ADHD 0.889
Area STSd anterior Area TF $<0.001$ ADHD 0.887
Area TF Area IntraParietal 2 $<0.001$ ADHD 0.887
Area 6 anterior ParaHippocampal Area 3 $<0.001$ ADHD 0.887
Area IFJp Posterior InferoTemporal complex $<0.001$ ADHD 0.886
Frontal Opercular Area 4 VentroMedial Visual Area 3 $<0.001$ ADHD 0.886
Area IFSp Ventral Visual Complex $<0.001$ ADHD 0.884
Middle Insular Area VentroMedial Visual Area 3 $<0.001$ ADHD 0.883
Medial IntraParietal Area Area TF $<0.001$ ADHD 0.881
Ventral Visual Complex Frontal Eye Fields $<0.001$ ADHD 0.881
Area 7m Area TF $<0.001$ ADHD 0.879
Area 6m anterior Eighth Visual Area $<0.001$ ADHD 0.879
Area Lateral IntraParietal dorsal Ventral Visual Complex $<0.001$ ADHD 0.879
Area TA2 Posterior InferoTemporal complex $<0.001$ ADHD 0.878
Supplementary Table 20: Top connectivity features in high frequency range of ABIDE ASD prediction
Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
Dorsal Area 24d Area PGs $<0.001$ ASD 6.950
Area p32 Area PGs $<0.001$ ASD 6.565
Ventral Area 6 Area 8 Anterior ventral $<0.001$ ASD 6.479
Area p32 Area 8 Anterior ventral $<0.001$ ASD 6.313
Area 9 Middle Area 8 Anterior ventral $<0.001$ ASD 6.135
Area 9-46 dorsal Area PGs $<0.001$ ASD 6.057
Area ventral 23 a+b Area 8 Anterior ventral $<0.001$ ASD 6.041
Supplementary and Cingulate Eye Field Area PGs $<0.001$ ASD 5.959
Area 31pd Area 8 Anterior ventral $<0.001$ ASD 5.899
ParaHippocampal Area 3 Area 8Ad $<0.001$ ASD 5.819
Area 31 anterior Area PGs $<0.001$ ASD 5.671
Orbital Frontal Complex Area 47 medial $<0.001$ ASD 5.531
Area a24 Area 8 Anterior ventral $<0.001$ ASD 5.509
PreSubiculum Area PGs $<0.001$ ASD 5.178
Area 31p ventral Area 8 Anterior ventral $<0.001$ ASD 5.171
Area 23 dorsal Area PGs $<0.001$ ASD 5.107
Area V3A Area 47 medial $<0.001$ ASD 4.942
Area STSd anterior Area PGs $<0.001$ ASD 4.881
Area TemporoParietoOccipital Junction 3 Area 8Ad $<0.001$ ASD 4.745
Area 46 Area 47 medial $0.001$ ASD 4.683
Area IFSp Area 47 medial $0.001$ ASD 4.669
Area 10 rostral Area 8 Anterior ventral $0.001$ ASD 4.589
Area posterior 24 Area TE1 Middle $0.001$ ASD 4.574
Area posterior 24 Area 8 Anterior ventral $0.001$ ASD 4.550
Primary Visual Cortex Area 47 medial $0.001$ ASD 4.533
Area Posterior 24 prime Area 47 medial $0.001$ ASD 4.512
Premotor Eye Field Area 47 medial $0.001$ ASD 4.507
Ventral Area 6 Area PGs $0.001$ ASD 4.502
Inferior 6-8 Transitional Area Area 47 medial $0.001$ ASD 4.456
PreSubiculum Area 8 Anterior ventral $<0.001$ ASD 4.430
Area 9-46 dorsal Area 8 Anterior ventral $<0.001$ ASD 4.377
VentroMedial Visual Area 1 Area 47 medial $0.001$ ASD 4.365
Area 33 prime Area 47 medial $0.001$ ASD 4.293
Second Visual Area Area 47 medial $0.001$ ASD 4.274
Para-Insular Area Area 8 Anterior ventral $0.001$ ASD 4.227
ParaHippocampal Area 3 Area 8 Anterior ventral $0.001$ ASD 4.222
Area PFm Complex Area 8 Anterior ventral $0.001$ ASD 4.222
Third Visual Area Area 47 medial $0.001$ ASD 4.206
Area Lateral Occipital 3 Area 47 medial $0.001$ ASD 4.196
Area PGi Area 8 Anterior ventral $0.001$ ASD 4.193
Area p32 prime Area PGs $0.001$ ASD 4.178
Area 33 prime Area 8 Anterior ventral $0.001$ ASD 4.176
Superior Frontal Language Area Area 8 Anterior ventral $<0.001$ ASD 4.176
Supplementary and Cingulate Eye Field Area 8 Anterior ventral $0.001$ ASD 4.166
Middle Insular Area Area PGs $0.001$ ASD 4.161
Area dorsal 32 Area 8 Anterior ventral $<0.001$ ASD 4.079
Area anterior 32 prime Area 8 Anterior ventral $0.001$ ASD 4.073
Area posterior 24 Area 47 medial $0.001$ ASD 4.065
Area Lateral IntraParietal ventral Area 47 medial $0.002$ ASD 4.046
Area IFJp Area 47 medial $0.001$ ASD 4.034
Anterior 24 prime Area 47 medial $0.002$ ASD 4.026
Inferior 6-8 Transitional Area Area TE1 Middle $0.002$ ASD 4.011
Area 11 lateral Area 47 medial $0.002$ ASD 3.987
ParaHippocampal Area 3 Area 47l (47 lateral) $0.001$ ASD 3.980
Area IntraParietal 1 Area 8 Anterior ventral $0.002$ ASD 3.980
Superior 6-8 Transitional Area Area TE1 Middle $0.001$ ASD 3.968
Area PH Area 47 medial $<0.001$ ASD 3.963
Area Lateral Occipital 2 Area 8 Anterior ventral $0.001$ ASD 3.956
Eighth Visual Area Inferior 6-8 Transitional Area $0.002$ ASD 3.952
Area 6 anterior Area 47 medial $0.002$ ASD 3.948
Para-Insular Area Area 47 medial $0.002$ ASD 3.918
Area s32 Area PGs $0.002$ ASD 3.913
Area OP1-SII Inferior 6-8 Transitional Area $0.002$ ASD 3.897
Area a24 Area PGs $0.002$ ASD 3.895
Area TE1 Middle Area 8 Anterior ventral $0.001$ ASD 3.895
Area Lateral Occipital 2 Area 47 medial $0.002$ ASD 3.875
Area 6 medial anterior Area TE1 Middle $0.002$ ASD 3.866
Area dorsal 23 a+b Area 8 Anterior ventral $0.002$ ASD 3.858
Supplementary and Cingulate Eye Field Area TE1 Middle $0.002$ ASD 3.843
Area TemporoParietoOccipital Junction 2 Area 8 Anterior ventral $0.002$ ASD 3.840
Area posterior 9-46 ventral Area 47 medial $<0.001$ ASD 3.825
Pirform Cortex Area 47 medial $0.002$ ASD 3.804
Area Posterior Insular 1 Area 47 medial $0.002$ ASD 3.800
Fourth Visual Area Area 47 medial $0.002$ ASD 3.798
Area a24 Dorsal area 6 $0.002$ ASD 3.789
Area anterior 10p Area TE1 Middle $0.002$ ASD 3.788
Anterior 24 prime Area TE1 Middle $0.002$ ASD 3.785
Visual area 3-B Area 47 medial $0.002$ ASD 3.757
Area s32 Dorsal area 6 $0.002$ ASD 3.752
Area TemporoParietoOccipital Junction 3 Area 8 Anterior ventral $0.002$ ASD 3.744
RetroInsular Cortex Area 8 Anterior ventral $0.002$ ASD 3.736
Area 46 Area 8 Anterior ventral $<0.001$ ASD 3.731
Perirhinal Ectorhinal Cortex Area 47 medial $0.002$ ASD 3.731
Area p32 prime Area 47 medial $0.002$ ASD 3.720
Lateral Belt Complex Area 47 medial $0.002$ ASD 3.702
Anterior 24 prime Area 8 Anterior ventral $0.002$ ASD 3.698
Area OP1-SII Primary Motor Cortex $0.002$ ASD 3.694
Area 6 medial anterior Superior 6-8 Transitional Area $0.002$ ASD 3.693
Area Lateral IntraParietal dorsal Medial Area 7P (posterior) $0.002$ ASD 3.693
Area p32 prime Area 8 Anterior ventral $0.002$ ASD 3.682
Premotor Eye Field Area 8 Anterior ventral $0.002$ ASD 3.678
Entorhinal Cortex Area 47 medial $0.003$ ASD 3.654
Area 31 anterior Area 8 Anterior ventral $0.002$ ASD 3.629
Area 7m Area 8 Anterior ventral $0.002$ ASD 3.625
Area dorsal 32 Area TG dorsal $0.003$ ASD 3.608
Area Posterior 24 prime Area 8 Anterior ventral $0.002$ ASD 3.602
Area IntraParietal 1 Medial Area 7P (posterior) $0.002$ ASD 3.593
Area dorsal 23 a+b Inferior 6-8 Transitional Area $0.003$ ASD 3.579
Ventral Area 6 Area 47 medial $0.003$ ASD 3.574
Area 31pd Area PGs $0.001$ ASD 3.556
Supplementary Table 21: Top connectivity features in low frequency range of ABIDE ASD prediction
Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
Area PHT Area posterior 9-46 ventral $0.028$ ASD 2.155
ParaBelt Complex Area posterior 9-46 ventral $0.029$ ASD 2.154
Auditory 5 Complex Area posterior 9-46 ventral $0.030$ ASD 2.123
Area 9 Posterior Area posterior 9-46 ventral $0.030$ ASD 2.122
Primary Sensory Cortex Area posterior 9-46 ventral $0.031$ ASD 2.116
Area PH Area posterior 9-46 ventral $0.031$ ASD 2.115
Area 3a Area posterior 9-46 ventral $0.031$ ASD 2.109
Area 6mp Area posterior 9-46 ventral $0.032$ ASD 2.099
Lateral Belt Complex Area posterior 9-46 ventral $0.033$ ASD 2.078
Area TA2 Area posterior 9-46 ventral $0.034$ ASD 2.065
Area TE2 posterior Area posterior 9-46 ventral $0.034$ ASD 2.057
Area V3A Area posterior 9-46 ventral $0.034$ ASD 2.057
Area PFm Complex Area posterior 9-46 ventral $0.034$ ASD 2.053
Area Lateral Occipital 2 Area posterior 9-46 ventral $0.035$ ASD 2.046
Dorsal area 6 Area posterior 9-46 ventral $0.035$ ASD 2.045
VentroMedial Visual Area 1 Area posterior 9-46 ventral $0.035$ ASD 2.039
ParaHippocampal Area 3 Area posterior 9-46 ventral $0.036$ ASD 2.038
Area IFSp Area posterior 9-46 ventral $0.036$ ASD 2.037
Primary Auditory Cortex Area posterior 9-46 ventral $0.036$ ASD 2.032
Ventral Visual Complex Area posterior 9-46 ventral $0.037$ ASD 2.023
Area 1 Area posterior 9-46 ventral $0.036$ ASD 2.018
Area PGs Area posterior 9-46 ventral $0.037$ ASD 2.009
Area IntraParietal 2 Area posterior 9-46 ventral $0.037$ ASD 2.007
Area TemporoParietoOccipital Junction 1 Area posterior 9-46 ventral $0.037$ ASD 2.004
Primary Motor Cortex Area posterior 9-46 ventral $0.038$ ASD 2.001
ParaHippocampal Area 1 Area posterior 9-46 ventral $0.039$ ASD 1.999
Auditory 4 Complex Area posterior 9-46 ventral $0.039$ ASD 1.993
Area 31a Area posterior 9-46 ventral $0.039$ ASD 1.987
Area TE1 posterior Area posterior 9-46 ventral $0.039$ ASD 1.985
Superior Temporal Visual Area Area posterior 9-46 ventral $0.039$ ASD 1.980
Area 25 Area posterior 9-46 ventral $0.040$ ASD 1.978
Area V6A Area posterior 9-46 ventral $0.039$ ASD 1.977
Area OP2-3-VS Area posterior 9-46 ventral $0.040$ ASD 1.975
Ventral Area 24d Area posterior 9-46 ventral $0.041$ ASD 1.975
Area 2 Area posterior 9-46 ventral $0.040$ ASD 1.963
ProStriate Area Area posterior 9-46 ventral $0.041$ ASD 1.960
RetroInsular Cortex Area posterior 9-46 ventral $0.041$ ASD 1.960
Area 9 anterior Area posterior 9-46 ventral $0.041$ ASD 1.954
VentroMedial Visual Area 3 Area posterior 9-46 ventral $0.042$ ASD 1.953
ParaHippocampal Area 2 Area posterior 9-46 ventral $0.042$ ASD 1.947
Area anterior 9-46 ventral Area posterior 9-46 ventral $0.043$ ASD 1.941
Medial Belt Complex Area posterior 9-46 ventral $0.043$ ASD 1.940
Insular Granular Complex Area posterior 9-46 ventral $0.043$ ASD 1.936
Eighth Visual Area Area posterior 9-46 ventral $0.044$ ASD 1.933
Superior Frontal Language Area Area posterior 9-46 ventral $0.044$ ASD 1.931
Area TemporoParietoOccipital Junction 3 Area posterior 9-46 ventral $0.044$ ASD 1.921
Area 43 Area posterior 9-46 ventral $0.044$ ASD 1.920
Posterior InferoTemporal complex Area posterior 9-46 ventral $0.045$ ASD 1.915
Area 7PC Area posterior 9-46 ventral $0.045$ ASD 1.913
Area STSv posterior Area posterior 9-46 ventral $0.044$ ASD 1.912
Area V3B Area posterior 9-46 ventral $0.046$ ASD 1.908
Area Lateral IntraParietal dorsal Area posterior 9-46 ventral $0.045$ ASD 1.905
Area Lateral Occipital 1 Area posterior 9-46 ventral $0.046$ ASD 1.904
Area Posterior Insular 1 Area posterior 9-46 ventral $0.046$ ASD 1.900
Dorsal Transitional Visual Area Area posterior 9-46 ventral $0.046$ ASD 1.900
Fusiform Face Complex Area posterior 9-46 ventral $0.047$ ASD 1.890
Second Visual Area Area posterior 9-46 ventral $0.048$ ASD 1.887
PeriSylvian Language Area Area posterior 9-46 ventral $0.047$ ASD 1.887
Area s32 Area posterior 9-46 ventral $0.047$ ASD 1.885
Area 46 Area posterior 9-46 ventral $0.047$ ASD 1.884
Hippocampus Area posterior 9-46 ventral $0.048$ ASD 1.880
VentroMedial Visual Area 2 Area posterior 9-46 ventral $0.048$ ASD 1.879
Area 8Av Area posterior 9-46 ventral $0.048$ ASD 1.877
Area OP4-PV Area posterior 9-46 ventral $0.048$ ASD 1.877
Posterior Insular Area 2 Area posterior 9-46 ventral $0.048$ ASD 1.876
Area TemporoParietoOccipital Junction 2 Area posterior 9-46 ventral $0.047$ ASD 1.876
Area 55b Area posterior 9-46 ventral $0.048$ ASD 1.876
Area 23d Area posterior 9-46 ventral $0.048$ ASD 1.874
PreSubiculum Area posterior 9-46 ventral $0.048$ ASD 1.871
Anterior IntraParietal Area Area posterior 9-46 ventral $0.048$ ASD 1.870
Area PFcm Area posterior 9-46 ventral $0.049$ ASD 1.869
Ventral Area 6 Area posterior 9-46 ventral $0.049$ ASD 1.867
Area 33 prime Area posterior 9-46 ventral $0.049$ ASD 1.866
Area STSd posterior Area posterior 9-46 ventral $0.048$ ASD 1.865
Parieto-Occipital Sulcus Area 1 Area posterior 9-46 ventral $0.049$ ASD 1.862
Seventh Visual Area Area posterior 9-46 ventral $0.050$ ASD 1.860
Area OP1-SII Area posterior 9-46 ventral $0.049$ ASD 1.860
Area STSd posterior Area Lateral IntraParietal dorsal $0.049$ ASD 1.838
Supplementary Table 22: Top connectivity features in ultralow frequency range of ABIDE ASD prediction
Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
Medial Belt Complex Area 7PC $0.002$ HC 3.728
Area TA2 (temporal region A, area 2) Area 7PC $0.005$ HC 3.157
ParaBelt Complex Area Frontal Opercular 5 $0.005$ HC 3.119
Dorsal area 6 Area Frontal Opercular 5 $0.009$ HC 2.781
ParaHippocampal Area 1 Area 7PC $0.009$ HC 2.752
Auditory 4 Complex Area 7PC $0.011$ HC 2.739
ParaHippocampal Area 1 Area posterior 9-46 ventral $0.010$ HC 2.729
Auditory 4 Complex Area Frontal Opercular 5 $0.011$ HC 2.696
Area PH Area Frontal Opercular 5 $0.004$ HC 2.570
Superior 6-8 Transitional Area Area 7PC $0.013$ HC 2.551
Posterior Insular Area 2 Area Frontal Opercular 5 $0.014$ HC 2.546
Medial Belt Complex Area Lateral IntraParietal dorsal $0.005$ HC 2.541
ParaHippocampal Area 1 Area IFJa $0.014$ HC 2.520
Medial Belt Complex Area posterior 9-46 ventral $0.014$ HC 2.518
Area anterior 9-46 ventral Area 7PC $0.014$ HC 2.515
Area OP2-3-VS (Ventral-Superior part of 2-3 area of Parietal Operculum) Area Frontal Opercular 5 $0.014$ HC 2.503
Medial Belt Complex Area IFJa $0.017$ HC 2.444
Area 8B Lateral Area 7PC $0.006$ HC 2.440
Area 6 medial anterior Area Lateral IntraParietal dorsal $0.005$ HC 2.436
ParaBelt Complex Area 7PC $0.017$ HC 2.433
Middle Temporal Area Area 7PC $0.017$ HC 2.402
Area 52 Area Lateral IntraParietal dorsal $0.006$ HC 2.398
Area 46 Area Frontal Opercular 5 $0.018$ HC 2.394
Medial Belt Complex Frontal Opercular Area 2 $0.018$ HC 2.379
Dorsal area 6 Area Lateral IntraParietal dorsal $0.006$ HC 2.374
Area 6 medial posterior Area Lateral IntraParietal dorsal $0.019$ HC 2.364
Posterior Insular Area 2 Area 7PC $0.019$ HC 2.357
Area OP4-PV (Posterior-Ventral part of area 4 of Parietal Operculum) Area Frontal Opercular 5 $0.020$ HC 2.335
Inferior 6-8 Transitional Area Area Frontal Opercular 5 $0.021$ HC 2.328
Frontal Opercular Area 1 Area Lateral IntraParietal dorsal $0.006$ HC 2.319
Area posterior 10p Area 7PC $0.021$ HC 2.310
Primary Auditory Cortex Area 7PC $0.023$ HC 2.308
Dorsal area 6 Area 7PC $0.021$ HC 2.305
Anterior Agranular Insula Complex Area Frontal Opercular 5 $0.023$ HC 2.290
ParaHippocampal Area 1 Area Lateral IntraParietal dorsal $0.007$ HC 2.287
Auditory 4 Complex Area IFJa $0.022$ HC 2.276
Auditory 4 Complex Area posterior 9-46 ventral $0.024$ HC 2.267
Area IFJa Area Frontal Opercular 5 $0.022$ HC 2.265
Area 9 Middle Area PGs $0.022$ HC 2.264
Area posterior 9-46 ventral Area Frontal Opercular 5 $0.024$ HC 2.256
Area anterior 9-46 ventral Area Frontal Opercular 5 $0.023$ HC 2.255
Area PF Opercular Area Frontal Opercular 5 $0.023$ HC 2.254
Frontal Opercular Area 1 Area 7PC $0.023$ HC 2.248
PreSubiculum Area PGs $0.023$ HC 2.247
Superior 6-8 Transitional Area Area Lateral IntraParietal dorsal $0.024$ HC 2.241
Area TA2 (temporal region A, area 2) Area Lateral IntraParietal dorsal $0.023$ HC 2.241
Area OP1-SII Area Frontal Opercular 5 $0.024$ HC 2.229
Area s32 Area 7PC $0.024$ HC 2.228
Medial Belt Complex Area Frontal Opercular 5 $0.008$ HC 2.226
Area PH Area IFJa $0.024$ HC 2.224
Area Posterior Insular 1 Area Lateral IntraParietal dorsal $0.024$ HC 2.215
VentroMedial Visual Area 1 Area Frontal Opercular 5 $0.025$ HC 2.213
Area 33 prime Area Frontal Opercular 5 $0.025$ HC 2.211
Area 43 Area IFJa $0.026$ HC 2.209
Area OP2-3-VS (Ventral-Superior part of 2-3 area of Parietal Operculum) Area Lateral IntraParietal dorsal $0.025$ HC 2.206
Frontal Opercular Area 1 Area Frontal Opercular 5 $0.025$ HC 2.205
Pirform Cortex Area Lateral IntraParietal dorsal $0.009$ HC 2.187
Area 9 anterior Area PGs $0.026$ HC 2.183
Area 8B Lateral Area PGs $0.026$ HC 2.183
Area PF Complex Area Frontal Opercular 5 $0.027$ HC 2.179
Anterior Agranular Insula Complex Area posterior 9-46 ventral $0.027$ HC 2.176
Area 1 Area Frontal Opercular 5 $0.027$ HC 2.173
Premotor Eye Field Area Frontal Opercular 5 $0.027$ HC 2.173
ParaHippocampal Area 1 Area Frontal Opercular 5 $0.027$ HC 2.172
Area PFt Area Frontal Opercular 5 $0.027$ HC 2.169
Area 13 lateral Area Frontal Opercular 5 $0.027$ HC 2.163
Rostral Area 6 Area 7PC $0.028$ HC 2.159
Anterior Agranular Insula Complex Area 7PC $0.029$ HC 2.143
Area 6 medial posterior Area PGs $0.028$ HC 2.139
Area 43 Area Frontal Opercular 5 $0.028$ HC 2.138
Area OP4-PV (Posterior-Ventral part of area 4 of Parietal Operculum) Area Lateral IntraParietal dorsal $0.028$ HC 2.138
Posterior Insular Area 2 Area posterior 9-46 ventral $0.029$ HC 2.128
ParaHippocampal Area 1 Sixth Visual Area $0.030$ HC 2.114
Parieto-Occipital Sulcus Area 1 Area Lateral IntraParietal dorsal $0.009$ HC 2.111
Area posterior 10p Area Frontal Opercular 5 $0.031$ HC 2.108
Ventral Area 6 Area Frontal Opercular 5 $0.030$ HC 2.106
Area STSd anterior Area Lateral IntraParietal dorsal $0.009$ HC 2.100
Area anterior 9-46 ventral Area Lateral IntraParietal dorsal $0.031$ HC 2.093
Area Posterior Insular 1 Area PGs $0.031$ HC 2.093
Rostral Area 6 Area Frontal Opercular 5 $0.010$ HC 2.091
Area FST Area Frontal Opercular 5 $0.031$ HC 2.085
Area Superior temporal sulcus ventral anterior Area 7PC $0.031$ HC 2.081
Area Posterior Insular 1 Area posterior 9-46 ventral $0.032$ HC 2.076
Area TE1 Middle Area 7PC $0.032$ HC 2.070
Area 6 medial posterior Area 7PC $0.032$ HC 2.069
Area Posterior Insular 1 Area 7PC $0.032$ HC 2.067
Area Posterior Insular 1 Area Frontal Opercular 5 $0.033$ HC 2.063
Area TA2 (temporal region A, area 2) Area PGs $0.032$ HC 2.062
PreSubiculum Area Lateral IntraParietal dorsal $0.011$ HC 2.055
Area 43 Area 7PC $0.033$ HC 2.048
ParaHippocampal Area 2 Area 7PC $0.034$ HC 2.045
Inferior 6-8 Transitional Area Area Lateral IntraParietal dorsal $0.034$ HC 2.044
ParaHippocampal Area 3 Area Lateral IntraParietal dorsal $0.012$ HC 2.043
Area 10 rostral Area PGs $0.034$ HC 2.041
ParaBelt Complex Area Lateral IntraParietal dorsal $0.034$ HC 2.041
Medial Superior Temporal Area Area Frontal Opercular 5 $0.034$ HC 2.041
Superior 6-8 Transitional Area Area PGs $0.034$ HC 2.040
Entorhinal Cortex Area Lateral IntraParietal dorsal $0.034$ HC 2.037
VentroMedial Visual Area 1 Area IFJa $0.034$ HC 2.036
Area Lateral Occipital 3 Area Lateral IntraParietal dorsal $0.012$ HC 2.033

### v2

@@
 Abstract
-A central challenge in neuroscience is understanding how the brain’s complex nonlinear dynamics give rise to emergent properties underlying adaptive cognition and behavior across spatial and temporal scales.
-These dynamics exhibit scale-free and multifractal properties, shaping the adaptive reconfiguration of neural networks.
-However, conventional neuroimaging models remain constrained by assumptions of linearity and stationarity, limiting their capacity to fully characterize these intricate processes.
-Transformer-based architectures, known for capturing long-range dependencies in sequential data, naturally align with the brain’s intrinsically hierarchical and temporal organization.
-Here, we introduce Multi-Band Brain Net (MBBN), a transformer-based deep learning framework that explicitly models frequency-specific spatiotemporal brain dynamics from fMRI by integrating scale-free network principles with frequency-resolved multi-band self-attention mechanisms.
-Trained on three large-scale neuroimaging cohorts (UK Biobank, ABCD, ABIDE) totaling 45,951 individuals, MBBN reveals previously undetectable frequency-dependent intrinsic network interactions, shedding new light on distinct connectivity disruptions across psychiatric conditions (ADHD, ASD, depression).
-This extensive validation demonstrates robust generalizability and highlights core neural principles conserved across populations and developmental stages.
-Specifically, MBBN achieves up to 30.59% higher predictive accuracy than state-of-the-art deep learning and conventional connectome-based methods, demonstrating the advantage of frequency-informed spatiotemporal modeling in capturing latent neural computations.
-Furthermore, MBBN’s attention-based interpretability reveals novel frequency-specific biomarkers for neurodevelopmental disorders, offering mechanistic insights into the hierarchical organization of brain function.
-By providing a principled, interpretable framework for dissecting spatiotemporal learning mechanisms, MBBN advances methodological rigor and provides new insight into how dynamic neural computations underpin cognitive function and psychiatric vulnerability, with implications for brain decoding, cognitive neuroscience, and precision psychiatry.
+Understanding how the brain’s complex nonlinear dynamics give rise to cognitive function remains a central challenge in neuroscience.
+While brain functional dynamics exhibits scale-free and multifractal properties across temporal scales, conventional neuroimaging analytics assume linearity and stationarity, failing to capture frequency-specific neural computations.
+Here, we introduce Multi-Band Brain Net (MBBN), the first transformer-based framework to explicitly model frequency-specific spatiotemporal brain dynamics from fMRI.
+MBBN integrates biologically-grounded frequency decomposition with multi-band self-attention mechanisms, enabling discovery of previously undetectable frequency-dependent network interactions.
+Trained on 49,673 individuals across three large-scale cohorts (UK Biobank, ABCD, ABIDE), MBBN sets a new state-of-the-art in predicting psychiatric and cognitive outcomes (depression, ADHD, ASD), showing particular strength in classification tasks with up to 52.5% higher AUROC and provides a novel framework for predicting cognitive intelligence scores.
+Frequency-resolved analyses uncover disorder-specific signatures: in ADHD, high-frequency fronto-sensorimotor connectivity is attenuated and opercular somatosensory nodes emerge as dynamic hubs; in ASD, orbitofrontal-somatosensory circuits show focal high-frequency disruption together with enhanced ultra-low-frequency coupling between the temporo-parietal junction and prefrontal cortex.
+By integrating scale-aware neural dynamics with deep learning, MBBN delivers more accurate and interpretable biomarkers, opening avenues for precision psychiatry and developmental neuroscience.
 keywords:
@@
 Understanding the brain’s dynamic processes is essential for unraveling the mechanisms underlying adaptive cognition and behavior.
-These processes influence critical functions such as learning, decision-making, and social interactions, and their dysregulation is closely linked to mental health conditions ([1, 2, 3, 4]).
-While functional magnetic resonance imaging (fMRI) has been instrumental in probing brain activity, many conventional analytical methods remain constrained by assumptions of linearity, stationary, static connectivity.
+These processes influence critical functions such as learning, decision-making, and social interactions, with mounting evidence that their dysregulation is closely linked to mental health conditions ([1, 2, 3, 4]).
+While functional magnetic resonance imaging (fMRI) has been instrumental in probing brain activity, many conventional analytical methods remain constrained by assumptions of linearity, stationarity and static connectivity.
 Approaches ranging from mass-univariate statistical tests and seed-based correlations to static functional connectivity maps often treat the brain as a relatively fixed system, thus failing to capture the temporal fluctuations, nonlinear interactions, and adaptive reconfigurations that characterize real neural dynamics.
-Recent advances in artificial intelligence (AI), particularly transformer-based architectures, have begun to address these shortcomings by modeling brain activity as a continuously evolving network of interdependent processes ([5, 6, 7, 8]).
-However, a critical gap remains in understanding and modeling two fundamental properties of neural dynamics: scale-free dynamics and multifractal structures.
-These properties are essential for comprehending the brain’s hierarchical complexity and adaptability ([9, 10, 2]).
-Scale-free dynamics, characterized by power-law relationships (Equation 1), reveal how neural activity maintains consistent patterns across different timescales and modalities.
+Recent advances in artificial intelligence, particularly transformer-based architectures originally designed for sequential data, have shown promise for modeling brain activity as a continuously evolving network of interdependent processes.
+Yet, these approaches largely ignore the frequency-specific nature of neural oscillations ([5, 6, 7, 8]).
+This oversight is particularly problematic given that existing neuroimaging models, including recent deep learning approaches, treat brain signals as broadband phenomena, overlooking the fundamental frequency-specific organization of neural computation.
+Indeed, this organization is best understood through two fundamental properties that describe the brain’s complexity: its multifractal structure and its scale-free dynamics.
+Multifractality, a key structural property, points to the brain’s nested, hierarchical organization where multiple scaling dimensions and power-law behaviors coexist ([9, 10, 11, 12, 13]).
+Crucially, it is within this complex structure that scale-free dynamics—a key functional property—become apparent.
+While a broadband signal is not inherently scale-free, decomposing it into distinct frequency components reveals power-law relationships (Equation 1) that signify consistent activity patterns across different timescales.
 These dynamics play crucial roles in brain maturation and state transitions, with temporal variability serving as a key regulatory mechanism ([14, 15]).
 The power-law exponent ($\beta$) varies across brain regions and states, providing valuable insights into neural network heterogeneity and serving as a potential marker for neurological and psychiatric disorders ([16, 17]).
-Multifractality extends this framework by incorporating multiple scaling dimensions and power-law behaviors ([15, 16, 17]), offering a sophisticated approach to understanding the brain’s nested, hierarchical organization ([18, 19]).
-This property manifests in distinct frequency-dependent patterns that reflect neural adaptability and functional specialization ([20]).
-Evidence from fMRI, ECoG, and EEG studies suggests that these multifractal principles drive frequency-specific changes in connectivity and topological organization, particularly in properties such as small-worldness and modularity ([21, 22]).
+Evidence from fMRI, ECoG, and EEG studies further suggests that these principles drive frequency-specific changes in connectivity and topological organization, such as small-worldness and modularity ([18, 19]).
 Understanding these patterns is crucial for developing more effective diagnostic and therapeutic strategies.
-To address these challenges, we introduce the Multi-Band Brain Net (MBBN), a novel framework based on Bidirectional Encoder Representations from Transformers (BERT) that integrates scale-free property-based frequency decomposition with advanced self-attention mechanisms and network theory concepts like communicability (Figure 1).
+To address these challenges, we introduce Multi-Band Brain Net (MBBN), the first transformer-based framework specifically designed to model frequency-resolved spatiotemporal brain dynamics from fMRI.
+Unlike existing approaches that analyze broadband signals, MBBN explicitly decomposes neural activity into biologically meaningful frequency bands using scale-free principles, then applies specialized attention mechanisms to capture frequency-specific connectivity patterns incorporating network theory concepts like communicability (Figure 1, 4).
 MBBN divides fMRI BOLD signals into distinct frequency bands, enabling the capture of band-specific functional and topological properties.
@@
 C.
-The frequency decomposition process is demonstrated using a representative sample from the UKB dataset.
-Although the knee frequencies ($f_{1}$, $f_{2}$) may vary across individuals, the overall pattern of frequency division remains consistent.
-The blue line represents the original power spectral density (PSD) of the fMRI signal.
-The pink, yellow-green, and yellow lines show the fitted power-law functions for the high, ultra-low, and low-frequency ranges, respectively.
+Frequency decomposition using scale-free principles on representative UKB data.
+Individual-specific knee frequencies ($f_{1}$, $f_{2}$) separate three biologically distinct bands: ultralow (pink), low (yellow-green), and high (yellow) frequencies, each exhibiting characteristic power-law scaling.
 The dotted red and green lines indicate the frequency boundaries: ultra-low to low ($f_{1}$), and low to high ($f_{2}$) frequencies, respectively.
 2 Results
-2.1 Comparison of Model Performance Across Various Downstream Tasks
+2.1 MBBN demonstrates superior performance across diverse neuroimaging tasks
 Figure 2: Comparative Analysis of fMRI Prediction Models.
-Across datasets and tasks, the proposed MBBN demonstrates superior predictive performance compared to baseline models including state-of-the-art models in the field (e.g., Brain Network Transformer ([5])), as well as vanilla BERT (without multi-fractality division).
-[CC] indicates models based on Pearson’s correlation-based connectivity; [TD] represents, models accounting for temporal dynamics.
+Across datasets and tasks, the proposed MBBN demonstrates superior predictive performance compared to baseline models including state-of-the-art models in the field (e.g., Brain Network Transformer ([5])), as well as vanilla BERT (without multifractal decomposition).
+[CC] indicates models based on Pearson’s correlation-based connectivity; [TD] represents models accounting for temporal dynamics.
 The x-axis denotes the models, and the y-axis represents performance metrics, with AUROC used for classification tasks and MAE for regression tasks.
-(A–B) show results for sex prediction, (C–D) for fluid intelligence prediction, (E–F) for depression prediction, G for ADHD prediction, and H for ASD prediction.
-Specifically: (A, B, C, G) correspond to the ABCD dataset using the HCP-MMP1 atlas.
-(D, E, F) correspond to the UKB dataset using the HCP-MMP1 atlas.
-H correspond to the ABIDE dataset using the HCP-MMP1 atlas.
-To evaluate model performance, we trained MBBN from scratch on multiple downstream tasks spanning biological, cognitive, and clinical domains, using both the HCP-MMP1 and Schaefer atlases across the UKB and ABCD datasets.
-MBBN consistently outperformed baseline models across all tasks.
-For biological phenotype prediction (sex classification), MBBN achieved notable gains of up to 11.6% on UKB and 18.88% on ABCD using the HCP-MMP1 atlas, with similar improvements using the Schaefer atlas (9.06% on UKB, 14.78% on ABCD).
-In cognitive assessment tasks, fluid intelligence prediction showed substantial improvements on the ABCD dataset (up to 18.88% with HCP-MMP1, 8.03% with Schaefer).
-For clinical applications, depression prediction on the UKB dataset demonstrated meaningful gains (up to 9.42% with HCP-MMP1, 6.98% with Schaefer).
-Comprehensive performance comparisons, including task-specific metrics and statistical analyses, are detailed in Appendix.
-2.2 Communicability-based masking loss improves fine-tuning performance over random masking loss.
-A key component of our methodology was a pretraining loss function designed to mask nodes with high communicability within the brain network.
+MBBN consistently outperforms baseline models across biological (A-B: sex prediction), cognitive (C-D: fluid intelligence), and clinical (E-H: depression, ADHD, ASD) prediction tasks.
+[CC] indicates correlation-based connectivity models; [TD] represents temporal dynamics models.
+AUROC is shown for classification tasks, MAE for regression tasks.
+Data from ABCD (A-C, G), UKB (D-F), and ABIDE (H) datasets using HCP-MMP1 atlas.
+A-F denotes from-scratch experiments and G-H denotes fine-tuning experiments.
+Task MBBN vs Best Baseline (%) vs Worst Baseline (%)
+ABCD Sex 0.916 (AUROC) +2.346 +19.271
+ABCD Fluid Intelligence 0.717 (MAE) +0.693 +9.584
+ABCD Depression 0.556 (AUROC) +1.460 +7.961
+UKB Sex 0.980 (AUROC) +3.158 +14.086
+UKB Fluid Intelligence 1.630 (MAE) -3.099 -1.305
+UKB Depression 2.21 (MAE) -1.231 +0.924
+ABCD ADHD 0.645 (AUROC) +1.735 +7.860
+ABIDE ASD 0.810 (AUROC) +24.615 +52.542
+Table 1: Performance gain of MBBN between best and worst baseline model. We illustrated finetuned MBBN for comparison of ABCD ADHD classification and ABIDE ASD classification.
+To evaluate its performance, MBBN was trained from scratch on a diverse set of downstream tasks spanning biological, cognitive, and clinical domains, utilizing both the HCP-MMP1 and Schaefer atlases with the UKB and ABCD datasets.
+While MBBN demonstrated unparalleled performance in classification tasks, its advantage in regression tasks (e.g., fluid intelligence and depression score prediction) was less pronounced, suggesting that the model’s architecture is particularly optimized for distinguishing categorical phenotypes rather than predicting continuous scores.
+The primary results presented in Figure 2 A-F focus on experiments conducted with the HCP-MMP1 atlas; corresponding analyses using the Schaefer atlas are detailed in the Appendix E.1.
+For biological phenotype prediction, MBBN achieved substantial improvements in sex classification with gains up to 14.09% (UKB) and 19.27% (ABCD) using HCP-MMP1 atlas.
+Cognitive assessment showed marked improvements, with fluid intelligence prediction demonstrating up to 9.58% enhancement on ABCD.
+Also clinical prediction tasks revealed MBBN’s therapeutic potential, achieving 0.92% improvement in depression prediction (UKB) and exceptional performance in psychiatric disorders (detailed in Supplementary Materials).
+2.2 Communicability-based pretraining enables superior clinical prediction performance
+The exceptional performance of MBBN, particularly in clinical prediction tasks, stems from our novel communicability-based pretraining strategy that masks nodes with high communicability within the brain network.
 This strategy was predicated on the hypothesis that selectively reconstructing these highly influential nodes during pretraining would better represent the complex communication patterns within the brain.
@@
 These datasets and tasks were chosen for their established clinical significance and high prevalence rates.
-Fine-tuning experiments revealed a clear advantage for the MBBN-high model, which employed a pretraining strategy that selectively masked nodes with high communicability.
-This approach resulted in significant performance improvements compared to baselines: a 5.55% gain in ADHD classification accuracy and a remarkable 30.59% gain in ASD classification accuracy.
+Fine-tuning experiments (Figure 2 G-H) revealed a clear advantage for the MBBN-high model, which employed a pretraining strategy that selectively masked nodes with high communicability.
+This approach resulted in significant performance improvements compared to baselines: a 7.86% gain in ADHD classification AUROC and a 52.54% gain in ASD classification AUROC.
 In contrast, models using random node masking (MBBN-random) or masking nodes with low communicability (MBBN-low) showed inferior performance, even compared to the MBBN model trained from scratch (MBBN-from scratch).
 These findings underscore the critical role of communicability-based pretraining in effectively leveraging the underlying network structure of the data to enhance predictive performance.
-Complete results and statistical details are available in Appendix.
-2.3 Key connectivity for mental illness classification revealed by attention mechanisms
-Figure 3: Self-Attention-Based Connectivity Learned from Resting FMRI Distinguishing ADHD and ASD from Healthy Controls.
-For presentational purposes, this figure highlights statistically significant connections with the top 10 largest effect sizes that differentiate individuals with ADHD (A-C) or ASD (D-F) from healthy controls (HC) (FDR corrected P’s ¡ 0.05).
-[Red: Higher in the disorder group; Blue: Lower in the disorder group.].
-A detailed list of the top 100 connectivity is provided in the Appendix F.
-(A-C) show ADHD connectivity patterns compared to HC across frequency bands: high (A), low (B), and ultralow (C).
-(D-F) illustrate ASD patterns compared to HC: high (D), low (E), and ultralow (F).
-The HCP-MMP1 atlas was used.
-We utilized a Grad-CAM-based attribution score to identify self-attention-weighted-connectivity with large effect sizes distinguishing healthy controls from individuals with mental illness and analyzed their roles.
+Detailed results and statistical details are available in Appendix E.1.
+These results demonstrate that communicability-based loss better captures a complex information flow of the brain than conventional masking loss, and MBBN’s frequency-resolved approach captures clinically relevant neural signatures that are missed by conventional broadband analysis methods.
+2.3 Frequency-specific connectivity patterns reveal distinct psychiatric biomarkers
+Figure 3: Frequency-specific connectivity disruptions in ADHD and ASD.
+Statistically significant connections (FDR-corrected $p<0.05$, top 10 effect sizes) distinguishing ADHD (A-C) and ASD (D-F) from healthy controls across high, low, and ultralow frequency bands.
+Red: increased connectivity in disorder; Blue: decreased connectivity.
+Detailed results in Supplementary Materials.
+To identify frequency-specific biomarkers underlying psychiatric disorders, we applied Grad-CAM-based attribution analysis to extract self-attention-weighted connectivity patterns that distinguished patient groups from healthy controls.
+This approach revealed distinct frequency-dependent neural signatures for ADHD and ASD that were previously undetectable using conventional broadband connectivity analysis.
 Details are described in Section 3.9.
@@
 These findings reveal frequency-specific neural connectivity alterations in ADHD.
-The results of the t-test distinguishing between the HC and ADHD groups showed that 6.55% of the statistically significant connections ($p<0.05$) were in the high-frequency range, with the 75% quantile value of Cohen’s d score being 0.71.
-In contrast, the low-frequency range accounted for only 1.67%, with the 75% quantile value of Cohen’s d being 0.17.
-The ultralow frequency range accounted for 8.39%, with the 75% quantile value of Cohen’s d being 0.71.
-In the high-frequency range (Figure 3 A), distinct connectivity patterns emerged involving key regions such as ParaHippocampal Area 3 and VentroMedial Visual Area, highlighting potential disruptions in visual processing and memory formation.
-Similarly, the low-frequency range (Figure 3 B) showed significant differences centered around Area PH (Parahippocampal Area) and Primary Auditory Cortex , pointing to altered neural system how sound information is integrated and processed in memoryn.
-In the ultralow-frequency range (Figure 3 C), connectivity changes in VentroMedial Visual Area 3 and Area TF suggest impairments related to visual processing and attention regulation.
-In the ABIDE test dataset (HCP-MMP1 atlas), the results of the t-test distinguishing between the HC and ASD groups showed that 7.88% of the statistically significant connections ($p<0.05$) were in the high-frequency range, with the 75% quantile value of Cohen’s d score being 1.19.
-In contrast, the low-frequency range accounted for only 0.24%, with the 75% quantile value of Cohen’s d being 0.74.
-The ultralow frequency range accounted for 0.91%, with the 75% quantile value of Cohen’s d being 0.84.
-ASD-related connectivity patterns exhibited localized differences compared to healthy controls (HC), with varying effects across frequency ranges (Figure 3).
-In the high-frequency range (Figure 3 D), connectivity differences were observed in regions such as PGi (PG inferior) and 47m (Area 47 medial), potentially linked to challenges in socio-emotional and executive functions.
-In the low-frequency range (Figure 3 E), IFSa showed prominent differences, indicating possible impairments in executive and communication functions.
-Finally, in the ultralow-frequency range (Figure 3 F), connectivity alterations in Lateral Area 7P and the Insular Granular Complex suggest disruptions in sensory and emotional processing.
+ADHD exhibits frequency-dependent connectivity alterations with high-frequency disruptions predominating.
+Comparing ADHD patients to healthy controls in the ABCD dataset, we identified distinct patterns of connectivity changes across frequency bands (Figure 3 A-C).
+Statistical analysis revealed that high-frequency connections showed the most extensive alterations, with 8.43% of all significant connections ($p<0.05$) occurring in this band and a 75th percentile Cohen’s d of 0.332, indicating moderate effect sizes.
+In contrast, low-frequency and ultralow-frequency bands showed more limited but still significant changes, accounting for 3.46% (Cohen’s d 75th percentile = 0.220) and 3.56% (Cohen’s d 75th percentile = 0.117) of significant connections, respectively.
+Anatomically, high-frequency disruptions centered on Area 10v (ventral part of Brodmann Area 10), indicating impaired executive control networks.
+Low-frequency alterations consistently involved Area OP2-3-VS (opercular parts of Areas 2 and 3 and adjacent Ventral Somatosensory / Sylvian region) and Auditory 4 Complex, suggesting disrupted sensorimotor-auditory integration critical for attention regulation.
+Ultralow-frequency changes in Medial Belt Complex and ParaHippocampal regions, while showing smaller effect sizes, point to subtle alterations in memory and emotional processing circuits.
+ASD demonstrated a contrasting frequency profile with ultralow and low frequencies most severely affected (Figure 3 D-F).
+T-test comparisons between ASD patients and healthy controls in the ABIDE dataset revealed a different pattern from ADHD.
+Ultralow-frequency connections showed the most extensive disruptions, representing 4.54% of all significant connections with large effect sizes (75th percentile Cohen’s d = 0.717).
+Low-frequency alterations were similarly prominent, accounting for 4.04% of significant connections with substantial effect sizes (Cohen’s d 75th percentile = 0.714).
+Notably, high-frequency changes were minimal, comprising only 1.81% of significant connections, though with moderate effect sizes (Cohen’s d 75th percentile = 0.575).
+High-frequency changes, while limited in extent, localized specifically to Area 11l (lateral part of Brodmann Area 11), consistent with known social-cognitive deficits.
+Low-frequency disruptions prominently affected Area 10pp (frontopolar cortex), suggesting impaired higher-order executive functions.
+Most significantly, the extensive ultralow-frequency alterations involved Area TPOJ3 (temporo-parieto-occipital junction) and multiple frontal regions, indicating fundamental disruptions in social cognition and language networks—core features of ASD pathophysiology.
 In both the ABCD and ABIDE datasets, analysis of false positive and false negative subjects (based on predicted values) showed no statistically significant connections in the self-attention-based connectivity analysis.
+These frequency-resolved patterns reveal disorder-specific neural signatures: ADHD primarily affects high-frequency sensorimotor networks with compensatory ultralow-frequency changes, while ASD shows pervasive low and ultralow-frequency disruptions in social-cognitive circuits.
+Importantly, false positive and false negative cases showed no significant connectivity patterns, confirming the specificity of these frequency-dependent biomarkers for accurate psychiatric classification.
 3 Methods
 3.1 Preprocessing
-The ABCD, UKB, and ABIDE datasets underwent initial preprocessing prior to analysis.
-In the ABCD study, a bandpass filter ranging from 0.009 to 0.08 Hz was applied to harmonize multi-site data [23].
-In the UKB study, a bandpass filter ranging from 0.008 to 0.1 Hz was used to minimize the effects of heart rate variability and hemodynamic response delays [24].
-In the ABIDE study, a finite impulse response (FIR) filter with a passband of 0.01–0.1 Hz [25] was employed to facilitate multi-site data integration.
+The ABCD, UKB, and ABIDE datasets underwent dataset-specific preprocessing to account for acquisition differences and multi-site harmonization requirements.
+In ABCD, a conservative bandpass filter (0.009-0.08 Hz) was applied following consortium protocols for adolescent populations ([20]).
+UKB employed a broader filter (0.008-0.1 Hz) optimized for adult populations to minimize heart rate variability effects ([21]).
+ABIDE used finite impulse response filtering (0.01-0.1 Hz) specifically designed for multi-site ASD data integration ([22]).
 For the present study, additional preprocessing was performed on the resting-state BOLD data.
 The following preprocessing procedures were applied to each participant’s resting-state BOLD data: First, skull-stripping, slice-timing correction, and spatial normalization to MNI space were conducted using fMRIPrep ([23]).
-Subsequently, head motion artifacts and signals from white matter were removed using component-based noise correction (CompCor) [27].
-The mean time-series signals from predefined regions of interest (ROIs) were extracted using the Nilearn package, with HCP-MMP1 symmetric ([28]) and Schaefer 400 ([29]) atlases employed for ROI definitions.
+Subsequently, head motion artifacts and signals from white matter were removed using component-based noise correction (CompCor) ([24]).
+The mean time-series signals from predefined regions of interest (ROIs) were extracted using the Nilearn package, with HCP-MMP1 asymmetric ([25]) and Schaefer 400 ([26]) atlases employed for ROI definitions.
 Participants whose data contained voxels with NaN values, typically arising from the regression of unreliable white-matter signals, were excluded from the analysis.
@@
 Features UKB ABCD ABIDE
-Gender (M/F) 40,699 (21,556/19,143) 8,896 (4,662/4,234) 141 (114/27)
+Gender (M/F) 40,699 (21,556/19,143) 8,833 (4,635/4,198) 141 (114/27)
 Age (year) 54.9$\pm$7.5 - 14.3$\pm$3.6
 Age (month) - 119.2$\pm$7.5 -
-Depression (ph9 score) 11.5$\pm$3.5 - -
-Diagnosis (disorder/controls) - 141 (77/64) -
-Fluid intelligence (z-score) - 0.01$\pm$0.91 -
+Depression (PHQ-9 score) 11.5$\pm$3.5 - -
+Depression (disorder/controls) - 3,113 (2,652/461) -
+ADHD (disorder/controls) - 4,289 (2,439/1,850) -
+ASD (disorder/controls) - - 141 (77/64)
 Table 2: Features of the UKB, ABCD, and ABIDE subjects. Disorder denotes ADHD in the ABCD dataset and ASD in the ABIDE dataset.
@@
 Optimal pretraining epochs were set to 1,000 for the HCP-MMP1 atlas and 400 for the Schaefer 400 atlas.
-Additional methodological details are provided in Appendix A.2.
-3.4 Dividing frequencies
-3.4.1 Frequency Dividing Strategy
+Additional methodological details are provided in Supplementary Materials A.2.
+3.4 Frequency-resolved decomposition based on scale-free principles
+Brain activity exhibits multifractal properties with distinct scaling behaviors across frequency ranges, reflecting the hierarchical organization of neural networks from local circuits to global integration pathways.
+Conventional broadband connectivity analysis obscures these frequency-specific dynamics, potentially missing critical neurophysiological signatures of psychiatric disorders.
+To capture these frequency-dependent neural processes, we developed a systematic two-step fitting procedure to identify physiologically meaningful frequency boundaries (knee frequencies) that demarcate distinct scaling regimes in the power spectral density of resting-state fMRI signals.
+Following frequency decomposition using the identified knee frequencies, we confirmed that each of the three segments exhibited distinct power-law scaling behavior, as evidenced by significantly different $\beta$ values across frequency bands ($p<0.001$).
+3.4.1 Step 1: Ultralow frequency boundary identification using Lorentzian fitting
+We first applied the Lorentzian equation (Equation 2) to the entire power spectral density to identify $f_{1}$, the primary knee frequency that separates ultralow frequency activity from higher-frequency neural processes:
 $Power(f)=\frac{A\cdot f_{1}^{2}}{f^{2}+f_{1}^{2}}$ (2)
-Let $w(f)$ be defined as:
+where $A$ represents the amplitude parameter and $f_{1}$ denotes the characteristic frequency at which the power spectrum transitions from flat (low-frequency plateau) to steep decay (high-frequency roll-off).
+This Lorentzian model effectively captures the fundamental 1/f characteristics of neural signals while identifying the critical transition point.
+The fitting was performed using SciPy’s curve_fit function, which is well-suited for this two-parameter optimization problem.
+Physiologically, $f_{1}$ corresponds to the boundary between slow cortical potentials and faster network dynamics, typically associated with default mode network fluctuations and global integration processes.
+3.4.2 Step 2: Low-high frequency boundary identification using spline multifractal modeling
+For frequencies above $f_{1}$, we applied a sophisticated spline multifractal equation to identify $f_{2}$, the second knee frequency that distinguishes low-frequency network connectivity from high-frequency local processing:
+To ensure smooth transitions between different scaling regimes and prevent fitting artifacts, we implemented a cubic spline weight function:
 $w(f)=\begin{cases}0&\text{if }\log(f)<\log(f_{2})-s\\ 1&\text{if }\log(f)>\log(f_{2})+s\\ \frac{1}{2}(1-\cos(\frac{\pi(\log(f)-\log(f_{2})+s)}{2s}))&\text{otherwise}\end{cases}$ (3)
-Then the spline multifractal function is:
+This weight function $w(f)$ provides a smooth, differentiable transition between two distinct power-law regimes.
+When $f\ll f_{2}$, $w(f)\approx 0$, ensuring the function follows low-frequency scaling behavior characterized by $\beta_{low}$.
+Conversely, when $f\gg f_{2}$, $w(f)\approx 1$, shifting to high-frequency scaling defined by $\beta_{high}$.
+The smoothness parameter $s$ controls the sharpness of this transition, preventing abrupt discontinuities that could introduce fitting artifacts.
+The complete spline multifractal function then becomes:
 $Power(f)=A\cdot f^{\beta_{low}(1-w(f))+\beta_{high}w(f)}$ (4)
 where:
-  • $A$ is the amplitude.
-  • $\beta_{low}$ is the low-frequency scaling exponent.
-  • $\beta_{high}$ is the high-frequency scaling exponent.
-  • $f_{2}$ is the knee frequency.
-  • $s$ is the smoothness parameter.
-To segment the resting-state fMRI data into distinct frequency bands, we applied both Lorentzian and spline multifractal equations to the power spectral density (PSD).
-These methods allowed us to identify knee frequencies ($f_{1}$, $f_{2}$) that demarcate the transitions between frequency bands and reveal their scale-free properties.
-The Lorentzian equation (Equation 2) was used to estimate $f_{1}$, the first knee frequency, by fitting the entire PSD.
-For frequencies above $f_{1}$, the spline multifractal equation (Equation 4) was applied to identify $f_{2}$, the second knee frequency.
-To ensure a smooth transition between different frequency regimes, we implemented cubic spline smoothing in the multifractal equation, preventing artifacts that could arise from abrupt transitions at the knee frequencies.
-The smoothing weight was adapted based on methodologies from ([32, 33]).
-The transition function $w(f)$ plays a crucial role in interpolating between two distinct power-law regimes.
-When $x\ll f_{2}$, $w(f)\approx 0$, ensuring that the function follows the lower-frequency power-law behavior characterized by $\beta_{\text{low}}$.
-Conversely, when $x\gg f_{2}$, $w(f)\approx 1$, shifting the function to the higher-frequency power-law behavior defined by $\beta_{\text{high}}$.
-In the intermediate range, $w(f)$ smoothly transitions between these regimes, with the parameter $s$ controlling the sharpness of the transition.
-This formulation enables a continuous and differentiable shift between scaling regimes while preserving the structural properties of the signal.
-These knee frequencies were used to segment the time-series data into three frequency bands: ultra-low frequency (below $f_{1}$), low frequency (between $f_{1}$ and $f_{2}$), and high frequency (above $f_{2}$).
-The segmentation process was implemented using the Finite impulse response (FIR) and Boxcar filter from the nitime package.
-The optimal filter types (FIR and Boxcar) were determined through systematic hyperparameter tuning on validation sets, evaluating their performance impact on the overall model.
-The Lorentzian function, having only two parameters, was efficiently fitted using the curve fit function from the SciPy package, which is well-suited for low-dimensional fitting.
-In contrast, the spline multifractal function, which involves five parameters, was fitted using the iminuit package, which is optimized for high-dimensional parameter estimation.
-In all three datasets, ANOVA analysis of $f_{1}$ and $f_{2}$ computed separately for each dataset, revealed a statistically significant difference between the two values ($p<0.001$), indicating that $f_{1}$ and $f_{2}$ reflect distinct frequency characteristics.
-To assess the scale-free properties of these frequency bands, we computed $\beta$ values for each range, as shown in Equation 1.
-The results, summarized in Table 3, demonstrated significant differences in $\beta$ values across bands ($p<0.001$), supporting the hypothesis that each frequency component exhibits distinct scale-free dynamics.
-These results highlight that different frequency bands in the brain exhibit unique functional and topological properties.
-Specifically, lower frequencies are associated with long-range, global integration of brain activity, while higher frequencies correspond to more localized processing and rapid adaptation ([34, 35]).
-This finding reinforces the concept of multifractal behavior in brain signals, which has been previously observed in ECoG and EEG recordings ([20]).
-These $\beta$ values, denoted as $\beta_{ultralow}$, $\beta_{low}$, $\beta_{high}$, correspond to the ultra-low, low, and high-frequency bands, respectively.
-3.4.2 Frequency Divison and Physiological Relevance
-The values of $f_{1}$ and $f_{2}$ varied slightly across datasets (Table 2), yet consistently fell within specific ranges, reinforcing their physiological relevance.
-$f_{1}$ (0.046 Hz for ABCD, 0.062 Hz UKB datasets, and 0.018 Hz for ABIDE) corresponds to the time constant of an exponentially decaying function ([36]).
-Meanwhile, $f_{2}$ marks the transition to higher-frequency brain processes.
-$f_{2}$ was 0.078Hz for ABCD, 0.096Hz for UKB, and 0.030Hz for ABIDE.
-In ABCD and UKB, the $f_{2}$ values approaches the upper bound of conventional resting-state fMRI bandpass filtering (0.009-0.08Hz for ABCD, 0.008-0.1Hz for UKB).
-However, contemporary fMRI preprocessing pipelines commonly employ the Butterworth filter, which applies a gradual roll-off rather than an abrupt cutoff at upper bound (0.08Hz for ABCD, 0.1 Hz for UKB).
-This design minimizes phase distortion while effectively suppressing physiological noise from respiratory and cardiac ([37]).
-Consequently, the proximity of $f_{2}$ to upper bound of bandpass filtering does not imply a loss of meaningful higher-frequency information, as these signals are gradually attenuated rather than completely eliminated.
-The significant differences in $\beta$ values across frequency bands further validate the frequency-dependent multifractal properties of fMRI signals.
-This demonstrates the potential of frequency-specific analyses to uncover multi-scale brain network organization and adaptive complexity in brain activity.
-Additionally, the reliability of $f_{1}$ and $f_{2}$ across different datasets emphasizes the generalizability of this approach, making it a powerful tool for further research into brain dynamics.
+  • $A$ is the amplitude scaling factor
+  • $\beta_{low}$ is the low-frequency scaling exponent (typically steeper, reflecting long-range correlations)
+  • $\beta_{high}$ is the high-frequency scaling exponent (typically flatter, reflecting local processing)
+  • $f_{2}$ is the second knee frequency (transition point between network and local dynamics)
+  • $s$ is the smoothness parameter controlling transition width
+This five-parameter model was fitted using the iminuit package, specifically designed for high-dimensional parameter estimation with robust convergence properties.
+Physiologically, $f_{2}$ represents the transition from large-scale network connectivity to localized neural processing, typically corresponding to the boundary between task-negative and task-positive network activity.
+3.4.3 Frequency band segmentation and filtering implementation
+Based on the identified knee frequencies $f_{1}$ and $f_{2}$, we segmented each participant’s time-series data into three physiologically distinct frequency bands:
+  1. 1.
+    Ultra-low frequency band ($<f_{1}$): Captures slow cortical potentials, default mode network activity, and global integration processes
+  2. 2.
+    Low frequency band ($f_{1}$ to $f_{2}$): Represents intermediate-scale network connectivity and inter-regional communication
+  3. 3.
+    High frequency band ($>f_{2}$): Reflects local circuit dynamics, sensorimotor processing, and rapid cognitive operations
+The segmentation was implemented using finite impulse response (FIR) filters from the nitime package for ultra-low and low-frequency bands, and Boxcar filters for high-frequency extraction.
+The choice of filter types was optimized through systematic hyperparameter tuning on validation sets, evaluating their impact on downstream classification performance.
+FIR filters were preferred for lower frequencies due to their superior phase preservation properties, while Boxcar filters provided computational efficiency for high-frequency components without compromising signal quality.
+3.4.4 Validation of frequency-specific scaling properties
 Dataset $f_{1}$ $f_{2}$
@@
 Table 4: Different $\beta$ value among frequency ranges and datasets
-The spline multifractal equation and the Lorentzian equation showed superior fitting performance across all datasets as depicted in Table 4.
-The consistent and high goodness-of-fit metrics across different datasets ($R^{2}>0.66$ for all cases) demonstrate the effectiveness of our dual-equation fitting approach, validating its reliability for characterizing the frequency-domain properties of resting-state fMRI data.
 Dataset Lorentzian Equation Spline Multifractal Equation
@@
   • Note: Values are presented as mean±$\pm$±standard deviation. $R^{2}$ values indicate the proportion of variance explained by each model.
+To validate that our frequency decomposition captured distinct neurophysiological processes, we computed scaling exponents ($\beta$ values) for each frequency band using power-law fitting:
+Statistical analysis revealed significant differences in $\beta$ values across all frequency bands ($p<0.001$), confirming that each band exhibits unique scale-free dynamics.
+Ultralow frequencies showed the flattest scaling (smallest $\beta$), consistent with long-range temporal correlations.
+High frequencies exhibited intermediate scaling, reflecting network-level integration.
+Low frequencies displayed the steepest scaling (largest $\beta$), characteristic of local, rapidly decorrelated processes.
+These findings support the hypothesis that frequency-specific decomposition captures fundamentally different aspects of brain organization, from global integration to local specialization.
 3.5 Model Design
 3.5.1 Model Architecture
-Our model consists of two primary modules: the temporal module and the spatial module.
+MBBN integrates temporal and spatial processing through a dual-module architecture that consists of the temporal module and the spatial module.
 These modules work synergistically to encode both the temporal and spatial characteristics of frequency-divided timeseries data, as shown in Figure 1 A.
@@
 In parallel, a spatial loss is computed to enforce distinct spatial connectivity patterns across frequency bands.
-The total loss function, which combines these two components, is defined as:
+The composite loss function enforces both task performance and frequency-specific spatial patterns:
 $Loss_{total}=Loss_{task}(y_{true},\frac{\hat{y_{ultralow}}+\hat{y_{low}}+\hat{y_{high}}}{3})+\lambda\cdot Loss_{spatial}$ (5)
+where task-specific loss ensures predictive accuracy while spatial loss encourages distinct connectivity patterns across frequency bands.
 In Equation 5, $Loss_{task}$ represents the cross-entropy loss for classification tasks or L1 loss for regression tasks.
-The $Loss_{spatial}$ term, as described in Section 9, emphasizes the distinct spatial connectivity patterns inherent to each frequency band.
+The $Loss_{spatial}$ term, as described in Equation 9, emphasizes the distinct spatial connectivity patterns inherent to each frequency band.
 $\lambda$ was determined by hyperparameter tuning.
-The chosen $\lambda$ value and the filter type (described in Section 3.4.1) were shown in Table 5.
+The chosen $\lambda$ value and the filter type were shown in Table 6.
 Dataset Filter Type Spatial Factor $\lambda$
@@
 Although some previous studies have used convolutional filters as a preprocessing step before entering signals into BERT ([31]), we chose to bypass this step.
-Although convolutional preprocessing can be computationally efficient, it may sacrifice the preservation of fine-grained temporal dynamics within the fMRI data, which could impact model performance (see Appendix D).
+Although convolutional preprocessing can be computationally efficient, it may sacrifice the preservation of fine-grained temporal dynamics within the fMRI data, which could impact model performance (see Supplementary Materials D).
 By directly inputting the time-series data into the original BERT structure, we aimed to retain the full complexity of the temporal dynamics, allowing the model to effectively capture both short-term and long-term dependencies.
@@
 $Loss_{spatial}=-log(Loss_{high-low}+Loss_{high-ultralow}+Loss_{ultralow-low})$ (9)
-3.8 Pretraining Strategy Based on Communicability
+The negative logarithm amplifies subtle connectivity differences, ensuring the model learns distinct spatial patterns even when frequency bands show similar overall connectivity structures.
+We confirmed that this approach indeed derives neurobiologically plausible network properties (see Appendix B).
+3.8 Communicability-based Pretraining Strategy
 3.8.1 Masking Strategy and Pretraining Loss
@@
 To achieve this, we incorporated the concept of communicability into the pretraining process.
-A detailed definition of communicability used in this study is provided in Appendix A.1.
+A detailed definition of communicability used in this study is provided in Supplementary Materials A.1.
+High-communicability nodes act as critical hubs in brain networks, integrating information across distributed regions.
+By masking these influential nodes during pretraining, we force the model to learn their functional roles through contextual inference, similar to masked language modeling but tailored to brain network topology.
 Our pretraining strategy draws inspiration from masking techniques commonly used in language models to capture contextual relationships and data patterns ([30]).
@@
   • Randomly masking regions of interest (ROIs).
-Detailed model performance results (Appendix E.1) show that masking high-communicability nodes led to the most effective learning of the data structure, resulting in superior performance in downstream tasks.
+Detailed model performance results (Supplementary Materials E.1) show that masking high-communicability nodes led to the most effective learning of the data structure, resulting in superior performance in downstream tasks.
 Figure 4: MBBN Pretraining Workflow.
@@
 Pretraining involved two complementary masking strategies: temporal masking and spatial masking (Figure 4).
-Spatial masking involved nullifying the timeseries data of the top $k$ nodes with the highest communicability scores, where $k$ is less than the total number of ROIs.
+Spatial masking involved masking the timeseries data of the top $k$ nodes with the highest communicability scores, where $k$ is less than the total number of ROIs.
 Temporal masking, on the other hand, initialized all ROI signals to zero within predefined time windows.
@@
 Traditional attribution methods are not directly applicable to this model, as the relationship between attention matrices and target values is indirect.
-To address this challenge, we developed a novel approach inspired by GradCAM ([43]), specifically tailored to the spatial attention module described in Section 3.7.
+To address this challenge, we developed a GradCAM ([34])-inspired approach to identify frequency-specific connections driving psychiatric classification.
+This method highlights connections where gradient magnitude (importance) aligns with attention strength (activation), revealing clinically relevant biomarkers.
+$Heatmap_{k}=\frac{\partial{Loss}_{spatial}}{\partial f_{k}(input_{k})}\odot f_{k}(input_{k})$ (12)
 GradCAM typically highlights important regions by computing the gradient of the loss for the target class and weighting it by the output feature maps.
@@
 To address this, we selected only a few top connections based on the absolute values of Cohen’s d for further evaluation of their impact.
-An exhaustive list of all significant connections can be found in Appendix F.
-$Heatmap_{k}=\frac{\partial{Loss}_{spatial}}{\partial f_{k}(input_{k})}\odot f_{k}(input_{k})$ (12)
+An exhaustive list of all significant connections can be found in Supplementary Materials F.
 3.10 Performance Comparison with Baseline Models
-To evaluate the effectiveness of MBBN in learning spatiotemporal dynamics, we compared its performance with baseline models, including functional connectivity-based models (XGBoost [45], BNT [5]) and temporal dynamics models (BolT ([6]), vanilla BERT ([39])).
+To evaluate the effectiveness of MBBN in learning spatiotemporal dynamics, we compared its performance with baseline models, including functional connectivity-based models (XGBoost ([36]), BNT ([5])) and temporal dynamics models (BolT ([6]), vanilla BERT ([30])).
 These models were selected based on their relevance to brain dynamics interpretation, either through static functional connectivity or sequential processing of temporal data.
@@
 While MBBN demonstrated superior performance across several downstream tasks, including the ABCD ADHD classification task, with significant improvements in accuracy and AUC scores, not all tasks showed a statistically significant advantage.
-Detailed performance comparison is described in Appendix E.1.
+Detailed performance comparison is described in Supplementary Materials E.1.
 To ensure a fair evaluation, we conducted statistical tests (paired t-tests) to compare the performance of MBBN with baseline models on each task.
@@
 4 Discussion
-Discussions should be brief and focused.
-In some disciplines use of Discussion or ‘Conclusion’ is interchangeable.
-It is not mandatory to use both.
-Some journals prefer a section ‘Results and Discussion’ followed by a section ‘Conclusion’.
-Please refer to Journal-level guidance for any specific requirements.
-This study introduces MBBN, a novel computational framework that integrates neuroscientific principles with advanced deep learning methodologies to analyze brain dynamics.
-MBBN employs a scale-free, frequency-based decomposition technique that encodes distinct frequency components as individual features, effectively capturing the multi-fractal dynamics of brain activity.
-Additionally, the model leverages a communicability-based masking strategy during pretraining, embedding the underlying network topology into its learned representations.
-Through a self-attention mechanism, MBBN combines biologically informed spatial relationships with temporal context, enabling the nuanced modeling of connectivity patterns and their evolution across multiple frequency ranges.
-This unified framework provides a robust platform for accurate, interpretable, and personalized predictions of cognitive and clinical outcomes.
-Compared to previous transformer-based models for brain imaging data (e.g., [8, 7, 5, 6]), MBBN uniquely integrates the multi-fractal ([46, 47, 48]) and scale-free properties of neural signals ([36, 49]), essential for understanding the hierarchical and dynamic nature of brain activity ([50, 51, 52]).
-Unlike models relying solely on static or sequential representations ([5, 6]), or employing simplistic masking strategies ([7, 8]), MBBN’s biologically grounded approach—decomposing multifractal neural signals into interpretable frequency components and utilizing a communicability-driven masking scheme—enables it to effectively capture and reconstruct critical network hubs.
-By doing so, MBBN provides a more nuanced and biologically relevant representation of spatiotemporal connectivity.
-Crucially, correlation analysis confirms that the connectivity derived from MBBN’s self-attention mechanism represents a novel measure, capturing information distinct from and complementary to traditional functional connectivity (FC).
-To elucidate this, we directly compared attention-based connectivity with FC.
-This revealed a strong negative correlation when analyzing raw, non-frequency-divided signals (Spearman $\rho\approx-0.6$), indicating that attention-based connectivity captures fundamentally different, complementary information rather than merely replicating static FC patterns.
-Of note, the correlations became considerably weaker within specific frequency bands (ranging from $\rho\approx-0.1\backsim-0.2$) (Supplementary Table 2).
-These findings strongly suggest that MBBN successfully extracts unique, frequency-dependent connectivity patterns that are distinct from both traditional FC and the aggregated signal.
-This divergence not only validates our attention mechanism as providing a novel and meaningful measure of neural interaction but also underscores the importance of frequency-specific analysis for uncovering nuanced brain dynamics potentially missed by conventional methods.
+This study presents MBBN, a novel deep learning framework that integrates biologically grounded principles with frequency-aware self-attention to model spatiotemporal brain dynamics.
+By leveraging scale-free and multifractal properties of neural signals, MBBN captures essential patterns of brain organization that are often overlooked by conventional transformer-based models.
+Additionally, MBBN employs communicability-based masking and frequency-aware self-attention to capture biologically meaningful connectivity patterns across multiple frequency ranges, enabling accurate and interpretable predictions of cognitive and clinical outcomes.
+Unlike previous approaches that rely on broadband or static representations ([5, 6]), MBBN applies a frequency decomposition strategy to encode neural oscillations as discrete features, enabling a richer characterization of dynamic connectivity patterns.
+This frequency-resolved approach is particularly critical given that neural communication is mediated by distinct oscillatory dynamics across frequency bands, with high-frequency oscillations supporting local processing ([37]) and low-frequency rhythms facilitating long-range integration ([38]).
+Traditional broadband or narrowband analyses often obscure these nuanced patterns by averaging across spectrally diverse signals, potentially masking functionally specific effects.
+The MBBN framework overcomes these limitations by capturing such frequency-specific properties, allowing for a more precise and physiologically grounded understanding of brain network dynamics.
+Furthermore, our communicability-based ([39, 40]) masking method ensures that topological structures within brain networks are embedded into learned representations, improving both robustness and interpretability.
+Importantly, MBBN-derived connectivity patterns complement rather than replicate traditional functional connectivity.
+While moderate anticorrelation existed at the broadband level, frequency-specific analyses revealed unique network dynamics (Supplementary Table 2), confirming that MBBN captures neurophysiologically distinct information critical for clinical interpretation.
 The biological plausibility of MBBN is further supported by its incorporation of the scale-free property, characterized by a power-law exponent ($\beta$), which differentiates brain states and regions based on their temporal correlation profiles ([2, 41, 42, 43, 44, 45]).
@@
 This approach establishes MBBN as a foundational framework for advancing neuroimaging analysis.
-MBBN demonstrated superior performance across diverse datasets, including UK Biobank, ABCD, and ABIDE, underscoring its robustness across different age groups and developmental stages.
-While UK Biobank includes primarily middle-aged adults, ABCD and ABIDE focus on children and adolescents during critical brain maturation periods.
-Unlike most functional brain transformer models ([8, 7, 5]), which are rarely validated on heterogeneous cohorts, MBBN effectively handles lifespan variability, suggesting that it captures fundamental principles of brain organization.
-This enables accurate predictions of clinical and cognitive outcomes across ages and provides insights into developmental trajectories and age-specific neurological markers.
-Our findings revealed distinct, frequency-specific connectivity alterations distinguishing individuals with ADHD and ASD from healthy controls.
-In ADHD, problematic connectivity was observed in regions essential for visual and audio processing, memory formation, and attention regulation with disruptions varying significantly across frequency ranges.
-In the high-frequency range, we observed significant reductions in connectivity between the Parahippocampal area 3 and multiple early visual regions, as well as between the Ventromedial visual area 2 and primary/third visual cortices.
-These findings imply impaired integration between visual information and memory or contextual processing systems, which may be linked to difficulties in sustaining attention, interpreting visual scenes, and integrating incoming stimuli in ADHD ([63, 64]).
-In the low-frequency range, decreased connectivity emerged between the Parahippocampal area and anterior temporal, perirhinal, and orbitofrontal regions—networks typically involved in episodic memory, semantic integration, and affective regulation ([65, 66]).
-Additionally, weakened connections between the primary auditory cortex and multimodal association areas point to disruptions in sensory processing ([67]).
-These alterations suggest that ADHD may be characterized by deficits in slower, long-range communication necessary for high-level cognitive integration and stable internal state representation.
-Conversely, the ultralow-frequency range showed increased connectivity in ADHD, particularly between the Ventromedial visual area 3 and spatial attention regions (e.g., Lateral intraparietal area ventral, medial area 7A), as well as between emotion-related regions such as Area TF (the posterior subdivision of the Parahippocampal cortex), Area 25 (Subgenual cingulate cortex), and the anterior cingulate cortex (s32).
-This pattern may reflect excessive internal processing, emotional dysregulation, and inefficient attentional gating ([68]).
-Together, these findings indicate that ADHD involves a dynamic reconfiguration of functional brain networks across multiple timescales—with reduced integration in fast and intermediate-range communication, and hyperconnectivity in the slowest neural rhythms—potentially contributing to the core symptoms of inattention, emotional lability, and sensory processing anomalies.
-Our frequency-specific analysis of the ABIDE dataset for ASD prediction revealed divergent patterns of functional connectivity alterations across spectral bands, underscoring the multi-timescale nature of network disruptions in autism spectrum disorder.
-In the high-frequency range, the ASD group exhibited markedly increased connectivity, particularly involving Area PGs (Posterior superior parietal lobule ), Area 8 anterior ventral (within the prefrontal cortex), and Area 8Ad (Anterior dorsal prefrontal cortex) with regions implicated in executive control and higher-order social cognition (e.g., Area p32, Area 9 Middle, Dorsal Area 24d).
-These patterns suggest hyperconnectivity in frontoparietal and cingulo-opercular circuits, potentially reflecting compensatory or maladaptive neural synchronization related to cognitive control and social attention ([69, 70]).
-In the low-frequency range, increased connectivity was observed between multiple sensory and associative areas (e.g., Area PH, Primary Sensory Cortex, Auditory and ParaBelt Complexes) and Area posterior 9–46 ventral.
-This may reflect overintegration or inefficient filtering in low-frequency sensory-cognitive loops, contributing to the sensory hypersensitivity and perceptual integration challenges frequently reported in ASD ([71, 72]).
-Conversely, the ultralow-frequency range showed decreased connectivity in the ASD group compared to controls, notably in connections between Parahippocampal area 1, Area PH, and auditory complexes with Area 7PC (Precuneus) and Area Frontal Opercular 5.
-These connections are implicated in high-level integrative functions such as spatial awareness, self-referential processing, and multimodal integration.
-Reduced ultralow-frequency connectivity may reflect a breakdown in global neural coordination and internal state regulation, core deficits commonly implicated in autism ([73, 74]).
-Together, these findings reveal a frequency-dependent reconfiguration of functional brain networks in ASD, with fronto-executive hyperconnectivity at higher frequencies and reduced long-range integrative coupling at the slowest rhythms—highlighting potential biomarkers and targets for intervention across multiple neurophysiological timescales.
-Connectivity patterns observed across frequency bands revealed a transition from local to more widespread disruptions in both ADHD and ASD.
-In ADHD, high-frequency alterations were primarily local, focusing on visual and memory-related circuits, while ultralow-frequency changes spanned broader fronto-parietal and cingulo-limbic networks.
-Similarly, ASD showed local hyperconnectivity in high and low-frequency bands, especially in frontal and sensory areas, with broader hypoconnectivity emerging only in the ultralow-frequency range.
-These results suggest that while both disorders exhibit frequency-specific connectivity alterations, the spatial scale of disruption expands with slower neural oscillations.
-In sum, MBBN’s multifaceted and biologically informed approach to modeling spatiotemporal brain dynamics represents a significant advancement over existing transformer-based frameworks.
-Its comprehensive integration of neuroscientific principles, self-attention mechanisms, and frequency-specific analyses positions MBBN as a promising foundation for more accurate, interpretable, and clinically meaningful neuroscience research.
+Across diverse cohorts—including UK Biobank, ABCD, and ABIDE—MBBN consistently demonstrated strong predictive performance, capturing neural organizational principles across age groups.
+Unlike many existing models that are validated only on narrow demographics, MBBN successfully generalized across developmental stages, revealing its potential as a lifespan-inclusive neuroimaging tool.
+Our frequency-specific analyses uncovered distinct connectivity disruptions in ADHD and ASD.
+In ADHD, we observed attenuated high-frequency connectivity in fronto-sensorimotor networks, which may underlie core symptoms such as inattention and executive dysfunction ([51, 52]).
+At lower frequencies, altered sensory and language processing circuits were evident ([53, 54, 55, 56]), and in the ultralow-frequency range, both hypo- and hyperconnectivity patterns suggested inefficient compensatory mechanisms, particularly involving the parahippocampal and cingulate regions ([57, 58, 59, 60, 61]).
+A striking finding was the frequency-dependent shift in Area OP2-3-VS (Opercular parts of Areas 2 and 3 and adjacent Ventral Somatosensory / Sylvian region) connectivity in ADHD.
+This region consistently emerged as a central node across different frequency bands, suggesting it may serve as a dynamic hub that potentially coordinates neural communication based on oscillatory states.
+Rather than subserving a fixed function, Area OP2-3-VS appears to assume context-dependent roles, possibly regulating cross-frequency information flow or facilitating inter-network switching—mechanisms that could be critical for coherent cognitive function.
+In ADHD, disruptions in these putative regulatory processes may contribute to frequency-specific connectivity alterations and the disorder’s heterogeneous symptomatology, underscoring the potential importance of spectral dynamics in characterizing functional hubs.
+In ASD, high-frequency hubs involving orbitofrontal and somatosensory circuits were prominent, consistent with known deficits in emotional and social processing ([62, 63, 64, 65]).
+At low frequencies, the posterior frontopolar cortex emerged as a key integrative hub; disruptions here may impair the coordination of memory, decision-making, and multisensory integration essential for goal-directed behavior ([66, 67, 68, 69]).
+Interestingly, ultralow-frequency analyses revealed heightened connectivity between the temporo-parietal junction and prefrontal regions, possibly reflecting compensatory cognitive strategies or atypical resource allocation during social cognition tasks ([70, 71, 72, 73]).
+Collectively, these results illustrate that MBBN not only outperforms prior models in predictive capacity but also provides interpretable biomarkers aligned with known neurodevelopmental patterns.
+By aligning transformer-based modeling with core neuroscientific principles—particularly the relevance of neural oscillations—MBBN offers a scalable, biologically plausible framework for understanding individualized brain dynamics.
+While MBBN offers notable methodological advancements, some limitations warrant consideration.
+First, our frequency decomposition employs theory-based spectral bands.
+Although grounded in established neurophysiological knowledge, this approach may not capture finer-grained oscillatory phenomena that deviate from these canonical bands.
+Future work could explore more adaptive approaches to move beyond theory-based bands, for example, by employing more granular spectral tiling (e.g., via wavelet packet decomposition [74]) or by developing a fully learnable filter-bank layer within the MBBN architecture.
+This might allow for end-to-end optimization of frequency bands for specific downstream tasks.
+Second, the communicability-based masking, while effective in promoting network-aware learning, is computationally demanding.
+To enhance scalability for ever-larger datasets, future iterations could investigate computationally efficient approximations of communicability, such as those based on polynomial expansions or random walks ([40]).
+Third, the assessment of test-retest reliability for the derived frequency-specific connectivity patterns would be crucial.
+Establishing such reliability is not only fundamental to evaluating the robustness of the MBBN model itself but also essential for validating the temporal stability of these patterns.
+This is particularly important for applications involving the tracking of neurodevelopmental trajectories or disease progression, where consistent and dependable measures are paramount ([75]).
+Future extensions of MBBN could incorporate multimodal data (e.g., structural MRI, diffusion MRI) and expand into task-based paradigms to examine how frequency-specific connectivity evolves under cognitive demands.
+Additionally, exploring cross-frequency coupling may offer deeper insights into hierarchical brain dynamics beyond individual frequency bands.
+MBBN establishes a new standard for interpretable, frequency-resolved brain modeling.
+By integrating neurophysiological principles with transformer architectures, it offers both predictive power and mechanistic insights into brain function, laying the foundation for precision neuroscience applications.
 Data and Code Availability
@@
 Ultralow frequency 0.0350 0.0991 4.8600 1.4794 3.2850 8.3929 0.9229 9.0941
-Raw (Non-dividing) 0.0630 -0.0000 7.7804 1.6143 4.8197 5.9527 0.9254 6.4329
+Raw (Non-dividing) 0.0630 $<0.0001$ 7.7804 1.6143 4.8197 5.9527 0.9254 6.4329
 Supplementary Table 1: Comparison of Network Properties between FC and Attention Matrix
@@
   2. 2.
-    Contrasting patterns with FC: While FC exhibits its highest modularity in the raw (unfiltered) condition (0.0630), attention-based connectivity shows essentially no modularity in the raw condition (-0.0000).
+    Contrasting patterns with FC: While FC exhibits its highest modularity in the raw (unfiltered) condition (0.0630), attention-based connectivity shows essentially no modularity in the raw condition ($<0.0001$).
 This striking difference suggests that our attention mechanism captures frequency-specific network organization that is lost when analyzing the full frequency spectrum together.
@@
 Ultra-low Frequency 0.870$\pm$0.010 0.779$\pm$0.013 0.935$\pm$0.003 0.862$\pm$0.007
-Divided Frequencies 0.916$\pm$0.004 0.844$\pm$0.006 0.975$\pm$0.001 0.920$\pm$0.001
+Divided Frequencies 0.916$\pm$0.004 0.844$\pm$0.006 0.980$\pm$0.001 0.920$\pm$0.001
 Supplementary Table 3: Results of BERT models trained with single frequency range in the timeseries data extracted by HCP-MMP1 atlas
@@
 MBBN with conv 0.825$\pm$0.009 0.726$\pm$0.009 0.947$\pm$0.003 0.870$\pm$0.002
-MBBN without conv 0.916$\pm$0.004 0.844$\pm$0.006 0.975$\pm$0.001 0.920$\pm$0.001
+MBBN without conv 0.916$\pm$0.004 0.844$\pm$0.006 0.980$\pm$0.001 0.920$\pm$0.001
 Supplementary Table 4: Results of BERT models with a convolutional layer before BERT model
@@
 Random 3 Freq 0.900$\pm$0.006 0.803$\pm$0.008 0.972$\pm$0.002 0.918$\pm$0.003
-Ours 0.916$\pm$0.004 0.844$\pm$0.006 0.975$\pm$0.001 0.920$\pm$0.001
+Ours 0.916$\pm$0.004 0.844$\pm$0.006 0.980$\pm$0.001 0.920$\pm$0.001
 Supplementary Table 5: Results of BERT variants trained with different frequency-dividing strategies in the timeseries data extracted by HCP-MMP1 atlas
@@
 Temporal 0.909$\pm$0.007 0.831$\pm$0.011 0.964$\pm$0.001 0.915$\pm$0.001
-Spatiotemporal (MBBN) 0.916$\pm$0.004 0.844$\pm$0.006 0.975$\pm$0.001 0.921$\pm$0.006
+Spatiotemporal (MBBN) 0.916$\pm$0.004 0.844$\pm$0.006 0.980$\pm$0.001 0.921$\pm$0.006
 Supplementary Table 6: Results of spatial and temporal modules of MBBN in the timeseries data extracted by HCP-MMP1 atlas
@@
 Appendix E Comparison with other models
-E.1 Perormance comparison
+E.1 Performance comparison
 ABCD sex UKB sex
 Models (data form) AUROC ↑ Accuracy ↑ AUROC ↑ Accuracy ↑
-XGBOOST (CC) 0.768$\pm$0.004 0.696$\pm$0.001 0.859$\pm$003 0.768$\pm$0.005
-BNT (CC) 0.895$\pm$0.001 0.814$\pm$0.010 0.950$\pm$0.004 0.864$\pm$0.017
-BolT (TD) 0.883$\pm$0.001 0.795$\pm$0.001 0.950$\pm$0.000 0.878$\pm$0.001
-vanilla BERT (TD) 0.884$\pm$0.006 0.778$\pm$0.027 0.942$\pm$0.014 0.876$\pm$0.006
-MBBN (TD) 0.916$\pm$0.004 0.844$\pm$0.006 0.975$\pm$0.001 0.921$\pm$0.006
+XGBOOST (CC) 0.785$\pm$0.005 0.709$\pm$0.003 0.859$\pm$003 0.768$\pm$0.005
+BNT (CC) 0.872$\pm$0.005 0.795$\pm$0.007 0.950$\pm$0.004 0.864$\pm$0.017
+BolT (TD) 0.886$\pm$0.002 0.795$\pm$0.002 0.950$\pm$0.000 0.878$\pm$0.001
+vanilla BERT (TD) 0.864$\pm$0.004 0.777$\pm$0.006 0.942$\pm$0.014 0.876$\pm$0.006
+MBBN (TD) 0.913$\pm$0.002 0.823$\pm$0.009 0.980$\pm$0.001 0.921$\pm$0.006
 Supplementary Table 8: Results of models trained from scratch predicting biological phenotypes. Atlas is HCP MMP 1. CC denotes correlation-based connectivity and TD denotes temporal dynamics.
@@
 Models AUROC ↑ Accuracy ↑ AUROC ↑ Accuracy ↑
-XGBOOST (CC) 0.812$\pm$0.004 0.734$\pm$0.003 0.905$\pm$0.004 0.884$\pm$0.002
+XGBOOST (CC) 0.718$\pm$0.004 0.659$\pm$0.003 0.905$\pm$0.004 0.884$\pm$0.002
 BNT (CC) 0.920$\pm$0.006 0.839$\pm$0.012 0.982$\pm$0.001 0.934$\pm$0.092
@@
 XGBOOST (CC) 2.417$\pm$0.025 12.342$\pm$0.151 0.548$\pm$0.026 0.829$\pm$0.000
-BNT (CC) 2.276$\pm$0.015 13.989$\pm$0.682 0.536$\pm$0.008 0.849$\pm$0.003
-BolT (TD) 2.217$\pm$0.022 12.729$\pm$0.290 0.541$\pm$0.016 0.843$\pm$0.003
-vanilla BERT (TD) 2.194$\pm$0.039 12.751$\pm$0.780 0.5147$\pm$0.014 0.847$\pm$0.003
-MBBN (TD) 2.210 $\pm$0.041 13.205$\pm$0.956 0.556$\pm$0.025 0.718$\pm$0.028
+BNT (CC) 2.276$\pm$0.015 13.989$\pm$0.682 0.546$\pm$0.015 0.847$\pm$0.006
+BolT (TD) 2.217$\pm$0.022 12.729$\pm$0.290 0.588$\pm$0.006 0.847$\pm$0.001
+vanilla BERT (TD) 2.194$\pm$0.039 12.751$\pm$0.780 0.546$\pm$0.015 0.818$\pm$0.051
+MBBN (TD) 2.210 $\pm$0.041 13.205$\pm$0.956 0.590$\pm$0.013 0.750$\pm$0.008
 Supplementary Table 10: Results of models trained from scratch predicting clinical outcomes, atlas is HCP-MMP1. CC denotes correlation-based connectivity and TD denotes temporal dynamics.
@@
 Models (data form) MAE ↓ MSE ↓ MAE ↓ MSE ↓
-XGBOOST (CC) 1.609$\pm$0.009 4.078$\pm$0.086 0.787$\pm$0.002 0.987$\pm$0.033
-BNT (CC) 1.581$\pm$0.023 3.993$\pm$0.516 0.793$\pm$0.008 1.061$\pm$0.093
+XGBOOST (CC) 1.609$\pm$0.009 4.078$\pm$0.086 0.720$\pm$0.003 0.827$\pm$0.007
+BNT (CC) 1.581$\pm$0.023 3.993$\pm$0.516 0.763$\pm$0.020 1.061$\pm$0.093
 BolT (TD) 1.589$\pm$0.002 3.973$\pm$0.006 0.728$\pm$0.004 0.845$\pm$0.010
-vanilla BERT (TD) 1.597$\pm$0.015 4.081$\pm$0.094 0.722$\pm$0.005 0.838$\pm$0.036
-MBBN (TD) 1.630 $\pm$0.006 4.225$\pm$0.125 0.717$\pm$0.012 0.825$\pm$0.034
+vanilla BERT (TD) 1.597$\pm$0.015 4.081$\pm$0.094 0.699$\pm$0.016 0.788$\pm$0.036
+MBBN (TD) 1.630 $\pm$0.006 4.225$\pm$0.125 0.694$\pm$0.018 0.774$\pm$0.041
 Supplementary Table 12: Results of models trained from scratch predicting cognitive outcomes, atlas is HCPMMP1. CC denotes correlation-based connectivity and TD denotes temporal dynamics.
@@
 Models (data form) AUROC ↑ Accuracy ↑ AUROC ↑ Accuracy ↑
-XGBOOST (CC) 0.598$\pm$0.010 0.570$\pm$0.005 0.531$\pm$0.020 0.487$\pm$0.006
-BNT (CC) 0.634$\pm$0.008 0.606$\pm$0.014 0.575$\pm$0.013 0.567$\pm$0.019
-BolT (TD) 0.603$\pm$0.005 0.592$\pm$0.010 0.650$\pm$0.064 0.598$\pm$0.057
-Vanilla BERT (TD) 0.609$\pm$0.020 0.592$\pm$0.015 0.583$\pm$0.004 0.542$\pm$0.021
-MBBN [from scratch] (TD) 0.598$\pm$0.009 0.589$\pm$0.024 0.654$\pm$0.134 0.521$\pm$0.059
+XGBOOST (CC) 0.598$\pm$0.010 0.570$\pm$0.005 0.573$\pm$0.053 0.546$\pm$0.046
+BNT (CC) 0.634$\pm$0.008 0.606$\pm$0.014 0.691$\pm$0.168 0.617$\pm$0.062
+BolT (TD) 0.603$\pm$0.005 0.592$\pm$0.010 0.632$\pm$0.005 0.566$\pm$0.003
+Vanilla BERT (TD) 0.609$\pm$0.020 0.592$\pm$0.015 0.644$\pm$0.089 0.625$\pm$0.051
+MBBN [from scratch] (TD) 0.633$\pm$0.009 0.593$\pm$0.015 0.780$\pm$0.012 0.657$\pm$0.038
 MBBN [random] (TD) 0.603$\pm$0.032 0.589$\pm$0.026 0.601$\pm$0.180 0.534$\pm$0.082
 MBBN [low] (TD) 0.615$\pm$0.009 0.602$\pm$0.011 0.583$\pm$0.084 0.547$\pm$0.098
-MBBN [high] (TD) 0.628$\pm$0.001 0.609$\pm$0.011 0.672$\pm$0.079 0.646$\pm$0.078
+MBBN [high] (TD) 0.645$\pm$0.011 0.601$\pm$0.015 0.810$\pm$0.065 0.717$\pm$0.019
 Supplementary Table 14: Finetuning results of pre-trained BERT models with different pretraining loss (HCP-MMP1 atlas).
@@
 Vanilla BERT (TD) 0.604$\pm$0.017 0.600$\pm$0.008 0.664$\pm$0.148 0.667$\pm$0.059
-MBBN [from scratch] (TD) 0.614$\pm$0.020 0.610$\pm$0.011 0.800$\pm$0.067 0.563$\pm$0.051
+MBBN [from scratch] (TD) 0.614$\pm$0.020 0.610$\pm$0.011 0.800$\pm$0.067 0.683$\pm$0.062
 MBBN [random] (TD) 0.596$\pm$0.014 0.593$\pm$0.032 0.696$\pm$0.052 0.563$\pm$0.088
 MBBN [low] (TD) 0.584$\pm$0.033 0.574$\pm$0.034 0.500$\pm$0.051 0.479$\pm$0.059
-MBBN [high] (TD) 0.618$\pm$0.019 0.580$\pm$0.035 0.822$\pm$0.019 0.563$\pm$0.051
+MBBN [high] (TD) 0.618$\pm$0.019 0.580$\pm$0.035 0.822$\pm$0.019 0.768$\pm$0.051
 Supplementary Table 15: Finetuning results of pre-trained BERT models with different pretraining loss (Schaefer 400 atlas).
@@
 Models (data form) Number of parameters FLOPs
-XGBOOST (CC) 6100 400
-BNT (CC) 3,787,250 1,805,681,216
-BolT (TD) 9,597,062 4,226,217,120
-Vanilla BERT (TD) 10,050,757 23,497,736,096
-MBBN (TD) 11,140,695 70,528,086,432
+XGBOOST (CC) 6004 400
+BNT (CC) 11,678,810 13,084,233,536
+BolT (TD) 20,697,661 50,032,700,640
+Vanilla BERT (TD) 22,280,137 69,848,980,256
+MBBN (TD) 23,370,075 209,581,818,912
 Supplementary Table 16: The table shows the number of parameters and FLOPs of baseline models used in this study and MBBN (our model).
@@
 Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
-Fourth Visual Area ParaHippocampal Area 3 $<0.001$ HC 0.654
-Second Visual Area ParaHippocampal Area 3 $<0.001$ HC 0.645
-Polar 10p ProStriate Area $<0.001$ HC 0.643
-Primary Visual Cortex VentroMedial Visual Area 2 $<0.001$ HC 0.638
-Area PFt ParaHippocampal Area 3 $<0.001$ HC 0.636
-Third Visual Area VentroMedial Visual Area 2 $<0.001$ HC 0.636
-Second Visual Area VentroMedial Visual Area 2 $<0.001$ HC 0.633
-Second Visual Area ProStriate Area $<0.001$ HC 0.633
-Area Lateral Occipital 2 VentroMedial Visual Area 2 $<0.001$ HC 0.627
-Third Visual Area VentroMedial Visual Area 1 $<0.001$ HC 0.627
-Area 11l ProStriate Area $<0.001$ HC 0.626
-Third Visual Area ParaHippocampal Area 3 $<0.001$ HC 0.623
-Primary Visual Cortex ParaHippocampal Area 3 $<0.001$ HC 0.618
-Posterior OFC Complex Frontal Opercular Area 3 $<0.001$ HC 0.616
-Area 2 ParaHippocampal Area 3 $<0.001$ HC 0.613
-Posterior OFC Complex Anterior Agranular Insula Complex $<0.001$ HC 0.612
-Third Visual Area ProStriate Area $<0.001$ HC 0.611
-Fourth Visual Area VentroMedial Visual Area 2 $<0.001$ HC 0.61
-Area posterior 10p Frontal Opercular Area 4 $<0.001$ HC 0.606
-Area 47m Frontal Opercular Area 4 $<0.001$ HC 0.606
-Posterior OFC Complex Frontal Opercular Area 4 $<0.001$ HC 0.604
-Anterior IntraParietal Area VentroMedial Visual Area 1 $<0.001$ HC 0.601
-Area 47l (47 lateral) Frontal Opercular Area 4 $<0.001$ HC 0.597
-Area 10r Frontal Opercular Area 4 $<0.001$ HC 0.596
-Area 10r Frontal Opercular Area 3 $<0.001$ HC 0.589
-Area 11l Dorsal Transitional Visual Area $<0.001$ HC 0.589
-Primary Visual Cortex VentroMedial Visual Area 1 $<0.001$ HC 0.589
-Area TemporoParietoOccipital Junction 1 ProStriate Area $<0.001$ HC 0.587
-Second Visual Area VentroMedial Visual Area 1 $<0.001$ HC 0.587
-Orbital Frontal Complex Frontal Opercular Area 4 $<0.001$ HC 0.587
-Area 13l Anterior Agranular Insula Complex $<0.001$ HC 0.587
-Fourth Visual Area VentroMedial Visual Area 1 $<0.001$ HC 0.584
-Posterior InferoTemporal complex VentroMedial Visual Area 2 $<0.001$ HC 0.582
-Orbital Frontal Complex Anterior Agranular Insula Complex $<0.001$ HC 0.581
-Polar 10p Area 23c $<0.001$ HC 0.578
-Polar 10p VentroMedial Visual Area 1 $<0.001$ HC 0.578
-Fourth Visual Area Frontal Opercular Area 3 $<0.001$ HC 0.577
-Area 31pd VentroMedial Visual Area 1 $<0.001$ HC 0.575
-Area Lateral Occipital 1 ProStriate Area $<0.001$ HC 0.575
-Insular Granular Complex ProStriate Area $<0.001$ HC 0.574
-Primary Auditory Cortex ProStriate Area $<0.001$ HC 0.574
-Area 5L ParaHippocampal Area 3 $<0.001$ HC 0.572
-Area 47m ProStriate Area $<0.001$ HC 0.571
-Anterior IntraParietal Area ParaHippocampal Area 3 $<0.001$ HC 0.57
-Area Lateral Occipital 1 ParaHippocampal Area 3 $<0.001$ HC 0.569
-Area V4t ParaHippocampal Area 3 $<0.001$ HC 0.568
-Area dorsal 23 a+b ParaHippocampal Area 3 $<0.001$ HC 0.567
-Area 10v Frontal Opercular Area 4 $<0.001$ HC 0.567
-Area V3B Anterior Ventral Insular Area $<0.001$ HC 0.565
-Posterior InferoTemporal complex ParaHippocampal Area 3 $<0.001$ HC 0.564
-Area 31pd ParaHippocampal Area 3 $<0.001$ HC 0.563
-Area V3CD ProStriate Area $<0.001$ HC 0.562
-Area V3B Anterior Agranular Insula Complex $<0.001$ HC 0.562
-Area V3CD PeriSylvian Language Area $<0.001$ HC 0.562
-Middle Temporal Area VentroMedial Visual Area 1 $<0.001$ HC 0.561
-Medial Belt Complex ProStriate Area $<0.001$ HC 0.561
-Fourth Visual Area ProStriate Area $<0.001$ HC 0.561
-Area TemporoParietoOccipital Junction 2 VentroMedial Visual Area 1 $<0.001$ HC 0.56
-VentroMedial Visual Area 1 VentroMedial Visual Area 2 $<0.001$ HC 0.559
-Polar 10p Area V3B $<0.001$ HC 0.559
-Area 52 ProStriate Area $<0.001$ HC 0.559
-Third Visual Area Ventral Visual Complex $<0.001$ HC 0.557
-Area Lateral Occipital 1 PeriSylvian Language Area $<0.001$ HC 0.556
-Area 47l (47 lateral) ProStriate Area $<0.001$ HC 0.556
-Area 10v Frontal Opercular Area 3 $<0.001$ HC 0.555
-Area 10v Area 23c $<0.001$ HC 0.554
-Fourth Visual Area PeriSylvian Language Area $<0.001$ HC 0.553
-Area 11l Sixth Visual Area $<0.001$ HC 0.553
-Area OP4-PV ParaHippocampal Area 3 $<0.001$ HC 0.553
-Third Visual Area PeriSylvian Language Area $<0.001$ HC 0.552
-Lateral Area 7A ParaHippocampal Area 3 $<0.001$ HC 0.552
-Area 10r Middle Insular Area $<0.001$ HC 0.551
-Medial Area 7P ProStriate Area $<0.001$ HC 0.55
-Area PGs VentroMedial Visual Area 1 $<0.001$ HC 0.55
-Area Lateral IntraParietal dorsal Primary Visual Cortex $<0.001$ HC 0.55
-Area 47l (47 lateral) Anterior Ventral Insular Area $<0.001$ HC 0.549
-Area TF Area TF $<0.001$ HC 0.549
-Medial Area 7P ParaHippocampal Area 3 $<0.001$ HC 0.549
-Primary Visual Cortex ProStriate Area $<0.001$ HC 0.549
-PreCuneus Visual Area VentroMedial Visual Area 1 $<0.001$ HC 0.549
-Area 7PC ParaHippocampal Area 3 $<0.001$ HC 0.549
-Area Lateral Occipital 1 Anterior Ventral Insular Area $<0.001$ HC 0.548
-Area IntraParietal 0 ParaHippocampal Area 3 $<0.001$ HC 0.548
-Area Lateral Occipital 2 ParaHippocampal Area 3 $<0.001$ HC 0.547
-Area TE1 posterior Dorsal Transitional Visual Area $<0.001$ HC 0.547
-Area PGs ProStriate Area $<0.001$ HC 0.545
-Area 31a ParaHippocampal Area 3 $<0.001$ HC 0.545
-Area 10v VentroMedial Visual Area 1 $<0.001$ HC 0.545
-Area 23d ParaHippocampal Area 3 $<0.001$ HC 0.545
-Area FST ParaHippocampal Area 3 $<0.001$ HC 0.545
-Area TemporoParietoOccipital Junction 3 VentroMedial Visual Area 1 $<0.001$ HC 0.543
-Area 13l Frontal Opercular Area 4 $<0.001$ HC 0.543
-Area Lateral Occipital 1 VentroMedial Visual Area 1 $<0.001$ HC 0.543
-VentroMedial Visual Area 1 VentroMedial Visual Area 1 $<0.001$ HC 0.542
-Area V3CD Anterior Ventral Insular Area $<0.001$ HC 0.542
-Area V3CD ParaHippocampal Area 3 $<0.001$ HC 0.542
-Area Frontal Opercular 5 ProStriate Area $<0.001$ HC 0.541
-Area 31p ventral ParaHippocampal Area 3 $<0.001$ HC 0.541
-Area Lateral Occipital 3 Frontal Opercular Area 3 $<0.001$ HC 0.541
-Polar 10p Dorsal Transitional Visual Area $<0.001$ HC 0.541
+Area OP4-PV R Area 10v R HC $<0.001$ 0.564
+Primary Motor Cortex R Area 10v R HC $<0.001$ 0.561
+Primary Motor Cortex L Area 10v R HC $<0.001$ 0.56
+Medial Belt Complex R Area 10v R HC $<0.001$ 0.559
+Area OP1-SII R Area 10v R HC $<0.001$ 0.559
+Area PF Opercular R Area 10v R HC $<0.001$ 0.558
+Area 43 L Area 10v R HC $<0.001$ 0.557
+Area a24 R Ventral Area 6 L HC $<0.001$ 0.554
+Superior Temporal Visual Area R Area 10v R HC $<0.001$ 0.554
+Area Posterior 24 prime R Area 10v R HC $<0.001$ 0.554
+Area 1 R Area 10v R HC $<0.001$ 0.553
+PreSubiculum R Area 10v R HC $<0.001$ 0.551
+Anterior IntraParietal Area R Area 10v R HC $<0.001$ 0.55
+Area PFt R Area 10v R HC $<0.001$ 0.549
+Area posterior 24 L Ventral Area 6 L HC $<0.001$ 0.548
+Ventral Area 6 L Area 10v R HC $<0.001$ 0.547
+Area 2 R Area 10v R HC $<0.001$ 0.546
+Anterior Ventral Insular Area L Area 10v R HC $<0.001$ 0.545
+Area PFcm R Area 10v R HC $<0.001$ 0.544
+Frontal Opercular Area 3 L Area 10v R HC $<0.001$ 0.543
+Ventral IntraParietal Complex L Area 10v R HC $<0.001$ 0.541
+Area PFm Complex R Area 10v R HC $<0.001$ 0.536
+VentroMedial Visual Area 1 R Area 10v R HC $<0.001$ 0.536
+Area 7PC R Area 10v R HC $<0.001$ 0.536
+Area V3B R Area 10v R HC $<0.001$ 0.535
+Area posterior 24 L Area 10v R HC $<0.001$ 0.535
+RetroInsular Cortex R Area 10v R HC $<0.001$ 0.535
+Area TE1 anterior L Ventral Area 6 L HC $<0.001$ 0.534
+Ventral Area 24d R Area 10v R HC $<0.001$ 0.534
+Area 25 R Ventral Area 6 L HC $<0.001$ 0.534
+Area p32 prime R Area 10v R HC $<0.001$ 0.534
+Insular Granular Complex R Area 10v R HC $<0.001$ 0.534
+Area 43 R Area 10v R HC $<0.001$ 0.531
+Area 1 L Area 10v R HC $<0.001$ 0.53
+Area TA2 L Area 10v R HC $<0.001$ 0.53
+Primary Sensory Cortex R Area 10v R HC $<0.001$ 0.53
+Area PGi R Area 10v R HC $<0.001$ 0.53
+Para-Insular Area L Area 10v R HC $<0.001$ 0.529
+Frontal Eye Fields L Area 10v R HC $<0.001$ 0.529
+Area 46 R Area 10v R HC $<0.001$ 0.528
+Area OP2-3-VS R Area 10v R HC $<0.001$ 0.527
+Area 55b R Area 10v R HC $<0.001$ 0.527
+Area dorsal 32 R Ventral Area 6 L HC $<0.001$ 0.527
+VentroMedial Visual Area 2 R Area 10v R HC $<0.001$ 0.526
+Dorsal Transitional Visual Area R Area 10v R HC $<0.001$ 0.526
+Area dorsal 32 L Area 11l R HC $<0.001$ 0.525
+Frontal Opercular Area 2 R Area 10v R HC $<0.001$ 0.525
+Area OP2-3-VS L Area 10v R HC $<0.001$ 0.524
+Area 10r R Ventral Area 6 L HC $<0.001$ 0.524
+Area 10v L Ventral Area 6 L HC $<0.001$ 0.524
+Frontal Opercular Area 4 L Ventral Area 6 L HC $<0.001$ 0.524
+Rostral Area 6 L Area 10v R HC $<0.001$ 0.523
+Frontal Opercular Area 1 R Area 10v R HC $<0.001$ 0.523
+Area PF Complex R Area 10v R HC $<0.001$ 0.522
+Area STSv anterior L Ventral Area 6 L HC $<0.001$ 0.521
+Area s32 L Ventral Area 6 L HC $<0.001$ 0.52
+Area TemporoParietoOccipital Junction 1 L Area 10v R HC $<0.001$ 0.52
+Frontal Opercular Area 4 L Area 10v R HC $<0.001$ 0.52
+Area 13l R Area 10v R HC $<0.001$ 0.52
+Area 3a R Area 10v R HC $<0.001$ 0.519
+Area PGi R Primary Motor Cortex R HC $<0.001$ 0.518
+Dorsal Area 24d R Area 10v R HC $<0.001$ 0.518
+Area dorsal 32 L Ventral Area 6 L HC $<0.001$ 0.518
+Area Posterior 24 prime L Area 10v R HC $<0.001$ 0.518
+Medial Belt Complex R Primary Motor Cortex R HC $<0.001$ 0.518
+Area posterior 24 R Ventral Area 6 L HC $<0.001$ 0.517
+Area PF Opercular L Area 10v R HC $<0.001$ 0.517
+Area IntraParietal 2 R Area 10v R HC $<0.001$ 0.517
+Area 47m R Area 10v R HC $<0.001$ 0.517
+Area 25 L Ventral Area 6 L HC $<0.001$ 0.517
+RetroSplenial Complex R Area 10v R HC $<0.001$ 0.516
+Parieto-Occipital Sulcus Area 1 L Area 10v R HC $<0.001$ 0.516
+Sixth Visual Area R Area 10v R HC $<0.001$ 0.516
+Area Frontal Opercular 5 L Area 11l R HC $<0.001$ 0.516
+Orbital Frontal Complex L Ventral Area 6 L HC $<0.001$ 0.515
+ParaHippocampal Area 3 R Area 10v R HC $<0.001$ 0.515
+Area p32 R Ventral Area 6 L HC $<0.001$ 0.514
+Area 10d L Ventral Area 6 L HC $<0.001$ 0.514
+Area a24 L Ventral Area 6 L HC $<0.001$ 0.514
+Area 43 L Ventral Area 6 L HC $<0.001$ 0.514
+Frontal Opercular Area 3 R Area 10v R HC $<0.001$ 0.513
+Supplementary and Cingulate Eye Field R Area 10v R HC $<0.001$ 0.513
+Dorsal area 6 L Area 10v R HC $<0.001$ 0.513
+Area PFt L Area 10v R HC $<0.001$ 0.513
+Area V6A R Area 10v R HC $<0.001$ 0.513
+Insular Granular Complex L Area 10v R HC $<0.001$ 0.513
+Area PFcm L Area 10v R HC $<0.001$ 0.513
+Anterior 24 prime R Area 10v R HC $<0.001$ 0.513
+Anterior Ventral Insular Area L Ventral Area 6 L HC $<0.001$ 0.513
+ParaHippocampal Area 1 R Area 10v R HC $<0.001$ 0.512
+PeriSylvian Language Area L Area 10v R HC $<0.001$ 0.511
+RetroSplenial Complex R Area Posterior Insular 1 L HC $<0.001$ 0.511
+Area 9-46d R Area 10v R HC $<0.001$ 0.511
+Area posterior 24 L Area 11l R HC $<0.001$ 0.511
+VentroMedial Visual Area 3 R Area 10v R HC $<0.001$ 0.511
+Area 47s L Area 11l R HC $<0.001$ 0.511
+Parieto-Occipital Sulcus Area 2 R Area 10v R HC $<0.001$ 0.51
+Area TG dorsal L Area 10v R HC $<0.001$ 0.51
+Area 2 L Area 10v R HC $<0.001$ 0.51
+Supplementary and Cingulate Eye Field L Area 10v R HC $<0.001$ 0.51
 Supplementary Table 18: Top connectivity features in low frequency range of ABCD ADHD prediction
-Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
-Area TG dorsal Area PH $<0.001$ HC 0.684
-Area TE1 anterior Area PH $<0.001$ HC 0.669
-Area anterior 47r Area PH $<0.001$ HC 0.661
-Area TG Ventral Area PH $<0.001$ HC 0.65
-Primary Auditory Cortex Area $<0.001$ HC 0.646
-Perirhinal Ectorhinal Cortex Area PH $<0.001$ HC 0.628
-Primary Auditory Cortex Area V3CD $<0.001$ HC 0.625
-Area TE2 anterior Area PH $<0.001$ HC 0.62
-Orbital Frontal Complex Area PH $<0.001$ HC 0.614
-Primary Auditory Cortex Area anterior 10p $<0.001$ HC 0.61
-Primary Auditory Cortex Area PH $<0.001$ HC 0.602
-Area TG dorsal Area V4t $<0.001$ HC 0.602
-Dorsal Area 24d Area PH $<0.001$ HC 0.601
-Area 10d Area PH $<0.001$ HC 0.594
-ParaHippocampal Area 3 Area PH $<0.001$ HC 0.593
-Area anterior 47r Area V4t $<0.001$ HC 0.592
-Area Area PH $<0.001$ HC 0.585
-ProStriate Area Area anterior 47r $<0.001$ HC 0.585
-Area TE2 posterior Area V4t $<0.001$ HC 0.585
-Area TG dorsal Area $<0.001$ HC 0.579
-Area V4t Area PH $<0.001$ HC 0.579
-Area 45 Area $<0.001$ HC 0.577
-Area 47m Area $<0.001$ HC 0.577
-Frontal Opercular Area 2 Area $<0.001$ HC 0.576
-Primary Auditory Cortex Area anterior 47r $<0.001$ HC 0.575
-Area TG Ventral Area V4t $<0.001$ HC 0.568
-Polar 10p Area PH $<0.001$ HC 0.566
-ProStriate Area Area V4t $<0.001$ HC 0.564
-Area 47s Area PH $<0.001$ HC 0.564
-Area TG dorsal Fusiform Face Complex $<0.001$ HC 0.563
-Area anterior 47r Area $<0.001$ HC 0.558
-Area STGa Area PH $<0.001$ HC 0.558
-Area 10v Area PH $<0.001$ HC 0.557
-Primary Auditory Cortex Area $<0.001$ HC 0.557
-Frontal Opercular Area 3 Area $<0.001$ HC 0.553
-Area OP1-SII Area anterior 47r $<0.001$ HC 0.552
-Area TemporoParietoOccipital Junction 1 Area anterior 47r $<0.001$ HC 0.552
-Superior Temporal Visual Area Area anterior 47r $<0.001$ HC 0.552
-Area TE2 posterior Area PH $<0.001$ HC 0.55
-ProStriate Area Area $<0.001$ HC 0.547
-Fusiform Face Complex Area anterior 47r $<0.001$ HC 0.542
-Area STGa Area $<0.001$ HC 0.541
-Area posterior 47r Area PH $<0.001$ HC 0.54
-Area 25 Area PH $<0.001$ HC 0.54
-Auditory 4 Complex Area V3CD $<0.001$ HC 0.539
-Area 10d Area V4t $<0.001$ HC 0.539
-Primary Auditory Cortex Area V4t $<0.001$ HC 0.538
-Area anterior 47r Area anterior 47r $<0.001$ HC 0.538
-ParaHippocampal Area 3 Area anterior 47r $<0.001$ HC 0.536
-Dorsal area 6 Area PH $<0.001$ HC 0.535
-Area anterior 10p Area V4t $<0.001$ HC 0.535
-Area 47m Area PH $<0.001$ HC 0.535
-Area 8B Area PH $<0.001$ HC 0.535
-Primary Visual Cortex Area PH $<0.001$ HC 0.534
-Area STSv anterior Area PH $<0.001$ HC 0.534
-Area TF Area PH $<0.001$ HC 0.534
-Area V6A Area PH $<0.001$ HC 0.533
-Area TG dorsal Area V3CD $<0.001$ HC 0.533
-Middle Temporal Area Area PH $<0.001$ HC 0.533
-Area OP2-3-VS Area anterior 47r $<0.001$ HC 0.533
-Area 47s Area $<0.001$ HC 0.533
-Posterior OFC Complex Area PH $<0.001$ HC 0.532
-Perirhinal Ectorhinal Cortex Area V4t $<0.001$ HC 0.532
-Area 10v Area $<0.001$ HC 0.532
-Area posterior 10p Area PH $<0.001$ HC 0.531
-Area anterior 10p Area PH $<0.001$ HC 0.531
-Area 47l (47 lateral) Area PH $<0.001$ HC 0.53
-Area 9 Middle Area PH $<0.001$ HC 0.53
-Area STSd posterior Area PH $<0.001$ HC 0.53
-Area TE2 posterior Area anterior 47r $<0.001$ HC 0.529
-Area TE1 anterior Area V4t $<0.001$ HC 0.529
-Area PHT Area PH $<0.001$ HC 0.528
-Area PFt Area anterior 47r $<0.001$ HC 0.527
-Area TE1 Middle Area PH $<0.001$ HC 0.526
-Area STSd anterior Area PH $<0.001$ HC 0.526
-Area 47s Area $<0.001$ HC 0.525
-RetroSplenial Complex Area PH $<0.001$ HC 0.525
-ParaHippocampal Area 1 Area PH $<0.001$ HC 0.525
-Entorhinal Cortex Area PH $<0.001$ HC 0.524
-Supplementary and Cingulate Eye Field Area $<0.001$ HC 0.524
-Area PGp Area PH $<0.001$ HC 0.524
-Area TG dorsal Area $<0.001$ HC 0.524
-Dorsal area 6 Area $<0.001$ HC 0.524
-Area anterior 47r Fourth Visual Area $<0.001$ HC 0.523
-Area PFcm Area PH $<0.001$ HC 0.523
-Supplementary and Cingulate Eye Field Area V3CD $<0.001$ HC 0.523
-Area 10r Area PH $<0.001$ HC 0.522
-Area Area PH $<0.001$ HC 0.521
-Eighth Visual Area Area $<0.001$ HC 0.521
-Area TE2 posterior Area $<0.001$ HC 0.52
-VentroMedial Visual Area 1 Polar 10p $<0.001$ HC 0.519
-Ventral Visual Complex Area TG Ventral $<0.001$ HC 0.519
-Area 31p ventral Area PH $<0.001$ HC 0.519
-Fourth Visual Area Area PH $<0.001$ HC 0.519
-ParaHippocampal Area 3 Fourth Visual Area $<0.001$ HC 0.518
-Area 7m Area PH $<0.001$ HC 0.518
-Area p32 Area PH $<0.001$ HC 0.517
-Auditory 4 Complex Area $<0.001$ HC 0.517
-Area FST Area PH $<0.001$ HC 0.517
-Area anterior 9-46v Area PH $<0.001$ HC 0.516
+Region 1 Region 2 Disorder/HC Corrected p-value Cohen’s d
+PeriSylvian Language Area L Area OP2-3-VS R HC $<0.001$ 0.441
+Dorsal Area 24d L Area OP2-3-VS R HC $<0.001$ 0.441
+Area 31pd L Auditory 4 Complex R HC $<0.001$ 0.44
+Area STSv posterior R Auditory 4 Complex R HC $<0.001$ 0.44
+Area STSv posterior R Area OP2-3-VS R HC $<0.001$ 0.439
+Dorsal area 6 R Area OP2-3-VS R HC $<0.001$ 0.435
+Area TE1 Middle L Frontal Opercular Area 2 R HC $<0.001$ 0.435
+Area TE1 anterior R Auditory 4 Complex R HC $<0.001$ 0.434
+Area STSd anterior L Area OP2-3-VS R HC $<0.001$ 0.433
+Auditory 5 Complex R Area OP2-3-VS R HC $<0.001$ 0.431
+Superior Temporal Visual Area L Insular Granular Complex L HC $<0.001$ 0.43
+RetroInsular Cortex L Area OP2-3-VS R HC $<0.001$ 0.429
+Area TE1 anterior R Area OP2-3-VS R HC $<0.001$ 0.429
+Area STSd posterior L Area OP2-3-VS R HC $<0.001$ 0.428
+Area 23d L Frontal Opercular Area 2 R HC $<0.001$ 0.424
+Area TG dorsal L Insular Granular Complex L HC $<0.001$ 0.423
+Sixth Visual Area R Area OP2-3-VS R HC $<0.001$ 0.423
+PeriSylvian Language Area L Insular Granular Complex L HC $<0.001$ 0.423
+PeriSylvian Language Area L Frontal Opercular Area 2 R HC $<0.001$ 0.422
+Entorhinal Cortex R Area OP2-3-VS R HC $<0.001$ 0.422
+PeriSylvian Language Area L Auditory 4 Complex R HC $<0.001$ 0.422
+Area STSv anterior L Area OP2-3-VS R HC $<0.001$ 0.421
+Area s32 L Frontal Opercular Area 2 R HC $<0.001$ 0.421
+Area s32 L Area OP2-3-VS R HC $<0.001$ 0.421
+Area TE1 Middle L Area OP2-3-VS R HC $<0.001$ 0.42
+Area 8Av L Insular Granular Complex L HC $<0.001$ 0.42
+Area STSd anterior R Area OP2-3-VS R HC $<0.001$ 0.418
+Area PHT R Area OP2-3-VS R HC $<0.001$ 0.417
+VentroMedial Visual Area 3 L Insular Granular Complex L HC $<0.001$ 0.417
+Area V6A L Area OP2-3-VS R HC $<0.001$ 0.417
+Area 47l (47 lateral) R Area OP2-3-VS R HC $<0.001$ 0.416
+Dorsal area 6 L Area OP2-3-VS R HC $<0.001$ 0.415
+Area 8Ad L Insular Granular Complex L HC $<0.001$ 0.414
+Area 8C L Insular Granular Complex L HC $<0.001$ 0.414
+Area TE1 anterior L Auditory 4 Complex R HC $<0.001$ 0.413
+Area V6A L Frontal Opercular Area 2 R HC $<0.001$ 0.413
+Area TG dorsal L Auditory 4 Complex R HC $<0.001$ 0.413
+Auditory 5 Complex L Area OP2-3-VS R HC $<0.001$ 0.412
+Area 10d R Frontal Opercular Area 2 R HC $<0.001$ 0.412
+Area 47s R Frontal Opercular Area 2 R HC $<0.001$ 0.411
+Area TE1 Middle L Auditory 4 Complex R HC $<0.001$ 0.411
+Area TE1 posterior L Insular Granular Complex L HC $<0.001$ 0.411
+Hippocampus R Auditory 4 Complex R HC $<0.001$ 0.411
+Area TE1 Middle L Insular Granular Complex L HC $<0.001$ 0.41
+Area STSd posterior L Auditory 4 Complex R HC $<0.001$ 0.41
+Primary Auditory Cortex R Auditory 4 Complex R HC $<0.001$ 0.41
+Dorsal Area 24d L Auditory 4 Complex R HC $<0.001$ 0.41
+Area dorsal 32 L Area OP2-3-VS R HC $<0.001$ 0.41
+Area TE1 posterior R Area OP2-3-VS R HC $<0.001$ 0.409
+Dorsal area 6 R Insular Granular Complex L HC $<0.001$ 0.409
+Primary Motor Cortex R Auditory 4 Complex R HC $<0.001$ 0.409
+Area TemporoParietoOccipital Junction 1 R Frontal Opercular Area 2 R HC $<0.001$ 0.409
+Area TE1 anterior R Frontal Opercular Area 2 R HC $<0.001$ 0.409
+Area TE1 Middle R Area OP2-3-VS R HC $<0.001$ 0.408
+Dorsal area 6 R Frontal Opercular Area 2 R HC $<0.001$ 0.408
+Lateral Belt Complex R Auditory 4 Complex R HC $<0.001$ 0.408
+Area 31pd L Insular Granular Complex L HC $<0.001$ 0.408
+Superior 6-8 Transitional Area L Insular Granular Complex L HC $<0.001$ 0.408
+Lateral Belt Complex R Area OP2-3-VS R HC $<0.001$ 0.407
+Area TG dorsal R Auditory 4 Complex R HC $<0.001$ 0.407
+Area 31pd L Area OP2-3-VS R HC $<0.001$ 0.407
+Area dorsal 32 L Frontal Opercular Area 2 R HC $<0.001$ 0.407
+Area 10d R Area OP2-3-VS R HC $<0.001$ 0.407
+Area STSd anterior L Insular Granular Complex L HC $<0.001$ 0.406
+Area 8B Lateral L Insular Granular Complex L HC $<0.001$ 0.405
+Area 31pd R Auditory 4 Complex R HC $<0.001$ 0.405
+Area 8Ad L Area OP2-3-VS R HC $<0.001$ 0.405
+Pirform Cortex L Area OP2-3-VS R HC $<0.001$ 0.405
+Frontal Opercular Area 1 R Area OP2-3-VS R HC $<0.001$ 0.405
+Area TemporoParietoOccipital Junction 1 R Auditory 4 Complex R HC $<0.001$ 0.405
+Dorsal area 6 R Medial Belt Complex R HC $<0.001$ 0.405
+Area 8B Lateral L Area OP2-3-VS R HC $<0.001$ 0.404
+Area PGs L Area OP2-3-VS R HC $<0.001$ 0.404
+Area TemporoParietoOccipital Junction 1 R Area OP2-3-VS R HC $<0.001$ 0.404
+Area 6mp R Auditory 4 Complex R HC $<0.001$ 0.403
+Area IFJa R Auditory 4 Complex R HC $<0.001$ 0.403
+Area 45 L Insular Granular Complex L HC $<0.001$ 0.403
+Area 31pd R Insular Granular Complex L HC $<0.001$ 0.403
+Ventral IntraParietal Complex R Area OP2-3-VS R HC $<0.001$ 0.403
+Auditory 5 Complex L Insular Granular Complex L HC $<0.001$ 0.403
+Auditory 4 Complex L Area OP2-3-VS R HC $<0.001$ 0.403
+Area PFcm R Auditory 4 Complex R HC $<0.001$ 0.403
+Area 6 anterior L Area OP2-3-VS R HC $<0.001$ 0.403
+Area dorsal 32 L Insular Granular Complex L HC $<0.001$ 0.403
+Area 10d L Area OP2-3-VS R HC $<0.001$ 0.403
+Area 9 Posterior L Insular Granular Complex L HC $<0.001$ 0.402
+Middle Temporal Area R Area OP2-3-VS R HC $<0.001$ 0.402
+Area 8Ad L Frontal Opercular Area 2 R HC $<0.001$ 0.402
+Area TE1 Middle L Posterior Insular Area 2 R HC $<0.001$ 0.401
+Area STSd anterior R Auditory 4 Complex R HC $<0.001$ 0.401
+Area TG dorsal R Frontal Opercular Area 2 R HC $<0.001$ 0.401
+Area TE1 posterior R Frontal Opercular Area 2 R HC $<0.001$ 0.401
+Area 7m R Auditory 4 Complex R HC $<0.001$ 0.401
+Entorhinal Cortex R Auditory 4 Complex R HC $<0.001$ 0.401
+Area 47l (47 lateral) L Area OP2-3-VS R HC $<0.001$ 0.4
+PreCuneus Visual Area L Insular Granular Complex L HC $<0.001$ 0.4
+Area STGa R Area OP2-3-VS R HC $<0.001$ 0.4
+Superior Temporal Visual Area L Area OP2-3-VS R HC $<0.001$ 0.4
+Area TG dorsal L Area OP2-3-VS R HC $<0.001$ 0.4
+Medial Belt Complex R Area OP2-3-VS R HC $<0.001$ 0.399
 Supplementary Table 19: Top connectivity features in ultralow frequency range of ABCD ADHD prediction
-Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
-Area Lateral IntraParietal ventral VentroMedial Visual Area 3 $<0.001$ ADHD 1.109
-Area s32 ParaHippocampal Area 3 $<0.001$ ADHD 1.062
-Area TF Area 25 $<0.001$ ADHD 1.028
-Inferior 6-8 Transitional Area VentroMedial Visual Area 3 $<0.001$ ADHD 1.016
-Area TF Area s32 $<0.001$ ADHD 1.007
-Ventral Visual Complex Area V3B $<0.001$ ADHD 1.002
-Area TA2 VentroMedial Visual Area 3 $<0.001$ ADHD 1.002
-Medial Area 7A VentroMedial Visual Area 3 $<0.001$ ADHD 1.0
-Area 25 Area TF $<0.001$ ADHD 0.994
-Area IntraParietal 0 VentroMedial Visual Area 3 $<0.001$ ADHD 0.994
-Area 25 ParaHippocampal Area 3 $<0.001$ ADHD 0.993
-Premotor Eye Field VentroMedial Visual Area 3 $<0.001$ ADHD 0.99
-Area TF Orbital Frontal Complex $<0.001$ ADHD 0.99
-Frontal Eye Fields VentroMedial Visual Area 3 $<0.001$ ADHD 0.979
-Area 13l Area TF $<0.001$ ADHD 0.975
-Area Lateral IntraParietal ventral Ventral Visual Complex $<0.001$ ADHD 0.969
-Ventral Area 6 VentroMedial Visual Area 3 $<0.001$ ADHD 0.968
-Frontal Eye Fields Eighth Visual Area $<0.001$ ADHD 0.968
-Area 10r Area TF $<0.001$ ADHD 0.963
-Medial Area 7A Eighth Visual Area $<0.001$ ADHD 0.96
-Lateral Area 7P VentroMedial Visual Area 3 $<0.001$ ADHD 0.959
-Inferior 6-8 Transitional Area Ventral Visual Complex $<0.001$ ADHD 0.955
-Premotor Eye Field Ventral Visual Complex $<0.001$ ADHD 0.955
-Area posterior 9-46v VentroMedial Visual Area 3 $<0.001$ ADHD 0.955
-Medial IntraParietal Area VentroMedial Visual Area 3 $<0.001$ ADHD 0.954
-Area 10r ParaHippocampal Area 3 $<0.001$ ADHD 0.953
-Orbital Frontal Complex Area TF $<0.001$ ADHD 0.948
-Frontal Eye Fields Posterior InferoTemporal complex $<0.001$ ADHD 0.948
-Ventral IntraParietal Complex VentroMedial Visual Area 3 $<0.001$ ADHD 0.945
-Area a24 Area TF $<0.001$ ADHD 0.944
-Area IFSp VentroMedial Visual Area 3 $<0.001$ ADHD 0.939
-Area 6 anterior VentroMedial Visual Area 3 $<0.001$ ADHD 0.938
-Frontal Eye Fields Ventral Visual Complex $<0.001$ ADHD 0.934
-Area 8BM Area TF $<0.001$ ADHD 0.934
-Area IntraParietal 1 Area TF $<0.001$ ADHD 0.934
-Area s32 Area TF $<0.001$ ADHD 0.934
-Posterior InferoTemporal complex Orbital Frontal Complex $<0.001$ ADHD 0.933
-Area V3B VentroMedial Visual Area 3 $<0.001$ ADHD 0.932
-IntraParietal Sulcus Area 1 VentroMedial Visual Area 3 $<0.001$ ADHD 0.932
-Area TF Area V3B $<0.001$ ADHD 0.932
-Area IFJa Posterior InferoTemporal complex $<0.001$ ADHD 0.93
-IntraParietal Sulcus Area 1 Ventral Visual Complex $<0.001$ ADHD 0.927
-Area a24 ParaHippocampal Area 3 $<0.001$ ADHD 0.927
-Area 6 anterior Ventral Visual Complex $<0.001$ ADHD 0.926
-Area 46 Eighth Visual Area $<0.001$ ADHD 0.925
-Area Lateral IntraParietal ventral Fourth Visual Area $<0.001$ ADHD 0.923
-Inferior 6-8 Transitional Area ParaHippocampal Area 3 $<0.001$ ADHD 0.923
-Area IFJp VentroMedial Visual Area 3 $<0.001$ ADHD 0.921
-Lateral Area 7P Ventral Visual Complex $<0.001$ ADHD 0.92
-Inferior 6-8 Transitional Area VentroMedial Visual Area 2 $<0.001$ ADHD 0.92
-Area dorsal 32 Area TF $<0.001$ ADHD 0.92
-Area TF Area a24 $<0.001$ ADHD 0.919
-Area IntraParietal 1 VentroMedial Visual Area 3 $<0.001$ ADHD 0.918
-Posterior OFC Complex Area TF $<0.001$ ADHD 0.917
-Inferior 6-8 Transitional Area Eighth Visual Area $<0.001$ ADHD 0.917
-Area 10r Posterior InferoTemporal complex $<0.001$ ADHD 0.916
-Area IFSp Posterior InferoTemporal complex $<0.001$ ADHD 0.916
-Area 6 anterior Area Lateral Occipital 2 $<0.001$ ADHD 0.916
-Anterior 24 prime VentroMedial Visual Area 3 $<0.001$ ADHD 0.913
-Area IFJa Area TF $<0.001$ ADHD 0.912
-Area s32 Posterior InferoTemporal complex $<0.001$ ADHD 0.911
-Area IntraParietal 0 ParaHippocampal Area 3 $<0.001$ ADHD 0.91
-Area 55b VentroMedial Visual Area 3 $<0.001$ ADHD 0.907
-Inferior 6-8 Transitional Area Fourth Visual Area $<0.001$ ADHD 0.907
-Anterior Ventral Insular Area Posterior InferoTemporal complex $<0.001$ ADHD 0.906
-Area Lateral IntraParietal ventral Eighth Visual Area $<0.001$ ADHD 0.906
-Area p32 prime VentroMedial Visual Area 3 $<0.001$ ADHD 0.905
-Medial IntraParietal Area Ventral Visual Complex $<0.001$ ADHD 0.904
-Area Lateral IntraParietal ventral VentroMedial Visual Area 1 $<0.001$ ADHD 0.904
-Area p32 Area TF $<0.001$ ADHD 0.903
-Area TA2 Eighth Visual Area $<0.001$ ADHD 0.902
-Area 8Ad Area TF $<0.001$ ADHD 0.902
-Area Lateral IntraParietal ventral VentroMedial Visual Area 2 $<0.001$ ADHD 0.902
-Area 6 anterior Eighth Visual Area $<0.001$ ADHD 0.901
-Premotor Eye Field Eighth Visual Area $<0.001$ ADHD 0.899
-Area STSv anterior Area TF $<0.001$ ADHD 0.899
-Medial Area 7A Ventral Visual Complex $<0.001$ ADHD 0.897
-Posterior OFC Complex Posterior InferoTemporal complex $<0.001$ ADHD 0.896
-Area 55b Posterior InferoTemporal complex $<0.001$ ADHD 0.895
-Area 25 Posterior InferoTemporal complex $<0.001$ ADHD 0.894
-Area PF Opercular VentroMedial Visual Area 3 $<0.001$ ADHD 0.893
-Area IFJp VentroMedial Visual Area 1 $<0.001$ ADHD 0.891
-Ventral IntraParietal Complex Ventral Visual Complex $<0.001$ ADHD 0.891
-Anterior Agranular Insula Complex VentroMedial Visual Area 3 $<0.001$ ADHD 0.891
-Area 55b Ventral Visual Complex $<0.001$ ADHD 0.89
-Entorhinal Cortex ParaHippocampal Area 3 $<0.001$ ADHD 0.889
-Area 46 VentroMedial Visual Area 3 $<0.001$ ADHD 0.889
-Area STSd anterior Area TF $<0.001$ ADHD 0.887
-Area TF Area IntraParietal 2 $<0.001$ ADHD 0.887
-Area 6 anterior ParaHippocampal Area 3 $<0.001$ ADHD 0.887
-Area IFJp Posterior InferoTemporal complex $<0.001$ ADHD 0.886
-Frontal Opercular Area 4 VentroMedial Visual Area 3 $<0.001$ ADHD 0.886
-Area IFSp Ventral Visual Complex $<0.001$ ADHD 0.884
-Middle Insular Area VentroMedial Visual Area 3 $<0.001$ ADHD 0.883
-Medial IntraParietal Area Area TF $<0.001$ ADHD 0.881
-Ventral Visual Complex Frontal Eye Fields $<0.001$ ADHD 0.881
-Area 7m Area TF $<0.001$ ADHD 0.879
-Area 6m anterior Eighth Visual Area $<0.001$ ADHD 0.879
-Area Lateral IntraParietal dorsal Ventral Visual Complex $<0.001$ ADHD 0.879
-Area TA2 Posterior InferoTemporal complex $<0.001$ ADHD 0.878
+Region 1 Region 2 Disorder/HC Corrected p-value Cohen’s d
+Medial Belt Complex R Area PGp R HC $<0.001$ 0.373
+Area dorsal 32 R ParaHippocampal Area 1 R ADHD $<0.001$ 0.372
+Area FST R ParaHippocampal Area 1 R ADHD $<0.001$ 0.368
+Area dorsal 32 R Area OP2-3-VS R ADHD $<0.001$ 0.367
+Medial Belt Complex R Seventh Visual Area R HC $<0.001$ 0.363
+Area PGi L ParaHippocampal Area 1 R ADHD $<0.001$ 0.361
+ParaHippocampal Area 3 R Area TF R HC 0.001 0.346
+Area 47s R Area PGp R HC 0.001 0.344
+Area 8BM R Area OP2-3-VS R ADHD 0.001 0.343
+Area FST R Area OP2-3-VS R ADHD 0.001 0.343
+Area 11l L Area PGp R HC 0.001 0.343
+Area 8Av R ParaHippocampal Area 1 R ADHD 0.001 0.343
+Area 11l L Anterior IntraParietal Area L HC 0.001 0.342
+Area 47l (47 lateral) L Anterior IntraParietal Area L HC 0.001 0.341
+Area PGi L Area OP2-3-VS R ADHD 0.001 0.339
+Area 8BM R ParaHippocampal Area 1 R ADHD 0.001 0.339
+Area FST R Area Posterior Insular 1 L ADHD 0.001 0.338
+Area TemporoParietoOccipital Junction 2 L ParaHippocampal Area 1 R ADHD 0.001 0.337
+Area 47s R Anterior IntraParietal Area L HC 0.001 0.336
+Area OP4-PV R Area PGp R HC 0.001 0.336
+Pirform Cortex R Area PGp R HC 0.001 0.336
+Dorsal Transitional Visual Area R ParaHippocampal Area 1 R ADHD 0.001 0.335
+Area dorsal 23 a+b R ParaHippocampal Area 1 R ADHD 0.001 0.334
+Fourth Visual Area L Area OP2-3-VS R ADHD 0.001 0.334
+Area STSv posterior R Area 47s R ADHD 0.001 0.333
+Area Posterior Insular 1 L Medial Superior Temporal Area L HC 0.001 0.33
+Area 47s R Seventh Visual Area R HC 0.001 0.33
+Area STSv posterior R Polar 10p R ADHD 0.001 0.33
+Area 11l L Seventh Visual Area R HC 0.001 0.33
+Area OP4-PV R Area TF R HC 0.001 0.329
+Medial Belt Complex R Sixth Visual Area L HC 0.001 0.328
+Area STSv posterior R Anterior Agranular Insula Complex R ADHD 0.002 0.328
+Area STSv posterior R ParaHippocampal Area 1 R ADHD 0.002 0.328
+Area 5L R ParaHippocampal Area 1 R ADHD 0.002 0.328
+Area IFSa R Area OP2-3-VS R ADHD 0.002 0.327
+Medial Belt Complex R Medial Superior Temporal Area L HC 0.002 0.326
+Area STGa R Area TF R HC 0.002 0.325
+Area 11l L Area 7PC L HC 0.002 0.325
+Area V3A L Area OP2-3-VS R ADHD 0.002 0.324
+Area posterior 9-46v R Area OP2-3-VS R ADHD 0.002 0.324
+Area anterior 9-46v R Area OP2-3-VS R ADHD 0.002 0.324
+Fourth Visual Area R Area 10v L ADHD 0.002 0.323
+Medial Belt Complex R Anterior IntraParietal Area L HC 0.002 0.323
+Area STSv posterior R Orbital Frontal Complex L ADHD 0.002 0.323
+Area PFm Complex L Area OP2-3-VS R ADHD 0.002 0.322
+Area STSv posterior R Area OP2-3-VS R ADHD 0.002 0.322
+Medial Area 7A R ParaHippocampal Area 1 R ADHD 0.002 0.321
+Medial Belt Complex R Area V3A L HC 0.002 0.32
+Area OP4-PV R Medial Superior Temporal Area L HC 0.002 0.32
+Area STSv posterior R Area 10v L ADHD 0.002 0.32
+Area 11l L Fourth Visual Area L HC 0.002 0.32
+Area STSd posterior L ParaHippocampal Area 1 R ADHD 0.002 0.319
+Area dorsal 32 R Area Posterior 24 prime R ADHD 0.002 0.319
+Area dorsal 32 R Hippocampus R ADHD 0.002 0.319
+Area Posterior Insular 1 L Anterior IntraParietal Area L HC 0.002 0.318
+Area PFm Complex L ParaHippocampal Area 1 R ADHD 0.002 0.318
+Third Visual Area R Area OP2-3-VS R ADHD 0.002 0.318
+Area FST L ParaHippocampal Area 1 R ADHD 0.002 0.317
+Area FST R Pirform Cortex L ADHD 0.002 0.317
+Area FST R Area p32 L ADHD 0.002 0.317
+Anterior IntraParietal Area L ParaHippocampal Area 1 R ADHD 0.002 0.316
+Area 11l L Ventral IntraParietal Complex L HC 0.002 0.316
+Area STSv posterior R Area 47l (47 lateral) R ADHD 0.002 0.316
+Area STSd posterior L Area OP2-3-VS R ADHD 0.002 0.316
+Area ventral 23 a+b R ParaHippocampal Area 1 R ADHD 0.002 0.315
+Area STSv posterior R Pirform Cortex L ADHD 0.002 0.315
+Medial Belt Complex R Second Visual Area L HC 0.002 0.315
+Area Posterior Insular 1 L Area PGp R HC 0.002 0.315
+Area 47l (47 lateral) L Medial Superior Temporal Area L HC 0.002 0.315
+Area 8Av R Area OP2-3-VS R ADHD 0.002 0.314
+Pirform Cortex R Seventh Visual Area R HC 0.002 0.314
+Sixth Visual Area L ParaHippocampal Area 1 R ADHD 0.002 0.314
+Area 43 L Area OP2-3-VS R ADHD 0.002 0.313
+Area STSv posterior R Area p32 L ADHD 0.002 0.313
+Area TG dorsal R Area PGp R HC 0.002 0.313
+Medial Superior Temporal Area L ParaHippocampal Area 1 R ADHD 0.003 0.312
+Area dorsal 32 R Anterior Agranular Insula Complex R ADHD 0.002 0.312
+Area p32 R ParaHippocampal Area 1 R ADHD 0.003 0.312
+Area PGi L Hippocampus R ADHD 0.003 0.312
+Area STGa L Anterior IntraParietal Area L HC 0.002 0.312
+Medial Area 7A R Area OP2-3-VS R ADHD 0.003 0.311
+Dorsal area 6 R Medial Superior Temporal Area L HC 0.003 0.311
+Area TE1 Middle R ParaHippocampal Area 1 R ADHD 0.003 0.311
+Area p32 L Medial Superior Temporal Area L HC 0.003 0.31
+IntraParietal Sulcus Area 1 R ParaHippocampal Area 1 R ADHD 0.003 0.31
+Fusiform Face Complex R ParaHippocampal Area 1 R ADHD 0.003 0.309
+Area 11l L Medial Superior Temporal Area L HC 0.003 0.309
+Area 44 R ParaHippocampal Area 1 R ADHD 0.003 0.309
+Area 23c R ParaHippocampal Area 1 R ADHD 0.003 0.309
+Medial Belt Complex R Area V6A R HC 0.003 0.309
+Parieto-Occipital Sulcus Area 1 R Anterior IntraParietal Area L HC 0.003 0.308
+Medial Belt Complex R Ventral IntraParietal Complex L HC 0.003 0.308
+Dorsal area 6 R Seventh Visual Area R HC 0.003 0.308
+Area 47l (47 lateral) L Area PGp R HC 0.003 0.308
+Area STSv posterior R Area anterior 10p L ADHD 0.003 0.307
+Ventral Area 6 L Medial Superior Temporal Area L HC 0.003 0.306
+Area TemporoParietoOccipital Junction 1 R ParaHippocampal Area 1 R ADHD 0.003 0.306
+Pirform Cortex L Area PGp R HC 0.003 0.306
+Area STSv posterior R Area 47s L ADHD 0.003 0.305
+Parieto-Occipital Sulcus Area 2 L ParaHippocampal Area 1 R ADHD 0.003 0.305
 Supplementary Table 20: Top connectivity features in high frequency range of ABIDE ASD prediction
-Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
-Dorsal Area 24d Area PGs $<0.001$ ASD 6.950
-Area p32 Area PGs $<0.001$ ASD 6.565
-Ventral Area 6 Area 8 Anterior ventral $<0.001$ ASD 6.479
-Area p32 Area 8 Anterior ventral $<0.001$ ASD 6.313
-Area 9 Middle Area 8 Anterior ventral $<0.001$ ASD 6.135
-Area 9-46 dorsal Area PGs $<0.001$ ASD 6.057
-Area ventral 23 a+b Area 8 Anterior ventral $<0.001$ ASD 6.041
-Supplementary and Cingulate Eye Field Area PGs $<0.001$ ASD 5.959
-Area 31pd Area 8 Anterior ventral $<0.001$ ASD 5.899
-ParaHippocampal Area 3 Area 8Ad $<0.001$ ASD 5.819
-Area 31 anterior Area PGs $<0.001$ ASD 5.671
-Orbital Frontal Complex Area 47 medial $<0.001$ ASD 5.531
-Area a24 Area 8 Anterior ventral $<0.001$ ASD 5.509
-PreSubiculum Area PGs $<0.001$ ASD 5.178
-Area 31p ventral Area 8 Anterior ventral $<0.001$ ASD 5.171
-Area 23 dorsal Area PGs $<0.001$ ASD 5.107
-Area V3A Area 47 medial $<0.001$ ASD 4.942
-Area STSd anterior Area PGs $<0.001$ ASD 4.881
-Area TemporoParietoOccipital Junction 3 Area 8Ad $<0.001$ ASD 4.745
-Area 46 Area 47 medial $0.001$ ASD 4.683
-Area IFSp Area 47 medial $0.001$ ASD 4.669
-Area 10 rostral Area 8 Anterior ventral $0.001$ ASD 4.589
-Area posterior 24 Area TE1 Middle $0.001$ ASD 4.574
-Area posterior 24 Area 8 Anterior ventral $0.001$ ASD 4.550
-Primary Visual Cortex Area 47 medial $0.001$ ASD 4.533
-Area Posterior 24 prime Area 47 medial $0.001$ ASD 4.512
-Premotor Eye Field Area 47 medial $0.001$ ASD 4.507
-Ventral Area 6 Area PGs $0.001$ ASD 4.502
-Inferior 6-8 Transitional Area Area 47 medial $0.001$ ASD 4.456
-PreSubiculum Area 8 Anterior ventral $<0.001$ ASD 4.430
-Area 9-46 dorsal Area 8 Anterior ventral $<0.001$ ASD 4.377
-VentroMedial Visual Area 1 Area 47 medial $0.001$ ASD 4.365
-Area 33 prime Area 47 medial $0.001$ ASD 4.293
-Second Visual Area Area 47 medial $0.001$ ASD 4.274
-Para-Insular Area Area 8 Anterior ventral $0.001$ ASD 4.227
-ParaHippocampal Area 3 Area 8 Anterior ventral $0.001$ ASD 4.222
-Area PFm Complex Area 8 Anterior ventral $0.001$ ASD 4.222
-Third Visual Area Area 47 medial $0.001$ ASD 4.206
-Area Lateral Occipital 3 Area 47 medial $0.001$ ASD 4.196
-Area PGi Area 8 Anterior ventral $0.001$ ASD 4.193
-Area p32 prime Area PGs $0.001$ ASD 4.178
-Area 33 prime Area 8 Anterior ventral $0.001$ ASD 4.176
-Superior Frontal Language Area Area 8 Anterior ventral $<0.001$ ASD 4.176
-Supplementary and Cingulate Eye Field Area 8 Anterior ventral $0.001$ ASD 4.166
-Middle Insular Area Area PGs $0.001$ ASD 4.161
-Area dorsal 32 Area 8 Anterior ventral $<0.001$ ASD 4.079
-Area anterior 32 prime Area 8 Anterior ventral $0.001$ ASD 4.073
-Area posterior 24 Area 47 medial $0.001$ ASD 4.065
-Area Lateral IntraParietal ventral Area 47 medial $0.002$ ASD 4.046
-Area IFJp Area 47 medial $0.001$ ASD 4.034
-Anterior 24 prime Area 47 medial $0.002$ ASD 4.026
-Inferior 6-8 Transitional Area Area TE1 Middle $0.002$ ASD 4.011
-Area 11 lateral Area 47 medial $0.002$ ASD 3.987
-ParaHippocampal Area 3 Area 47l (47 lateral) $0.001$ ASD 3.980
-Area IntraParietal 1 Area 8 Anterior ventral $0.002$ ASD 3.980
-Superior 6-8 Transitional Area Area TE1 Middle $0.001$ ASD 3.968
-Area PH Area 47 medial $<0.001$ ASD 3.963
-Area Lateral Occipital 2 Area 8 Anterior ventral $0.001$ ASD 3.956
-Eighth Visual Area Inferior 6-8 Transitional Area $0.002$ ASD 3.952
-Area 6 anterior Area 47 medial $0.002$ ASD 3.948
-Para-Insular Area Area 47 medial $0.002$ ASD 3.918
-Area s32 Area PGs $0.002$ ASD 3.913
-Area OP1-SII Inferior 6-8 Transitional Area $0.002$ ASD 3.897
-Area a24 Area PGs $0.002$ ASD 3.895
-Area TE1 Middle Area 8 Anterior ventral $0.001$ ASD 3.895
-Area Lateral Occipital 2 Area 47 medial $0.002$ ASD 3.875
-Area 6 medial anterior Area TE1 Middle $0.002$ ASD 3.866
-Area dorsal 23 a+b Area 8 Anterior ventral $0.002$ ASD 3.858
-Supplementary and Cingulate Eye Field Area TE1 Middle $0.002$ ASD 3.843
-Area TemporoParietoOccipital Junction 2 Area 8 Anterior ventral $0.002$ ASD 3.840
-Area posterior 9-46 ventral Area 47 medial $<0.001$ ASD 3.825
-Pirform Cortex Area 47 medial $0.002$ ASD 3.804
-Area Posterior Insular 1 Area 47 medial $0.002$ ASD 3.800
-Fourth Visual Area Area 47 medial $0.002$ ASD 3.798
-Area a24 Dorsal area 6 $0.002$ ASD 3.789
-Area anterior 10p Area TE1 Middle $0.002$ ASD 3.788
-Anterior 24 prime Area TE1 Middle $0.002$ ASD 3.785
-Visual area 3-B Area 47 medial $0.002$ ASD 3.757
-Area s32 Dorsal area 6 $0.002$ ASD 3.752
-Area TemporoParietoOccipital Junction 3 Area 8 Anterior ventral $0.002$ ASD 3.744
-RetroInsular Cortex Area 8 Anterior ventral $0.002$ ASD 3.736
-Area 46 Area 8 Anterior ventral $<0.001$ ASD 3.731
-Perirhinal Ectorhinal Cortex Area 47 medial $0.002$ ASD 3.731
-Area p32 prime Area 47 medial $0.002$ ASD 3.720
-Lateral Belt Complex Area 47 medial $0.002$ ASD 3.702
-Anterior 24 prime Area 8 Anterior ventral $0.002$ ASD 3.698
-Area OP1-SII Primary Motor Cortex $0.002$ ASD 3.694
-Area 6 medial anterior Superior 6-8 Transitional Area $0.002$ ASD 3.693
-Area Lateral IntraParietal dorsal Medial Area 7P (posterior) $0.002$ ASD 3.693
-Area p32 prime Area 8 Anterior ventral $0.002$ ASD 3.682
-Premotor Eye Field Area 8 Anterior ventral $0.002$ ASD 3.678
-Entorhinal Cortex Area 47 medial $0.003$ ASD 3.654
-Area 31 anterior Area 8 Anterior ventral $0.002$ ASD 3.629
-Area 7m Area 8 Anterior ventral $0.002$ ASD 3.625
-Area dorsal 32 Area TG dorsal $0.003$ ASD 3.608
-Area Posterior 24 prime Area 8 Anterior ventral $0.002$ ASD 3.602
-Area IntraParietal 1 Medial Area 7P (posterior) $0.002$ ASD 3.593
-Area dorsal 23 a+b Inferior 6-8 Transitional Area $0.003$ ASD 3.579
-Ventral Area 6 Area 47 medial $0.003$ ASD 3.574
-Area 31pd Area PGs $0.001$ ASD 3.556
+Region 1 Region 2 Disorder/HC Corrected p-value Cohen’s d
+Area 2 L Area 11l R HC 0.002 2.232
+Area PFcm L Area 11l R HC 0.003 2.191
+RetroInsular Cortex L Area 11l R HC 0.003 2.162
+Ventral Area 24d L Area 11l R HC 0.003 2.162
+Area IntraParietal 2 L Area 11l R HC 0.003 2.149
+Area Posterior 24 prime R Area 11l R HC 0.003 2.147
+Area Posterior 24 prime L Area 11l R HC 0.003 2.147
+Ventral Area 24d R Area 11l R HC 0.003 2.145
+Area 1 L Area 11l R HC 0.003 2.141
+Area 44 R Area 11l R HC 0.003 2.14
+Primary Auditory Cortex L Area 11l R HC 0.003 2.127
+Area dorsal 32 R Area 11l R HC 0.004 2.11
+Area Frontal Opercular 5 L Area 11l R HC 0.004 2.106
+Area anterior 32 prime L Area 11l R HC 0.004 2.104
+Parieto-Occipital Sulcus Area 2 R Area 11l R HC 0.004 2.101
+Anterior IntraParietal Area L Area 11l R HC 0.004 2.101
+Anterior 24 prime R Area 11l R HC 0.004 2.099
+Area 43 R Area 11l R HC 0.004 2.097
+Orbital Frontal Complex R Area 11l R HC 0.004 2.091
+Area OP4-PV R Area 11l R HC 0.004 2.09
+Area 23d L Area 11l R HC 0.004 2.085
+Frontal Opercular Area 3 L Area 11l R HC 0.004 2.083
+Area PGi R Area 11l R HC 0.004 2.082
+Dorsal Area 24d L Area 11l R HC 0.004 2.082
+Primary Sensory Cortex L Area 11l R HC 0.004 2.08
+Second Visual Area R Sixth Visual Area L ASD 0.004 2.075
+Area PFt L Area 11l R HC 0.004 2.075
+Area 7PC L Area 11l R HC 0.004 2.073
+Area anterior 32 prime R Area 11l R HC 0.004 2.07
+Area 33 prime L Area 11l R HC 0.004 2.067
+Area OP2-3-VS L Area 11l R HC 0.004 2.066
+Area posterior 10p R Area 11l R HC 0.004 2.064
+Dorsal Transitional Visual Area R Area 11l R HC 0.004 2.064
+Area ventral 23 a+b L Area 11l R HC 0.004 2.055
+Area 23d R Area 11l R HC 0.004 2.054
+Area 7m R Area 11l R HC 0.004 2.053
+Area 10v R Area 11l R HC 0.004 2.05
+Lateral Area 7A L Area 11l R HC 0.004 2.05
+Area IntraParietal 1 L Area 11l R HC 0.004 2.049
+Area Frontal Opercular 5 R Area 11l R HC 0.004 2.049
+Anterior 24 prime L Area 11l R HC 0.004 2.049
+Area posterior 9-46v L Area 11l R HC 0.004 2.046
+Area p32 prime R Area 11l R HC 0.004 2.046
+Area ventral 23 a+b R Area 11l R HC 0.004 2.045
+Area posterior 24 L Area 11l R HC 0.004 2.045
+Area dorsal 32 L Area 11l R HC 0.004 2.044
+PeriSylvian Language Area L Area 11l R HC 0.004 2.044
+Auditory 4 Complex R Area 11l R HC 0.004 2.044
+Lateral Belt Complex R Area 11l R HC 0.004 2.043
+Area 8BM R Area 11l R HC 0.004 2.043
+Area PGs L Area 11l R HC 0.004 2.041
+Area 10d R Area 11l R HC 0.004 2.04
+Fourth Visual Area L Sixth Visual Area L ASD 0.005 2.039
+Rostral Area 6 R Area 11l R HC 0.005 2.038
+Area 33 prime R Area 11l R HC 0.005 2.035
+Insular Granular Complex L Area 11l R HC 0.005 2.034
+Ventral Area 6 R Area 11l R HC 0.005 2.034
+Seventh Visual Area R Sixth Visual Area L ASD 0.005 2.032
+Area dorsal 23 a+b L Area 11l R HC 0.005 2.027
+Orbital Frontal Complex L Area 11l R HC 0.005 2.027
+Area TE2 anterior L Area 11l R HC 0.005 2.027
+Medial Belt Complex R Area 11l R HC 0.005 2.026
+Area 45 R Area 11l R HC 0.005 2.025
+Medial Area 7P R Area 11l R HC 0.005 2.023
+Ventral Area 6 L Area 11l R HC 0.005 2.021
+Dorsal Area 24d R Area 11l R HC 0.005 2.02
+ProStriate Area L Area 11l R HC 0.005 2.02
+Area OP1-SII L Area 11l R HC 0.005 2.02
+Medial IntraParietal Area L Area 11l R HC 0.005 2.019
+Area anterior 47r R Area 11l R HC 0.005 2.018
+Superior 6-8 Transitional Area R Area 11l R HC 0.005 2.015
+Area 8Ad R Area 11l R HC 0.005 2.012
+Area IFSp L Area 11l R HC 0.005 2.012
+Superior Temporal Visual Area L Area 11l R HC 0.005 2.011
+Ventral IntraParietal Complex L Area 11l R HC 0.005 2.011
+Area p32 prime L Area 11l R HC 0.005 2.01
+PreSubiculum R Sixth Visual Area L ASD 0.005 2.007
+Area 8B Lateral R Area 11l R HC 0.005 2.005
+Area V6A L Area 11l R HC 0.005 2.004
+Area IFJp L Area 11l R HC 0.005 2.004
+Area Lateral IntraParietal ventral L Area 11l R HC 0.005 2.004
+Area 8Av R Area 11l R HC 0.005 2.004
+Area IFJa L Area 11l R HC 0.005 2.004
+Parieto-Occipital Sulcus Area 2 L Area 11l R HC 0.005 2.003
+Frontal Opercular Area 4 L Area 11l R HC 0.005 2.003
+Area PF Opercular L Area 11l R HC 0.005 2.001
+Area posterior 24 R Area 11l R HC 0.005 2.001
+Area PFm Complex R Area 11l R HC 0.005 2.0
+Area 5L R Area 11l R HC 0.005 2.0
+Area 10v L Area 11l R HC 0.005 1.998
+Area 5m ventral L Area 11l R HC 0.005 1.998
+Anterior Ventral Insular Area L Area 11l R HC 0.005 1.996
+Area posterior 9-46v R Area 11l R HC 0.005 1.994
+Area PGp L Area 11l R HC 0.005 1.992
+Frontal Opercular Area 2 L Area 11l R HC 0.005 1.991
+Area PFm Complex L Area 11l R HC 0.005 1.99
+Area Lateral IntraParietal dorsal L Area 11l R HC 0.005 1.988
+Parieto-Occipital Sulcus Area 1 L Area 11l R HC 0.005 1.988
+Area 31p ventral L Area 11l R HC 0.005 1.987
+Area V6A R Area 11l R HC 0.005 1.987
 Supplementary Table 21: Top connectivity features in low frequency range of ABIDE ASD prediction
 Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
-Area PHT Area posterior 9-46 ventral $0.028$ ASD 2.155
-ParaBelt Complex Area posterior 9-46 ventral $0.029$ ASD 2.154
-Auditory 5 Complex Area posterior 9-46 ventral $0.030$ ASD 2.123
-Area 9 Posterior Area posterior 9-46 ventral $0.030$ ASD 2.122
-Primary Sensory Cortex Area posterior 9-46 ventral $0.031$ ASD 2.116
-Area PH Area posterior 9-46 ventral $0.031$ ASD 2.115
-Area 3a Area posterior 9-46 ventral $0.031$ ASD 2.109
-Area 6mp Area posterior 9-46 ventral $0.032$ ASD 2.099
-Lateral Belt Complex Area posterior 9-46 ventral $0.033$ ASD 2.078
-Area TA2 Area posterior 9-46 ventral $0.034$ ASD 2.065
-Area TE2 posterior Area posterior 9-46 ventral $0.034$ ASD 2.057
-Area V3A Area posterior 9-46 ventral $0.034$ ASD 2.057
-Area PFm Complex Area posterior 9-46 ventral $0.034$ ASD 2.053
-Area Lateral Occipital 2 Area posterior 9-46 ventral $0.035$ ASD 2.046
-Dorsal area 6 Area posterior 9-46 ventral $0.035$ ASD 2.045
-VentroMedial Visual Area 1 Area posterior 9-46 ventral $0.035$ ASD 2.039
-ParaHippocampal Area 3 Area posterior 9-46 ventral $0.036$ ASD 2.038
-Area IFSp Area posterior 9-46 ventral $0.036$ ASD 2.037
-Primary Auditory Cortex Area posterior 9-46 ventral $0.036$ ASD 2.032
-Ventral Visual Complex Area posterior 9-46 ventral $0.037$ ASD 2.023
-Area 1 Area posterior 9-46 ventral $0.036$ ASD 2.018
-Area PGs Area posterior 9-46 ventral $0.037$ ASD 2.009
-Area IntraParietal 2 Area posterior 9-46 ventral $0.037$ ASD 2.007
-Area TemporoParietoOccipital Junction 1 Area posterior 9-46 ventral $0.037$ ASD 2.004
-Primary Motor Cortex Area posterior 9-46 ventral $0.038$ ASD 2.001
-ParaHippocampal Area 1 Area posterior 9-46 ventral $0.039$ ASD 1.999
-Auditory 4 Complex Area posterior 9-46 ventral $0.039$ ASD 1.993
-Area 31a Area posterior 9-46 ventral $0.039$ ASD 1.987
-Area TE1 posterior Area posterior 9-46 ventral $0.039$ ASD 1.985
-Superior Temporal Visual Area Area posterior 9-46 ventral $0.039$ ASD 1.980
-Area 25 Area posterior 9-46 ventral $0.040$ ASD 1.978
-Area V6A Area posterior 9-46 ventral $0.039$ ASD 1.977
-Area OP2-3-VS Area posterior 9-46 ventral $0.040$ ASD 1.975
-Ventral Area 24d Area posterior 9-46 ventral $0.041$ ASD 1.975
-Area 2 Area posterior 9-46 ventral $0.040$ ASD 1.963
-ProStriate Area Area posterior 9-46 ventral $0.041$ ASD 1.960
-RetroInsular Cortex Area posterior 9-46 ventral $0.041$ ASD 1.960
-Area 9 anterior Area posterior 9-46 ventral $0.041$ ASD 1.954
-VentroMedial Visual Area 3 Area posterior 9-46 ventral $0.042$ ASD 1.953
-ParaHippocampal Area 2 Area posterior 9-46 ventral $0.042$ ASD 1.947
-Area anterior 9-46 ventral Area posterior 9-46 ventral $0.043$ ASD 1.941
-Medial Belt Complex Area posterior 9-46 ventral $0.043$ ASD 1.940
-Insular Granular Complex Area posterior 9-46 ventral $0.043$ ASD 1.936
-Eighth Visual Area Area posterior 9-46 ventral $0.044$ ASD 1.933
-Superior Frontal Language Area Area posterior 9-46 ventral $0.044$ ASD 1.931
-Area TemporoParietoOccipital Junction 3 Area posterior 9-46 ventral $0.044$ ASD 1.921
-Area 43 Area posterior 9-46 ventral $0.044$ ASD 1.920
-Posterior InferoTemporal complex Area posterior 9-46 ventral $0.045$ ASD 1.915
-Area 7PC Area posterior 9-46 ventral $0.045$ ASD 1.913
-Area STSv posterior Area posterior 9-46 ventral $0.044$ ASD 1.912
-Area V3B Area posterior 9-46 ventral $0.046$ ASD 1.908
-Area Lateral IntraParietal dorsal Area posterior 9-46 ventral $0.045$ ASD 1.905
-Area Lateral Occipital 1 Area posterior 9-46 ventral $0.046$ ASD 1.904
-Area Posterior Insular 1 Area posterior 9-46 ventral $0.046$ ASD 1.900
-Dorsal Transitional Visual Area Area posterior 9-46 ventral $0.046$ ASD 1.900
-Fusiform Face Complex Area posterior 9-46 ventral $0.047$ ASD 1.890
-Second Visual Area Area posterior 9-46 ventral $0.048$ ASD 1.887
-PeriSylvian Language Area Area posterior 9-46 ventral $0.047$ ASD 1.887
-Area s32 Area posterior 9-46 ventral $0.047$ ASD 1.885
-Area 46 Area posterior 9-46 ventral $0.047$ ASD 1.884
-Hippocampus Area posterior 9-46 ventral $0.048$ ASD 1.880
-VentroMedial Visual Area 2 Area posterior 9-46 ventral $0.048$ ASD 1.879
-Area 8Av Area posterior 9-46 ventral $0.048$ ASD 1.877
-Area OP4-PV Area posterior 9-46 ventral $0.048$ ASD 1.877
-Posterior Insular Area 2 Area posterior 9-46 ventral $0.048$ ASD 1.876
-Area TemporoParietoOccipital Junction 2 Area posterior 9-46 ventral $0.047$ ASD 1.876
-Area 55b Area posterior 9-46 ventral $0.048$ ASD 1.876
-Area 23d Area posterior 9-46 ventral $0.048$ ASD 1.874
-PreSubiculum Area posterior 9-46 ventral $0.048$ ASD 1.871
-Anterior IntraParietal Area Area posterior 9-46 ventral $0.048$ ASD 1.870
-Area PFcm Area posterior 9-46 ventral $0.049$ ASD 1.869
-Ventral Area 6 Area posterior 9-46 ventral $0.049$ ASD 1.867
-Area 33 prime Area posterior 9-46 ventral $0.049$ ASD 1.866
-Area STSd posterior Area posterior 9-46 ventral $0.048$ ASD 1.865
-Parieto-Occipital Sulcus Area 1 Area posterior 9-46 ventral $0.049$ ASD 1.862
-Seventh Visual Area Area posterior 9-46 ventral $0.050$ ASD 1.860
-Area OP1-SII Area posterior 9-46 ventral $0.049$ ASD 1.860
-Area STSd posterior Area Lateral IntraParietal dorsal $0.049$ ASD 1.838
+Frontal Opercular Area 3 L Polar 10p R HC 0.003 2.128
+Frontal Opercular Area 2 L Polar 10p R HC 0.003 2.123
+Area V3A R Polar 10p R HC 0.003 2.117
+Area posterior 47r L Polar 10p R HC 0.004 2.109
+IntraParietal Sulcus Area 1 L Polar 10p R HC 0.004 2.101
+Area OP2-3-VS R Polar 10p R HC 0.004 2.1
+Frontal Opercular Area 3 R Polar 10p R HC 0.004 2.096
+Frontal Opercular Area 2 R Polar 10p R HC 0.004 2.095
+Area 44 L Polar 10p R HC 0.004 2.093
+Para-Insular Area R Polar 10p R HC 0.004 2.093
+Eighth Visual Area R Polar 10p R HC 0.004 2.091
+Ventral IntraParietal Complex R Polar 10p R HC 0.004 2.09
+Area IFSp L Polar 10p R HC 0.004 2.087
+Area anterior 9-46v L Polar 10p R HC 0.004 2.083
+Area STGa R Polar 10p R HC 0.004 2.082
+Medial IntraParietal Area L Polar 10p R HC 0.004 2.078
+ParaHippocampal Area 1 R Polar 10p R HC 0.004 2.075
+Rostral Area 6 R Polar 10p R HC 0.004 2.074
+Area V3B L Polar 10p R HC 0.004 2.074
+Pirform Cortex L Polar 10p R HC 0.004 2.074
+Anterior 24 prime L Polar 10p R HC 0.004 2.073
+Para-Insular Area L Polar 10p R HC 0.004 2.072
+Area TF R Polar 10p R HC 0.004 2.071
+Primary Auditory Cortex R Polar 10p R HC 0.004 2.071
+Area PFt R Polar 10p R HC 0.004 2.069
+Area V4t L Polar 10p R HC 0.004 2.069
+Area OP1-SII R Polar 10p R HC 0.004 2.069
+Area V3B R Polar 10p R HC 0.004 2.069
+Pirform Cortex R Polar 10p R HC 0.004 2.068
+Area IntraParietal 0 L Polar 10p R HC 0.004 2.067
+Area Frontal Opercular 5 L Polar 10p R HC 0.004 2.064
+Frontal Opercular Area 1 R Polar 10p R HC 0.004 2.064
+Area 55b L Polar 10p R HC 0.004 2.064
+Area OP2-3-VS L Polar 10p R HC 0.004 2.064
+ProStriate Area R Polar 10p R HC 0.004 2.063
+Area STSd posterior L Polar 10p R HC 0.004 2.062
+Area IntraParietal 0 R Polar 10p R HC 0.004 2.062
+Area 8B Lateral L Polar 10p R HC 0.004 2.061
+Area STSd anterior L Polar 10p R HC 0.004 2.061
+Lateral Belt Complex R Polar 10p R HC 0.004 2.061
+Lateral Area 7P L Polar 10p R HC 0.004 2.061
+IntraParietal Sulcus Area 1 R Polar 10p R HC 0.004 2.06
+Area OP1-SII L Polar 10p R HC 0.004 2.06
+Area 2 R Polar 10p R HC 0.004 2.059
+Area IFJa R Polar 10p R HC 0.004 2.059
+Ventral IntraParietal Complex L Polar 10p R HC 0.004 2.058
+Area Posterior Insular 1 R Polar 10p R HC 0.004 2.058
+Area 7PC R Polar 10p R HC 0.004 2.058
+Area IFJp R Polar 10p R HC 0.004 2.057
+Insular Granular Complex R Polar 10p R HC 0.004 2.057
+Area IFSa L Polar 10p R HC 0.004 2.057
+Dorsal Transitional Visual Area L Polar 10p R HC 0.004 2.056
+Posterior Insular Area 2 R Polar 10p R HC 0.004 2.055
+Area 33 prime R Polar 10p R HC 0.004 2.055
+Frontal Opercular Area 4 L Polar 10p R HC 0.004 2.054
+Area Lateral IntraParietal ventral R Polar 10p R HC 0.004 2.053
+Area 8BM L Polar 10p R HC 0.004 2.053
+Ventral Visual Complex R Polar 10p R HC 0.004 2.053
+Middle Insular Area L Polar 10p R HC 0.004 2.053
+Area 9 Middle R Polar 10p R HC 0.004 2.052
+Insular Granular Complex L Polar 10p R HC 0.004 2.051
+Sixth Visual Area R Polar 10p R HC 0.004 2.051
+Lateral Area 7A R Polar 10p R HC 0.004 2.051
+Ventral Visual Complex L Polar 10p R HC 0.004 2.05
+Area TG dorsal R Polar 10p R HC 0.004 2.05
+Frontal Opercular Area 4 R Polar 10p R HC 0.004 2.05
+Anterior IntraParietal Area R Polar 10p R HC 0.004 2.05
+Hippocampus R Polar 10p R HC 0.004 2.05
+PreSubiculum L Polar 10p R HC 0.004 2.05
+Area anterior 10p L Polar 10p R HC 0.004 2.048
+Area IFJp L Polar 10p R HC 0.004 2.047
+Seventh Visual Area R Polar 10p R HC 0.004 2.047
+VentroMedial Visual Area 2 R Polar 10p R HC 0.004 2.046
+Eighth Visual Area L Polar 10p R HC 0.004 2.045
+Area TA2 R Polar 10p R HC 0.004 2.044
+Fourth Visual Area R Polar 10p R HC 0.004 2.044
+Area PFcm R Polar 10p R HC 0.004 2.044
+Area PFm Complex R Polar 10p R HC 0.004 2.044
+Area PFm Complex L Polar 10p R HC 0.004 2.044
+Area TE1 anterior R Polar 10p R HC 0.004 2.043
+Posterior Insular Area 2 L Polar 10p R HC 0.004 2.043
+Area 8Ad L Polar 10p R HC 0.004 2.043
+Area dorsal 32 R Polar 10p R HC 0.004 2.042
+Area V4t R Polar 10p R HC 0.004 2.042
+Frontal Eye Fields L Polar 10p R HC 0.004 2.041
+Posterior InferoTemporal complex L Polar 10p R HC 0.004 2.041
+Area Lateral IntraParietal dorsal R Polar 10p R HC 0.004 2.041
+Area PF Complex R Polar 10p R HC 0.004 2.04
+Area 9 Middle L Polar 10p R HC 0.004 2.04
+Area STSd anterior R Polar 10p R HC 0.004 2.04
+Area 43 L Polar 10p R HC 0.004 2.04
+VentroMedial Visual Area 3 L Polar 10p R HC 0.004 2.04
+Area 8Av L Polar 10p R HC 0.004 2.039
+Area 6mp L Polar 10p R HC 0.004 2.039
+Area TE2 anterior R Polar 10p R HC 0.004 2.039
+Area 1 L Polar 10p R HC 0.004 2.037
+Area V6A L Polar 10p R HC 0.005 2.037
+ParaBelt Complex R Polar 10p R HC 0.005 2.037
+Area 44 R Polar 10p R HC 0.005 2.036
+VentroMedial Visual Area 3 R Polar 10p R HC 0.005 2.036
 Supplementary Table 22: Top connectivity features in ultralow frequency range of ABIDE ASD prediction
-Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
-Medial Belt Complex Area 7PC $0.002$ HC 3.728
-Area TA2 (temporal region A, area 2) Area 7PC $0.005$ HC 3.157
-ParaBelt Complex Area Frontal Opercular 5 $0.005$ HC 3.119
-Dorsal area 6 Area Frontal Opercular 5 $0.009$ HC 2.781
-ParaHippocampal Area 1 Area 7PC $0.009$ HC 2.752
-Auditory 4 Complex Area 7PC $0.011$ HC 2.739
-ParaHippocampal Area 1 Area posterior 9-46 ventral $0.010$ HC 2.729
-Auditory 4 Complex Area Frontal Opercular 5 $0.011$ HC 2.696
-Area PH Area Frontal Opercular 5 $0.004$ HC 2.570
-Superior 6-8 Transitional Area Area 7PC $0.013$ HC 2.551
-Posterior Insular Area 2 Area Frontal Opercular 5 $0.014$ HC 2.546
-Medial Belt Complex Area Lateral IntraParietal dorsal $0.005$ HC 2.541
-ParaHippocampal Area 1 Area IFJa $0.014$ HC 2.520
-Medial Belt Complex Area posterior 9-46 ventral $0.014$ HC 2.518
-Area anterior 9-46 ventral Area 7PC $0.014$ HC 2.515
-Area OP2-3-VS (Ventral-Superior part of 2-3 area of Parietal Operculum) Area Frontal Opercular 5 $0.014$ HC 2.503
-Medial Belt Complex Area IFJa $0.017$ HC 2.444
-Area 8B Lateral Area 7PC $0.006$ HC 2.440
-Area 6 medial anterior Area Lateral IntraParietal dorsal $0.005$ HC 2.436
-ParaBelt Complex Area 7PC $0.017$ HC 2.433
-Middle Temporal Area Area 7PC $0.017$ HC 2.402
-Area 52 Area Lateral IntraParietal dorsal $0.006$ HC 2.398
-Area 46 Area Frontal Opercular 5 $0.018$ HC 2.394
-Medial Belt Complex Frontal Opercular Area 2 $0.018$ HC 2.379
-Dorsal area 6 Area Lateral IntraParietal dorsal $0.006$ HC 2.374
-Area 6 medial posterior Area Lateral IntraParietal dorsal $0.019$ HC 2.364
-Posterior Insular Area 2 Area 7PC $0.019$ HC 2.357
-Area OP4-PV (Posterior-Ventral part of area 4 of Parietal Operculum) Area Frontal Opercular 5 $0.020$ HC 2.335
-Inferior 6-8 Transitional Area Area Frontal Opercular 5 $0.021$ HC 2.328
-Frontal Opercular Area 1 Area Lateral IntraParietal dorsal $0.006$ HC 2.319
-Area posterior 10p Area 7PC $0.021$ HC 2.310
-Primary Auditory Cortex Area 7PC $0.023$ HC 2.308
-Dorsal area 6 Area 7PC $0.021$ HC 2.305
-Anterior Agranular Insula Complex Area Frontal Opercular 5 $0.023$ HC 2.290
-ParaHippocampal Area 1 Area Lateral IntraParietal dorsal $0.007$ HC 2.287
-Auditory 4 Complex Area IFJa $0.022$ HC 2.276
-Auditory 4 Complex Area posterior 9-46 ventral $0.024$ HC 2.267
-Area IFJa Area Frontal Opercular 5 $0.022$ HC 2.265
-Area 9 Middle Area PGs $0.022$ HC 2.264
-Area posterior 9-46 ventral Area Frontal Opercular 5 $0.024$ HC 2.256
-Area anterior 9-46 ventral Area Frontal Opercular 5 $0.023$ HC 2.255
-Area PF Opercular Area Frontal Opercular 5 $0.023$ HC 2.254
-Frontal Opercular Area 1 Area 7PC $0.023$ HC 2.248
-PreSubiculum Area PGs $0.023$ HC 2.247
-Superior 6-8 Transitional Area Area Lateral IntraParietal dorsal $0.024$ HC 2.241
-Area TA2 (temporal region A, area 2) Area Lateral IntraParietal dorsal $0.023$ HC 2.241
-Area OP1-SII Area Frontal Opercular 5 $0.024$ HC 2.229
-Area s32 Area 7PC $0.024$ HC 2.228
-Medial Belt Complex Area Frontal Opercular 5 $0.008$ HC 2.226
-Area PH Area IFJa $0.024$ HC 2.224
-Area Posterior Insular 1 Area Lateral IntraParietal dorsal $0.024$ HC 2.215
-VentroMedial Visual Area 1 Area Frontal Opercular 5 $0.025$ HC 2.213
-Area 33 prime Area Frontal Opercular 5 $0.025$ HC 2.211
-Area 43 Area IFJa $0.026$ HC 2.209
-Area OP2-3-VS (Ventral-Superior part of 2-3 area of Parietal Operculum) Area Lateral IntraParietal dorsal $0.025$ HC 2.206
-Frontal Opercular Area 1 Area Frontal Opercular 5 $0.025$ HC 2.205
-Pirform Cortex Area Lateral IntraParietal dorsal $0.009$ HC 2.187
-Area 9 anterior Area PGs $0.026$ HC 2.183
-Area 8B Lateral Area PGs $0.026$ HC 2.183
-Area PF Complex Area Frontal Opercular 5 $0.027$ HC 2.179
-Anterior Agranular Insula Complex Area posterior 9-46 ventral $0.027$ HC 2.176
-Area 1 Area Frontal Opercular 5 $0.027$ HC 2.173
-Premotor Eye Field Area Frontal Opercular 5 $0.027$ HC 2.173
-ParaHippocampal Area 1 Area Frontal Opercular 5 $0.027$ HC 2.172
-Area PFt Area Frontal Opercular 5 $0.027$ HC 2.169
-Area 13 lateral Area Frontal Opercular 5 $0.027$ HC 2.163
-Rostral Area 6 Area 7PC $0.028$ HC 2.159
-Anterior Agranular Insula Complex Area 7PC $0.029$ HC 2.143
-Area 6 medial posterior Area PGs $0.028$ HC 2.139
-Area 43 Area Frontal Opercular 5 $0.028$ HC 2.138
-Area OP4-PV (Posterior-Ventral part of area 4 of Parietal Operculum) Area Lateral IntraParietal dorsal $0.028$ HC 2.138
-Posterior Insular Area 2 Area posterior 9-46 ventral $0.029$ HC 2.128
-ParaHippocampal Area 1 Sixth Visual Area $0.030$ HC 2.114
-Parieto-Occipital Sulcus Area 1 Area Lateral IntraParietal dorsal $0.009$ HC 2.111
-Area posterior 10p Area Frontal Opercular 5 $0.031$ HC 2.108
-Ventral Area 6 Area Frontal Opercular 5 $0.030$ HC 2.106
-Area STSd anterior Area Lateral IntraParietal dorsal $0.009$ HC 2.100
-Area anterior 9-46 ventral Area Lateral IntraParietal dorsal $0.031$ HC 2.093
-Area Posterior Insular 1 Area PGs $0.031$ HC 2.093
-Rostral Area 6 Area Frontal Opercular 5 $0.010$ HC 2.091
-Area FST Area Frontal Opercular 5 $0.031$ HC 2.085
-Area Superior temporal sulcus ventral anterior Area 7PC $0.031$ HC 2.081
-Area Posterior Insular 1 Area posterior 9-46 ventral $0.032$ HC 2.076
-Area TE1 Middle Area 7PC $0.032$ HC 2.070
-Area 6 medial posterior Area 7PC $0.032$ HC 2.069
-Area Posterior Insular 1 Area 7PC $0.032$ HC 2.067
-Area Posterior Insular 1 Area Frontal Opercular 5 $0.033$ HC 2.063
-Area TA2 (temporal region A, area 2) Area PGs $0.032$ HC 2.062
-PreSubiculum Area Lateral IntraParietal dorsal $0.011$ HC 2.055
-Area 43 Area 7PC $0.033$ HC 2.048
-ParaHippocampal Area 2 Area 7PC $0.034$ HC 2.045
-Inferior 6-8 Transitional Area Area Lateral IntraParietal dorsal $0.034$ HC 2.044
-ParaHippocampal Area 3 Area Lateral IntraParietal dorsal $0.012$ HC 2.043
-Area 10 rostral Area PGs $0.034$ HC 2.041
-ParaBelt Complex Area Lateral IntraParietal dorsal $0.034$ HC 2.041
-Medial Superior Temporal Area Area Frontal Opercular 5 $0.034$ HC 2.041
-Superior 6-8 Transitional Area Area PGs $0.034$ HC 2.040
-Entorhinal Cortex Area Lateral IntraParietal dorsal $0.034$ HC 2.037
-VentroMedial Visual Area 1 Area IFJa $0.034$ HC 2.036
-Area Lateral Occipital 3 Area Lateral IntraParietal dorsal $0.012$ HC 2.033
+Region 1 Region 2 Disorder/HC Corrected p-value Cohen’s d
+Area PHT L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 5.637
+Area 9 Posterior R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 5.334
+Area 8Ad R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.99
+Area 8B Lateral L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.835
+Orbital Frontal Complex L Area TF R ASD $<0.001$ 3.767
+Superior Frontal Language Area L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.566
+Superior 6-8 Transitional Area L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.566
+Area TE2 posterior L Area Posterior 24 prime R ASD $<0.001$ 3.549
+RetroInsular Cortex R Area anterior 32 prime L ASD $<0.001$ 3.54
+Frontal Opercular Area 2 L Ventral IntraParietal Complex R ASD $<0.001$ 3.499
+Fourth Visual Area L Area anterior 32 prime L ASD $<0.001$ 3.494
+Lateral Area 7A R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.38
+Frontal Eye Fields L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.355
+Area V3B R Area 9-46d L ASD 0.002 3.339
+Medial Area 7P L Area anterior 32 prime L ASD 0.002 3.328
+Fourth Visual Area L Area PFm Complex L ASD 0.002 3.295
+Sixth Visual Area R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.252
+Area posterior 9-46v R Area anterior 32 prime L ASD $<0.001$ 3.201
+Second Visual Area R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.176
+Area 8BM R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.173
+Area STSd posterior L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.169
+Second Visual Area R Area 23d R ASD $<0.001$ 3.143
+IntraParietal Sulcus Area 1 R Area 23d R ASD $<0.001$ 3.128
+Lateral Area 7P L RetroInsular Cortex L ASD $<0.001$ 3.08
+Lateral Area 7P R RetroInsular Cortex L ASD $<0.001$ 3.076
+Medial Area 7P L Primary Sensory Cortex R ASD $<0.001$ 3.064
+Second Visual Area L Area anterior 32 prime L ASD $<0.001$ 3.024
+Second Visual Area L Area PFm Complex L ASD $<0.001$ 3.012
+Area posterior 9-46v L RetroInsular Cortex L ASD $<0.001$ 3.009
+Area TE2 posterior L Area 5m R ASD $<0.001$ 2.983
+Area 5m L Area anterior 32 prime L ASD $<0.001$ 2.966
+Area PH L Area anterior 32 prime L ASD $<0.001$ 2.963
+Area 2 R Area anterior 32 prime L ASD $<0.001$ 2.962
+Area V3B R Area anterior 32 prime L ASD $<0.001$ 2.95
+Area 9 Middle L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.928
+Area 31a R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.921
+Perirhinal Ectorhinal Cortex R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.919
+Area 6m anterior R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.913
+Area PH L Area PFm Complex L ASD $<0.001$ 2.878
+Ventral Area 6 R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.867
+Parieto-Occipital Sulcus Area 1 L Medial Superior Temporal Area R ASD $<0.001$ 2.859
+Ventral Visual Complex R Area anterior 32 prime L ASD $<0.001$ 2.859
+Area anterior 32 prime R Area 23d R ASD $<0.001$ 2.844
+Area TemporoParietoOccipital Junction 3 R Area anterior 32 prime L ASD $<0.001$ 2.838
+ParaHippocampal Area 2 R Area anterior 32 prime L ASD $<0.001$ 2.821
+Area 9-46d R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.82
+Area V3B R Area Posterior 24 prime R ASD $<0.001$ 2.801
+Area PHT L Area dorsal 23 a+b L ASD $<0.001$ 2.801
+Area 47m L Area Lateral Occipital 1 L ASD $<0.001$ 2.796
+Ventral IntraParietal Complex L Medial Superior Temporal Area L ASD $<0.001$ 2.796
+Area IntraParietal 2 R Area anterior 32 prime L ASD $<0.001$ 2.789
+Area STSv posterior L Area 23d R ASD 0.007 2.781
+VentroMedial Visual Area 1 L Primary Sensory Cortex R ASD $<0.001$ 2.775
+Area posterior 24 L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.774
+Medial Area 7A R Primary Sensory Cortex R ASD $<0.001$ 2.767
+Primary Motor Cortex R Area anterior 32 prime L ASD 0.001 2.762
+Medial Area 7A L Primary Sensory Cortex R ASD 0.001 2.753
+Medial Area 7P L Area PFm Complex L ASD $<0.001$ 2.751
+Dorsal Transitional Visual Area L Medial Superior Temporal Area R ASD 0.001 2.744
+Area Lateral Occipital 2 R RetroInsular Cortex L ASD 0.001 2.736
+Area Lateral Occipital 2 R Medial Superior Temporal Area R ASD 0.001 2.734
+Area dorsal 32 L Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.723
+Area TG Ventral L RetroInsular Cortex L ASD 0.001 2.711
+Area Lateral IntraParietal dorsal R Area anterior 32 prime L ASD 0.001 2.696
+Primary Visual Cortex R Area 23d R ASD 0.001 2.688
+Area p32 R Medial Superior Temporal Area R ASD 0.005 2.685
+Area STSv posterior L Area PFm Complex L ASD 0.005 2.683
+Area PFt R Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.681
+Area 31a R Area PFm Complex L ASD 0.001 2.672
+Area TemporoParietoOccipital Junction 3 R Area PFm Complex L ASD 0.001 2.669
+Dorsal Transitional Visual Area R Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.658
+Frontal Opercular Area 3 L Area 8Av L ASD 0.001 2.645
+Anterior IntraParietal Area R Area anterior 32 prime L ASD 0.001 2.627
+Dorsal Transitional Visual Area R Area PFm Complex L ASD 0.001 2.62
+Posterior Insular Area 2 R Area anterior 32 prime L ASD 0.001 2.618
+Primary Visual Cortex L Area 23d R ASD 0.001 2.615
+Medial Area 7A L Area anterior 32 prime L ASD 0.001 2.608
+Primary Visual Cortex L Area PFm Complex L ASD 0.001 2.606
+Area 8Av L Primary Sensory Cortex R ASD 0.001 2.599
+Area TE2 posterior L RetroInsular Cortex L ASD 0.001 2.583
+Area Lateral Occipital 2 R Area dorsal 23 a+b L ASD 0.001 2.58
+RetroInsular Cortex R Area PFm Complex L ASD 0.001 2.578
+Area 1 R Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.576
+Primary Sensory Cortex R Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.572
+Area STSv anterior L Area 6 anterior L ASD 0.001 2.57
+Area V3B R Medial Belt Complex L ASD 0.001 2.568
+Seventh Visual Area L Medial Superior Temporal Area L ASD 0.003 2.565
+Area V3A R Ventral IntraParietal Complex R ASD 0.001 2.565
+Area STGa L Area 23d R ASD 0.001 2.558
+Area 2 L Ventral IntraParietal Complex R ASD 0.001 2.55
+Para-Insular Area R Area 23d R ASD 0.001 2.548
+Inferior 6-8 Transitional Area L RetroInsular Cortex L ASD 0.001 2.548
+Seventh Visual Area L Area PFm Complex L ASD 0.001 2.546
+Dorsal area 6 R Area 9-46d L ASD 0.001 2.541
+Area 8BM L Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.536
+Area TE2 posterior L Area Posterior 24 prime L ASD 0.001 2.535
+Area 52 L Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.533
+Area dorsal 32 R Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.532
+Ventral Area 6 R Area 23d R ASD 0.001 2.53
+Area 9 Posterior L Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.53